Linux Consultants HOWTO

Mr. Poet <poet@linuxports.com> v9.46, 12 September 1999

This document contains a listing of companies providing commercial Linux related support. If you want to find a Linux consultant or consulting firm in your area, this listing will probably be of help for you. It is maintained by Poet poet@linuxports.com. This document is also completely indexed, databased and searchable at the LinuxPorts website. If you would like information on advertising on the Consultants database pages please contact us at poet@linuxports.com.

Note to the reader:

Please send comments to Mr. Poet poet@linuxports.com.

Mr. Poet

About this document

This is the Linux Consultants HOWTO. It is a listing of companies providing commercial Linux related support. If you contact any companies listed in this document, please mention the Linux Consultants HOWTO.

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the Coordinator Tim Bynum linux-howto@metalab.unc.edu.

Linux Consultants HOWTO can be found on the World Wide Web at http://www.linuxports.com/. New versions of the Linux Consultants HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date!

This site also contains all information needed to submit to this howto or search this howto via a database.

Companies providing Linux support are invited to fill out the form at LinuxPorts Please note that we have temporarily stopped accepting 8-bit characters (German umlauts, etc).

Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999 Mr. Poet and LinuxPorts

A verbatim copy may be reproduced or distributed in any electronic medium without permission of the author. Physical medium copies, reproductions or distributions for commercial use must have explicit permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Linux Consultants HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy.

Commercial redistribution is allowed and encouraged; however, the author must be notified of any such distributions and all copyright information must be intact.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Tim Bynum, the Linux HOWTO coordinator, at linux-howto@metalab.unc.edu.

1 Consultant Listings

1.1 Consultant ID: Argentina, BASystem

Manuel ugarte 2328
Buenos Aires, NON United States 1428
Argentina

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Not Applicable Company Phone Number:

54-11 4781-9710 Company Fax:

54-11 4789-9228 Support Phone Number:

54-11 4781-9710 Contact Email Address:

clau@basystem.com.ar Company URL:

http://www.basystem.com.ar Primary Contact:

Norberto Mayo

1.2 Consultant ID: Argentina, CompuSoft

     Drago 23, Piso 5, Oficina 12
     Bahia Blanca, NON United States 8000
     Argentina

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems System Administration System Security Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not_Applicable Company Phone Number:

+54(291) 455-3570 Company Fax:

+54(291) 453-6735 Support Phone Number:

+54(291) 15 648-9284 Contact Email Address:

goburastero@arnet.com.ar Company URL:

http://www.arnet.com.ar Primary Contact:

Guillermo Burastero

1.3 Consultant ID: Argentina, CyberWare Information Technology

Zeballos 2774
Rosario, NON United States 2000
Argentina

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(54)(0341) 156-40-14 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

soporte@cyberware.com.ar Company URL:

http://www.cyberware.com.ar Primary Contact:

Javier Kohan

1.4 Consultant ID: Argentina, Generic Soft

     Defensa 767 ofic. 20
     Buenos Aires, NON United States 1065
     Argentina

Consulting Specialties:

CIFS (Samba) Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

5411-4300-7294 Company Fax:

5411-4300-7294 Support Phone Number:

5411-4300-7294 Contact Email Address:

soporte@generic.findhere.com Company URL:

http://generic.findhere.com Primary Contact:

Pablo Lazaro

1.5 Consultant ID: Argentina, Jorge Forte

     Calle 47 Nro 826 Dep 9
     La Plata, NON United States 1900
     Argentina

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

542214821914 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

soporte@forte.com.ar Company URL:

http://www.forte.com.ar Primary Contact:

Jorge Forte

1.6 Consultant ID: Argentina, NIXE S.R.L.

     Pte. Luis S. Peña 366 4 A
     Buenos Aires, NON United States 1110
     Argentina

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+54 (11) 4383-7831 Company Fax:

+54 (11) 4383-7831 Support Phone Number:

+54 (11) 15-4444-832 Contact Email Address:

support@nixe.com.ar Company URL:

http://www.nixe.com.ar Primary Contact:

Gustavo Gasparrini

1.7 Consultant ID: Argentina, Nixe S.R.L.

     Pte. Luis Saenz Pe#a 366 4A
     Buenos Aires, NON United States 1110
     Argentina

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+54 (11) 4383-7831 Company Fax:

+54 (11) 4383-7831 Support Phone Number:

+54 (11) 4383-7831 Contact Email Address:

support@nixe.com.ar Company URL:

http://Not_Applicable Primary Contact:

Gustavo Gasparrini

1.8 Consultant ID: Australia, Computer Clinic (WA) Pty Ltd

     4/226 Main St Osborne Park
     Perth WA, NON United States 6017
     Australia

Consulting Specialties:

Custom Systems Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Caldera Company Phone Number:

08 94406611 Company Fax:

08 94406610 Support Phone Number:

08 94406611 Contact Email Address:

support@cclinic.com.au Company URL:

http://www.cclinic.com.au Primary Contact:

Colin or Garth Hewitt or Atkinson

1.9 Consultant ID: Australia, Cybersource Pty. Ltd.

Level 8, 140 Queen Street
Melbourne, NON United States 3000
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+61 3 9642 5997 Company Fax:

+61 3 9642 5998 Support Phone Number:

+61 3 9642 5997 Contact Email Address:

rh-support@cyber.com.au Company URL:

http://www.cyber.com.au Primary Contact:

Con Zymaris

1.10 Consultant ID: Australia, Cyberspace Corporation

3A/426 Burwood Highway
Melbourne, NON United States 3152
Australia

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not Applicable Company Phone Number:

+61 (3) 9887 4700 Company Fax:

+61 (3) 9887 2756 Support Phone Number:

+61 (3) 9887 4700 Contact Email Address:

support@cyberspace.net.au Company URL:

http://www.cyberspace.net.au Primary Contact:

Keith Chamberlain

1.11 Consultant ID: Australia, DataSource Pty. Ltd

Not_Applicable
SYDNEY, NON United States 2570
Australia

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61418675892 Company Fax:

Not_Applicable Support Phone Number:

+6146556792 Contact Email Address:

Linux@sysmon.net Company URL:

http://www.sysmon.net Primary Contact:

Steve Deadman

1.12 Consultant ID: Australia, EMUSYS Unix Consulting

20 Ivanhoe Rd
Croydon, NON United States 2132
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 2 9798 7604 Company Fax:

+61 2 9798 2854 Support Phone Number:

0500 500 EMU Contact Email Address:

support@emusys.com.au Company URL:

http://www.emusys.com.au Primary Contact:

Anthony Rumble

1.13 Consultant ID: Australia, Echo Labs

     29 Weld St, Nedlands
     Perth, NON United States 6009
     Australia

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 41 356 0008 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

echo@iinet.net.au Company URL:

http://www.iinet.net.au/˜ echo Primary Contact:

Ralph Billes

1.14 Consultant ID: Australia, Farrow Norris

     58 St Georges Pde
     Sydney, NON United States 2220
     Australia

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 4 1724 3183 Company Fax:

+61 2 9546 4468 Support Phone Number:

+61 4 1724 3183 Contact Email Address:

support@fn.com.au Company URL:

http://http://www.fn.com.au/ Primary Contact:

James Farrow

1.15 Consultant ID: Australia, Hinterlands Consultancy I.T.

     1 Crusader Rd
     Galston NSW., NON United States 2159
     Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Intel and Alpha Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 2 9653 2837 Company Fax:

+61 2 9653 2937 Support Phone Number:

+61 2 9653 2837 Contact Email Address:

support@hinterlands.com.au Company URL:

http://www.hinterlands.com.au Primary Contact:

Chris Blown

1.16 Consultant ID: Australia, Informed Technology Pty Ltd

136/138 Railway Pde
West Leederville, NON United States 6007
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Debian Sparc Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61-8-9380-4244 Company Fax:

+61-8-9380-4354 Support Phone Number:

+61-8-9380-4244 Contact Email Address:

support@it.net.au Company URL:

http://www.it.net.au Primary Contact:

Andrew Howell

1.17 Consultant ID: Australia, John Pearson

     PO Box 3412 Rundle Mall
     Adelaide, NON United States 5000
     Australia

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 407391169 Company Fax:

+61 883644231 Support Phone Number:

+61 407391169 Contact Email Address:

huiac@camtech.net.au Company URL:

http://Not_Applicable Primary Contact:

John Pearson

1.18 Consultant ID: Australia, LINSUP.COM

P. O. Box 689
St. Ives, NON United States 2075
Australia

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Debian Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 2 9144-6131 Company Fax:

+61 2 9144-6138 Support Phone Number:

+61 2 9144-6131 Contact Email Address:

info@linsup.com Company URL:

http://www.linsup.com Primary Contact:

Richard Ames

1.19 Consultant ID: Australia, Linux Garden

     Unit 4, 142 Culloden Road, Marsfield
     Sydney, NSW, NON United States 2122
     Australia

Consulting Specialties:

Custom Programming Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

61 0417 869242 Company Fax:

Not_Applicable Support Phone Number:

61 0417 869242 Contact Email Address:

shyam@linuxgarden.com Company URL:

http://www.linuxgarden.com Primary Contact:

Shyam Govardhan

1.20 Consultant ID: Australia, Studio of Arts And Sciences (SAAS)

PO Box 214 Mosman
Sydney, NON United States NSW 2088
Australia

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Darwin and all MAC Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61-2-9453-9100 Company Fax:

+61-2-9453-9101 Support Phone Number:

+61-2-9453-9100 Contact Email Address:

saas@saas.nsw.edu.au Company URL:

http://www.saas.nsw.edu.au Primary Contact:

Roger Buck

1.21 Consultant ID: Australia, Sydnet Group Pty Ltd

PO Box 4664
North Rocks, NON United States 2151
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 2 9873 6400 Company Fax:

+61 2 9873 6411 Support Phone Number:

+61 2 9873 6400 Contact Email Address:

support@syd.net.au Company URL:

http://www.syd.net.au Primary Contact:

Mark Lipscombe

1.22 Consultant ID: Australia, TereDonn Computer Engineering

     Suite 1, 128 Bowen St, Spring Hill
     BRISBANE, ZZ 4004
     Australia

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

ALL Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61 7 32369366 Company Fax:

+61 7 32369930 Support Phone Number:

+61 414 663053 Contact Email Address:

service@tdce.com.au Company URL:

http://www.tdce.com.au Primary Contact:

Terence Giufre-Sweetser

1.23 Consultant ID: Australia, WWWalker Web Development

     PO Box 288
     Wentworthville, NON United States 2145
     Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

61-412-405-727 Company Fax:

61-2-97772058 Support Phone Number:

61-412-405727 Contact Email Address:

dwight@zip.com.au Company URL:

http://www.wwwalker.com.au Primary Contact:

Dwight Walker

1.24 Consultant ID: Australia, Win International Networking

Lisle Street
Mount Claremont, NON United States 6010
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Mandrake Specialty Distribution:

Other Other Specialty Distribution:

various/hand tooled Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+61-409-65535-9 Company Fax:

Not_Applicable Support Phone Number:

+61-409-65535-9 Contact Email Address:

leon@brooks.smileys.net Company URL:

http://Not_Applicable Primary Contact:

Leon Brooks

1.25 Consultant ID: Australia, linsup.com

P. O. Box 689
St. Ives, NON United States 2075
Australia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

61 2 9144-6131 Company Fax:

61 2 9144-6138 Support Phone Number:

61 2 9144-6131 Contact Email Address:

info@linsup.com Company URL:

http://www.linsup.com Primary Contact:

Richard Ames

1.26 Consultant ID: Austria, Bernhard Zwischenbrugger

     Lerchenfeldestrasse 88-90/17
     Vienna, NON United States 1080
     Austria

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

auto install one dis Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43/1/4022752 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

bzwische@email.archlab.tuwien.ac.at Company URL:

http://Not_Applicable Primary Contact:

Bernhard Zwischenbrugger

1.27 Consultant ID: Austria, Bitsniz Data Systems

     Kreuzgasse 60
     Vienna, NON United States A-1180
     Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Linux Certifications:

Not Applicable Company Phone Number:

+43-(0)1-4065098 Company Fax:

+43-(0)1-4065098-99 Support Phone Number:

+43-(0)1-4065098-20 Contact Email Address:

office@bitsniz.at Company URL:

http://www.bitsniz.at Primary Contact:

Franz Antonicek

1.28 Consultant ID: Austria, Dialog Data GmbH

Steyrergasse 76
Graz, NON United States A-8010
Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43(316)810091 Company Fax:

+43(316)826348 Support Phone Number:

+43(316)810091 Contact Email Address:

support@dialogdata.com Company URL:

http://www.dialogdata.com/ Primary Contact:

Norbert Friesl

1.29 Consultant ID: Austria, EDV-Beratung F.Ganter

     Liebiggasse 19
     Graz, NON United States A-8010
     Austria

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet (Web) Programming System Administration Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43 699 110 21 621 Company Fax:

+43 699 310 21 621 Support Phone Number:

Not_Applicable Contact Email Address:

ganter@ganter.at Company URL:

http://www.ganter.at Primary Contact:

Fritz Ganter

1.30 Consultant ID: Austria, Horus GmbH

     Jakob-Haringer-Straße 8
     Salzburg, NON United States A-5020
     Austria

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+43 (0)662 450363-0 Company Fax:

+43 (0)662 450363-30 Support Phone Number:

+43 (0)662 450363-0 Contact Email Address:

info@horus.com Company URL:

http://www.horus.com Primary Contact:

Thomas ’Dune’ Freina

1.31 Consultant ID: Austria, LINUX@work

     Hauptstr. 99
     Hagenberg, NON United States 4232
     Austria

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43 7236 2200 Company Fax:

+43 7236 2200 22 Support Phone Number:

+43 7236 2200 13 Contact Email Address:

support@linuxatwork.at Company URL:

http://www.linuxatwork.at Primary Contact:

Andreas Hupfau

1.32 Consultant ID: Austria, LinzNet

Floetzerweg 150
Linz, NON United States 4030
Austria

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43 732 370700 Company Fax:

+43 732 376554 Support Phone Number:

+43 732 370700 Contact Email Address:

support@linznet.at Company URL:

http://www.linznet.at Primary Contact:

Rainer Skarke

1.33 Consultant ID: Austria, M.T.G. Handelsges.m.b.H.

Raiffeisenstr. 16/9
Zwölfaxing, NON United States 2320
Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

custom distributions Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43/676/4091256 Company Fax:

+43/1/7065299 Support Phone Number:

+43/1/7070750 Contact Email Address:

office@mtg.co.at Company URL:

http://https://www.mtg.co.at/ Primary Contact:

Andreas Kostyrka

1.34 Consultant ID: Austria, PSI Providing Services for the Internet

Hütteldorfer Straße 193/24
Wien, NON United States 1140
Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Suse Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43-1-9117757 Company Fax:

+43-1-9117757 Support Phone Number:

+43-1-9117757 Contact Email Address:

info@psi.co.at Company URL:

http://www.psi.co.at Primary Contact:

Raphael Wegmann

1.35 Consultant ID: Austria, communication.center.one

     Hetzendorfer Str.23/20
     Vienna, ZZ A-1120
     Austria

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43-1-8045753 Company Fax:

+43-1-8031553 Support Phone Number:

+43-676-477-70-70 Contact Email Address:

support@ccone.at Company URL:

http://www.ccone.at Primary Contact:

Gerhard Beck

1.36 Consultant ID: Austria, g.a.m.s. edv-dienstleistungen gmbh

Stiegergasse 15-17/8
Vienna, NON United States 1150
Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43 1 895 84 99 Company Fax:

+43 1 895 84 99 60 Support Phone Number:

+43 1 895 84 99 Contact Email Address:

linux@gams.at Company URL:

http://www.gams.at Primary Contact:

Andreas Kainz

1.37 Consultant ID: Austria, iDAS

     Gasgasse 13/5
     Vienna, NON United States A-1150
     Austria

Consulting Specialties:

CIFS (Samba) Custom Systems Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

++43-664-2630131 Company Fax:

++43-1-8939122-26 Support Phone Number:

++43-664-2630131 Contact Email Address:

office@iDAS.co.at Company URL:

http://www.iDAS.co.at Primary Contact:

Thomas Kirchtag

1.38 Consultant ID: Austria, xS+S

     Karmarschgasse 51/2/20
     Wien, NON United States A-1100
     Austria

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

xS+S Linux Specialty Distribution:

Other Other Specialty Distribution:

xS+S Linux Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+43 1 6060114 0 Company Fax:

+43 1 6060114 71 Support Phone Number:

+43 1 6060114 10 Contact Email Address:

support@xss.co.at Company URL:

http://www.xss.co.at Primary Contact:

Andreas Haumer

1.39 Consultant ID: Bangladesh, Information Services Network Limited

52, New Eskaton Road TMC Building(3rd & 4th Floor)
Dhaka, NON United States 1000
Bangladesh

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

8802842785 Company Fax:

88029345460 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.bangla.net Primary Contact:

Sumon Ahmed Sabir

1.40 Consultant ID: Belgium, Alexandre Dulaunoy, Consultance OpenSource

4, rue d'Aineffe
Borlez Faimes, NON United States 4317
Belgium

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

self-made Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

3219566115 Company Fax:

3219567337 Support Phone Number:

3219566115 Contact Email Address:

adulau@unix.be.eu.org Company URL:

http://http://unix.be.eu.org/adulau Primary Contact:

Alexandre Dulaunoy

1.41 Consultant ID: Belgium, BARTH Group’s

     12, rue de l'Espoir
     Liège, NON United States 4030
     Belgium

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

32 4 367.49.16 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

BARTH.groups@skynet.be Company URL:

http://www.skynet.be Primary Contact:

Wathelet Michel

1.42 Consultant ID: Belgium, Better Access NV

Geldenaakse Vest 6
Leuven, NON United States 3000
Belgium

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

Other Other Specialty Distribution:

Custombuilt Linux So Reseller Authorizations:

Compaq HP Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32 16 29 80 45 Company Fax:

+32 16 29 80 46 Support Phone Number:

+32 16 29 80 45 Contact Email Address:

info@ba.be Company URL:

http://www.ba.be Primary Contact:

Jan Guldentops

1.43 Consultant ID: Belgium, CompuSense bvba

     Mr Vorsselmanslaan 27
     Kalmthout, NON United States B 2920
     Belgium

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security System Administration System Security Main Distribution:

RedHat Other Distribution:

SuSe Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Sparc Linux Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32 75 503503 Company Fax:

+32 3 6666579 Support Phone Number:

+32 75 503503 Contact Email Address:

herman.willekens@skynet.be Company URL:

http://www.skynet.be Primary Contact:

Herman Willekens

1.44 Consultant ID: Belgium, CoolNet

     rue Croix Claire 67
     NANDRIN, NON United States 4550
     Belgium

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Network Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32 75 93 51 71 Company Fax:

+32 4 372 01 38 Support Phone Number:

Not_Applicable Contact Email Address:

marc.brocha@coolnet.be Company URL:

http://www.coolnet.be Primary Contact:

marc brocha Not_Applicable

1.45 Consultant ID: Belgium, Double Barrel Consultancy & Productions

Sportstraat 28
Gent, ZZ 9000
Belgium

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32-9-2227764 Company Fax:

+32-9-2224976 Support Phone Number:

+32-9-2227764 Contact Email Address:

help@double-barrel.be Company URL:

http://www.double-barrel.be Primary Contact:

Michael Vergallen

1.46 Consultant ID: Belgium, FD Informatica Consulting

     Merellaan 42
     Kapellen, NON United States B2950
     Belgium

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32 (75) 96 04 93 Company Fax:

+32 (3) 605 10 56 Support Phone Number:

(only on request) Contact Email Address:

support@denkens.com Company URL:

http://www.denkens.com Primary Contact:

Frederik Denkens

1.47 Consultant ID: Belgium, SPIER bvba

     Knaptandstraat 96
     Sint-Niklaas, NON United States B9100
     Belgium

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Linux Certifications:

Not Applicable Company Phone Number:

+ 32 3 765 90 61 Company Fax:

+ 32 3 765 90 62 Support Phone Number:

+ 32 3 765 90 61 Contact Email Address:

support@spier.be Company URL:

http://www.spier.be Primary Contact:

Jan Lybeert

1.48 Consultant ID: Belgium, VirgoPlus

Rue Dartois, 1
Liège, NON United States 4000
Belgium

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32-4-253 00 59 Company Fax:

+32-4-253 00 49 Support Phone Number:

+32-4-253 00 85 Contact Email Address:

bruno@virgoplus.com Company URL:

http://www.virgoplus.com Primary Contact:

Christophe Ozer

1.49 Consultant ID: Belgium, information technology partners

     Klokhof 2
     Heusden-Zolder, NON United States 3550
     Belgium

Consulting Specialties:

E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

Suse/Caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+32 11 53 77 53 Company Fax:

+32 11 53 77 17 Support Phone Number:

+32 11 53 77 53 Contact Email Address:

guy.dillen@skynet.be Company URL:

http://www.skynet.be Primary Contact:

Guy Dillen

1.50 Consultant ID: Brazil, 4Web Internet

     R. James Watt, 142 - Cj 42
     Sao Paulo, NON United States 04576-050
     Brazil

Consulting Specialties:

E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55-11-5506-6844 Company Fax:

55-11-5506-6844 Support Phone Number:

55-11-5506-6844 Contact Email Address:

esaito@4web.com.br Company URL:

http://www.4web.com.br Primary Contact:

Eduardo Saito

1.51 Consultant ID: Brazil, Abner Graham Jacobsen

     Rua Bruno Becacici, 107
     Vitoria, NON United States 29042-0320
     Brazil

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Conectiva Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55 027 322-1837 Company Fax:

55 027 223-0562 Support Phone Number:

Not Available Contact Email Address:

ajacobsen@interlink.com.br Company URL:

http://www.interlink.com.br Primary Contact:

Not_Applicable Not_Applicable

1.52 Consultant ID: Brazil, Dextra Solucoes em Informatica

     Av Jose Souza Campos 1815/1205
     Campinas, NON United States 13025-320
     Brazil

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55-19-251-3644 Company Fax:

+55-19-253-3742 Support Phone Number:

NA Contact Email Address:

consultoria@dextra.com.br Company URL:

http://www.dextra.com.br Primary Contact:

Bill Coutinho

1.53 Consultant ID: Brazil, Dextra Solucoes

     Av J Souza Campos 1815 cj 1205
     Campinas, NON United States 13025-320
     Brazil

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55 19 251-3644 Company Fax:

55 19 253-3742 Support Phone Number:

55 19 251-3644 Contact Email Address:

consultoria@dextra.com.br Company URL:

http://www.dextra.com.br Primary Contact:

Bill Coutinho

1.54 Consultant ID: Brazil, Fernando Mauro Martins

     Rua Padre Justino, 44
     São Paulo, NON United States 05580-090
     Brazil

Consulting Specialties:

Custom Programming Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

none Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55 11 814-5679 Company Fax:

na Support Phone Number:

+55 11 814-5679 Contact Email Address:

lalo@webcom.com Company URL:

http://www.webcom.com Primary Contact:

Fernando Martins

1.55 Consultant ID: Brazil, INFOX Sistemas e Computadores Ltda

     Rua Boquim, 97 - 1. andar
     Aracaju, NON United States 49.010-280
     Brazil

Consulting Specialties:

System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55-79-211-5122 Company Fax:

55-79-211-0783 Support Phone Number:

Not_Applicable Contact Email Address:

infox@infox.com.br Company URL:

http://www.infox.com.br Primary Contact:

Fred Silva

1.56 Consultant ID: Brazil, Kernel LinuxRO

     Rua Jacaranda, 910 - Fundos
     Porto Velho, NON United States 78912-570
     Brazil

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

CONECTIVA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55699844175 Company Fax:

+55692273741 Support Phone Number:

+55699844175 Contact Email Address:

jbs@portovelho.br Company URL:

http://Not_Applicable Primary Contact:

Joao Silva

1.57 Consultant ID: Brazil, Lumina Informatica Ltda

     Caixa Postal 1393
     Blumenau, NON United States 89010-971
     Brazil

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not Applicable Company Phone Number:

+55 47 329-0065 Company Fax:

+55 47 329-0091 Support Phone Number:

+55 47 329-0065 Contact Email Address:

lumina@luminainfo.com.br Company URL:

http://www.luminainfo.com.br Primary Contact:

Odilon Mader Netto

1.58 Consultant ID: Brazil, Millennium Consultoria e Informática Ltda.

     758, R. Francisco Parolin
     Curitiba, NON United States 80.220-360
     Brazil

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55-41-333-5645 Company Fax:

+55-41-333-5645 Support Phone Number:

+55-41-913-2316 / +5 Contact Email Address:

millennium@millennium.etc.br Company URL:

http://www.millennium.etc.br Primary Contact:

André A. C. Santos

1.59 Consultant ID: Brazil, Onlinet Com. Serv. de Informática Ltda.

     R. Pará, 35 - Cid, Industrial
     Lorena, NON United States 12600-000
     Brazil

Consulting Specialties:

Internet Server Development Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55 (012) 553-1234 Company Fax:

+55 (012) 553-2568 Support Phone Number:

+55 (012) 553-1234 Contact Email Address:

suporte@onlinet.com.br Company URL:

http://www.onlinet.com.br Primary Contact:

Denys Sene

1.60 Consultant ID: Brazil, Pragana, Filhos E Cia Ltda

     CxPostal 721 CDC Aldeia
     Camaragibe, NON United States 54792-990
     Brazil

Consulting Specialties:

Custom Programming Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55-81-459-1776 Company Fax:

55-81-459-1776 Support Phone Number:

Not_Applicable Contact Email Address:

rpragana@acm.org Company URL:

http://Not_Applicable Primary Contact:

Rildo Pragana

1.61 Consultant ID: Brazil, Rodrigo Barbosa

     Phone or e-mail contact prefered
     Santa Rita do Sapucai, NON United States 37540-000
     Brazil

Consulting Specialties:

CIFS (Samba) Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55-35-9982-0006 Company Fax:

Not_Applicable Support Phone Number:

55-35-9982-0006 Contact Email Address:

rodrigob@linuxbr.com.br Company URL:

http://Not_Applicable Primary Contact:

Rodrigo Barbosa

1.62 Consultant ID: Brazil, TELEMATICA Coml. de Teleinformática Ltda.

     R. Princesa Isabel, 211
     Pelotas, NON United States 96015-590
     Brazil

Consulting Specialties:

CIFS (Samba) Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Conectiva Linux Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55 532 27-7266 Company Fax:

+55 532 27-7972 Support Phone Number:

+55 532 27-7266 Contact Email Address:

telematica@conesul.com.br Company URL:

http://Not_Applicable Primary Contact:

Vagner Farias

1.63 Consultant ID: Brazil, VisualBook

161, Hipolito da Costa
Rio de Janeiro, NON United States 20551040
Brazil

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55021-8720623 Company Fax:

55021-872-0576 Support Phone Number:

55021-2040333 Contact Email Address:

vb@visualbook.com.br Company URL:

http://www.visualbook.com.br Primary Contact:

Luiz Medeiros

1.64 Consultant ID: Brazil, aIPiWorks Network Consulting

R. Conceicao 233, Sala 1514, Ed. Centro Emp. ENCOL
Campinas, NON United States 13010-000
Brazil

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

RedHat / Mandrake / Specialty Distribution:

Other Other Specialty Distribution:

Mandrake / RedHat Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

055-19-2281876 Company Fax:

055-19-2281876 Support Phone Number:

055-019-2281876 Contact Email Address:

info@aipiworks.com.br Company URL:

http://www.aipiworks.com.br Primary Contact:

Mario R. Williams

1.65 Consultant ID: Brazil, iSNet Serviços de Informatica

     Av. Othon Gama d'Eça, 900/704
     Florianopolis, NON United States 88015-240
     Brazil

Consulting Specialties:

E-Commerce Main Distribution:

Other Other Distribution:

Conectiva Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+55 48 322-0249 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

info@isnet.com.br Company URL:

http://www.isnet.com.br Primary Contact:

Paulo Santos

1.66 Consultant ID: Bulgaria, University of Mining and Geology, Dept. Computer S

     Studentski Grad
     Sofia, NON United States 1700
     Bulgaria

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+359 2 62581 /ext.56 Company Fax:

+359 2 Support Phone Number:

+359 2 62581 /ext.56 Contact Email Address:

support@mgu.bg Company URL:

http://www.mgu.bg Primary Contact:

Volin Karagiozov

1.67 Consultant ID: Canada, 4 Office Automation

     375 Britannia Road East
     Mississauga, ON L4Z 3E2
     Canada

Consulting Specialties:

CIFS (Samba) Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

1-416-463-3423 Company Fax:

1-905-501-1515 Support Phone Number:

1-416-463-3423 Contact Email Address:

hyeung@4office.com Company URL:

http://www.4office.com Primary Contact:

Hillman Yeung

1.68 Consultant ID: Canada, 8D Technologies inc.

Pobox 778
Hudson, PQ J0P1H0
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

514.994.2645 Company Fax:

514.994.2654 Support Phone Number:

514.994.2645 Contact Email Address:

info@8D.com Company URL:

http://www.8D.com Primary Contact:

Patrick Bernard

1.69 Consultant ID: Canada, ADR Computing

     5-1115 West 10th Avenue
     Vancouver, BC V6H 1J2
     Canada

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 604 202 2800 Company Fax:

Not_Applicable Support Phone Number:

1 604 202 2800 Contact Email Address:

support@adrcomp.com Company URL:

http://www.adrcomp.com Primary Contact:

Chris Airriess

1.70 Consultant ID: Canada, Affinity Systems Inc.

Box 10, RR1
Innisfail, AB T4G 1T6
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-403-886-4638 Company Fax:

1-403-886-2663 Support Phone Number:

1-403-886-4638 Contact Email Address:

support@affinity-systems.ab.ca Company URL:

http://www.affinity-systems.ab.ca Primary Contact:

James Bourne

1.71 Consultant ID: Canada, Allt Computer Consulting

170 Willow Rd.
Guelph, ON N1H 1W8
Canada

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

519 763-0405 Company Fax:

508 448-0716 Support Phone Number:

n/a Contact Email Address:

allt@netloads.com Company URL:

http://www.netloads.com/allt Primary Contact:

Greg Allt

1.72 Consultant ID: Canada, Best Computers

10610 170 St
Edmonton, AB T5S-1P3
Canada

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

SuSE, Debian Specialty Distribution:

Other Other Specialty Distribution:

Secured Linux Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

780-413-9830 Company Fax:

780-413-9831 Support Phone Number:

780-413-9830 Contact Email Address:

support@best-comp.com Company URL:

http://http://www.best-comp.com/ Primary Contact:

Kurt Seifried

1.73 Consultant ID: Canada, CAL Consultants Inc.

     358 Danforth Ave., Suite 29
     Toronto, ON M4K 3Z2
     Canada

Consulting Specialties:

E-Commerce Internet Server Development Network Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Linux Certifications:

Not_Applicable Company Phone Number:

01.416.461.7634 Company Fax:

01.416.461.6908 Support Phone Number:

01.416.461.7634 Contact Email Address:

support@cal.ca Company URL:

http://www.cal.ca Primary Contact:

Graydon Clipperton

1.74 Consultant ID: Canada, Centurion Services

9351 Blundell Road
Richmond, BC V6Y 1K5
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not_Applicable Company Phone Number:

1-604-279-1857 Company Fax:

1-604-279-1800 Support Phone Number:

1-604-720-1857 Contact Email Address:

jwalter@rogers.wave.ca Company URL:

http://www.rogers.wave.ca Primary Contact:

Mr. Jan Walter

1.75 Consultant ID: Canada, Chris Newton

254 Barton
Fredericton, NB e3a 5a6
Canada

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

506-474-0454 Company Fax:

Not_Applicable Support Phone Number:

506-474-0454 Contact Email Address:

newton@unb.ca Company URL:

http://Not_Applicable Primary Contact:

Chris Newton

1.76 Consultant ID: Canada, Computing And Networking Centre

     POBOX 28117
     Dartmouth/Halifax, NS B2W6E2
     Canada

Consulting Specialties:

Network Administration Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

902-465-3734 Company Fax:

902-465-3734 Support Phone Number:

902-465-3734 Contact Email Address:

cnc98@sprint.ca Company URL:

http://Not_Applicable Primary Contact:

Ahmed Bentiba

1.77 Consultant ID: Canada, DFC International Computers Inc.

52 Mowatt Court
Thornhill, ON L3T 6V5
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Compaq IBM Linux Certifications:

RedHat Not Applicable Company Phone Number:

(905)731-6449 Company Fax:

(905)731-7684 Support Phone Number:

(905)731-6449 Contact Email Address:

support@dfc.com Company URL:

http://www.dfc.com Primary Contact:

David Fitzerman

1.78 Consultant ID: Canada, DFC International Computing Inc.

52 Mowatt Court
Thornhill, ON L3T 6V5
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

UltraPenguin Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

(905)731-6449 Company Fax:

(905)731-7684 Support Phone Number:

(905)731-6449 Contact Email Address:

support@dfc.com Company URL:

http://www.dfc.com Primary Contact:

David Fitzerman

1.79 Consultant ID: Canada, DPD Software Ltd

1455 Waverly St
Winnipeg, MB R2W 0RY
Canada

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slak Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

RedHat Company Phone Number:

204-489-4398 Company Fax:

204-489-4398 Support Phone Number:

204-489-4398 Contact Email Address:

kevind@dpdsoftware.mb.ca Company URL:

http://www.dpdsoftware.mb.ca Primary Contact:

Kevin Druet

1.80 Consultant ID: Canada, Diminishing Networks

5330 Entwhistle Dr
Nanaimo, BC V9V 1H2
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1(250) 758-2217 Company Fax:

1(250) 758-2217 Support Phone Number:

Not_Applicable Contact Email Address:

tech@mymail.penguinpowered.com Company URL:

http://www.soatusa.com/DimNet/ Primary Contact:

Adam Jordens

1.81 Consultant ID: Canada, G Desktop Solutions

     2037 Decarie Blvd #1
     Montreal, PQ H4A 3J2
     Canada

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(514) 996-5153 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

gldesk@gomer.mlink.net Company URL:

http://gomer.mlink.net/˜ gldesk/ Primary Contact:

Greg Patterson

1.82 Consultant ID: Canada, Gedris Linux Consulting

     183 St. Andrew Street
     Ottawa, ON K1N 5G3
     Canada

Consulting Specialties:

Custom Programming Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(613)562-0628 Company Fax:

Not_Applicable Support Phone Number:

(613)562-0628 Contact Email Address:

vic@worldchat.com Company URL:

http://www.worldchat.com Primary Contact:

Victor Gedris

1.83 Consultant ID: Canada, Global Proximity Corporation

RR 7
Woodstock, ON N4S 7W2
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+1 519 469 3439 Company Fax:

+1 519 469 8653 Support Phone Number:

+1 519 469 3439 Contact Email Address:

Support@Global.Proximity.ON.CA Company URL:

http://www.Global.Proximity.ON.CA Primary Contact:

Chris Tyler

1.84 Consultant ID: Canada, HK Computer Consultants

     35 Chalmers Drive
     Ajax, NON United States l155w8
     Canada

Consulting Specialties:

Custom Programming Custom Systems Network Administration System Administration Main Distribution:

Caldera Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

905-619-6888 Company Fax:

Not_Applicable Support Phone Number:

905-619-6888 Contact Email Address:

hallamh@excite.com Company URL:

http://Not_Applicable Primary Contact:

Hallam Holder

1.85 Consultant ID: Canada, Hard Data Ltd.

     11060 - 166 Avenue
     Edmonton, AB T5X 1Y3
     Canada

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

ALL Specialty Distribution:

Other Other Specialty Distribution:

Hard Hat Linux Reseller Authorizations:

Compaq HP Linux Certifications:

Not Applicable Company Phone Number:

01-780-456-9771 Company Fax:

01-780-456-9772 Support Phone Number:

01-780-456-1510 Contact Email Address:

support@harddata.com Company URL:

http://www.harddata.com Primary Contact:

Maurice Hilarius

1.86 Consultant ID: Canada, ITPS Groupe Inc.

     350 Price Arthur Apt.810
     Montreal, PQ H2X 3R4
     Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

RedHat Company Phone Number:

(514) 286-1354 Company Fax:

(514) 286-4827 Support Phone Number:

(514)-821-1568 Contact Email Address:

Boris.Kuschel@itpsgroup.com Company URL:

http://http://www.itpsgroup.com/ Primary Contact:

Boris Kuschel

1.87 Consultant ID: Canada, JDP Computer Systems Inc.

206 Finch Ave. East
Toronto, ON L9W 3S9
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-416-733-3655 Company Fax:

1-416-733-1839 Support Phone Number:

1-416-733-3655 Contact Email Address:

support@jdp.com Company URL:

http://www.jdp.com Primary Contact:

Gordon Gray

1.88 Consultant ID: Canada, Jeff Tranter

     1 Laurie Court
     Kanata, ON K2L 1S2
     Canada

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 613 836 7697 Company Fax:

Not_Applicable Support Phone Number:

+1 613 836 7697 Contact Email Address:

tranter@pobox.com Company URL:

http://www.pobox.com/˜ tranter Primary Contact:

Jeff Tranter

1.89 Consultant ID: Canada, LeadSource Consultants Inc.

1227 Mississauga Road
Mississauga, ON L5H 2J1
Canada

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

HoNet Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

905 278 6808 Company Fax:

905 278 6863 Support Phone Number:

905 278 6808 Contact Email Address:

support@leadsource.ca Company URL:

http://http://www.leadsource.ca Primary Contact:

Adam Mikolajewicz

1.90 Consultant ID: Canada, Logisoft Technologies inc

     718 pl Denise-pelletier
     Ste-Julie, PQ J3E 2G8
     Canada

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Mandrake Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(514) 570-7638 Company Fax:

Not_Applicable Support Phone Number:

(514) 580-0050 Contact Email Address:

support@logisoftech.com Company URL:

http://www.logisoftech.com Primary Contact:

Guillaume Bourque

1.91 Consultant ID: Canada, MOE Online Enterprises

50 Stephanie, Suite 1008
Toronto, ON M5T1B3
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 416 588 0672 Company Fax:

Not_Applicable Support Phone Number:

n/a Contact Email Address:

mozai@canada.com Company URL:

http://ativan.netdesign.net/˜ moses/ Primary Contact:

Moses Moore

1.92 Consultant ID: Canada, Macadamian Technologies Inc.

     700 Industrial Avenue Suite 220
     Ottawa, ON K1G 0Y9
     Canada

Consulting Specialties:

Custom Programming Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-877-7-SYNDEO Company Fax:

1-613-739-9859 Support Phone Number:

1-877-7-SYNDEO Contact Email Address:

info@macadamian.com Company URL:

http://www.macadamian.com Primary Contact:

Claude Montpetit

1.93 Consultant ID: Canada, Mediabase

     3 shenstone rd, (changing soon)
     North York, ON m2r 3b3 (changi
     Canada

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware 4.0 Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

4164685400 Company Fax:

na Support Phone Number:

Not_Applicable Contact Email Address:

mediabase@ihosts.net Company URL:

http://www.mediabase.ihosts.net Primary Contact:

Cyrus Hill

1.94 Consultant ID: Canada, Minek Consulting Inc.

     B403-1331 Homer St.
     Vancouver, BC V6B 5M5
     Canada

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-604-3185875 Company Fax:

1-604-899-2492 Support Phone Number:

1-604-3185875 Contact Email Address:

support@minek.com Company URL:

http://www.minek.com Primary Contact:

Dan Rokosz

1.95 Consultant ID: Canada, Montage IT Services

350, 708 - 11 Ave. SW
Calgary, AB T2R0E4
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

All Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

01.403.265.4461 Company Fax:

01.403.265.4462 Support Phone Number:

Not_Applicable Contact Email Address:

dboyd@cal.montage.ca Company URL:

http://www.montage.ca Primary Contact:

Darren Boyd

1.96 Consultant ID: Canada, MostlyLinux

     Not_Applicable
     Calgary, AB Not_Applicable
     Canada

Consulting Specialties:

CIFS (Samba) Custom Systems Network Administration System Administration System Security Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Small Business serve Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

403-285-1399 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Scott Barker

1.97 Consultant ID: Canada, Net Direct Inc.

557 Havelock Dr.
Waterloo, ON N2L 4Z1
Canada

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

519-579-5006 Company Fax:

519-745-9940 Support Phone Number:

519-579-5006 Contact Email Address:

support@netdirect.ca Company URL:

http://www.net.direct.ca Primary Contact:

John Van Ostrand

1.98 Consultant ID: Canada, Pilkington Software Inc.

     1046 Barrington St
     Halifax, NS B3H 2R1
     Canada

Consulting Specialties:

Custom Programming Custom Systems Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

902 423 6033 Company Fax:

902 423 6033 Support Phone Number:

902 423 6033 Contact Email Address:

support@cpsoft.com Company URL:

http://www.cpsoft.com Primary Contact:

Charles Pilkington

1.99 Consultant ID: Canada, Pryor and Pryor Inc.

     602 - 1230 Comox Street
     Vancouver, BC V6E 1K7
     Canada

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

604-685-2621 Company Fax:

604-683-3488 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://http://www.pryor-and-pryor.com Primary Contact:

Roger Pryor

1.100 Consultant ID: Canada, Quist Consulting

219 Donlea Drive
Toronto, ON M4G2N1
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Debian Specialty Distribution:

LinuxRouter Other Specialty Distribution:

UltraPenguin Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1.416.696.7600 Company Fax:

1.416.978.6620 Support Phone Number:

1.416.696.7600 Contact Email Address:

russ@quist.on.ca Company URL:

http://www.quist.on.ca Primary Contact:

Russell Suherland

1.101 Consultant ID: Canada, Real-Time Remedies Inc.

     33 Ridgefield Crescent
     Nepean, ON K2H 6S3
     Canada

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Intel x86 and NetWin Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

613-277-4944 Company Fax:

616-828-9659 Support Phone Number:

Not_Applicable Contact Email Address:

mlord@pobox.com Company URL:

http://Not_Applicable Primary Contact:

Mark Lord

1.102 Consultant ID: Canada, Roaring Penguin Software Inc.

     986 Eiffel Avenue
     Ottawa, ON K2C 0J2
     Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Network Administration System Administration

Main Distribution:

Other Other Distribution:

Any distro Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

613 851-4379 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

David Skoll

1.103 Consultant ID: Canada, Salmar Consulting Inc.

P.O. Box 52531, 1801 Lakeshore Road West
Mississauga, ON L5J 4S6
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

All Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

905-403-8835 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

info@salmar.com Company URL:

http://www.salmar.com Primary Contact:

Marcel Gagne

1.104 Consultant ID: Canada, Soft Touch Computers

Not_Applicable
Calgary, AB Not_Applicable
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

SUSE Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not Applicable Company Phone Number:

(403)259-6016 Company Fax:

(403)258-0267 Support Phone Number:

(403)259-6016 Contact Email Address:

admin@softouchcomputer.com Company URL:

http://www.softouchcomputer.com Primary Contact:

Al Green

1.105 Consultant ID: Canada, Stable Network Technologies Inc.

     141 Adelaide Street West, Suite 1200
     Toronto, ON M5H 3L5
     Canada

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 416 367 2745 Company Fax:

1 416 861 0650 Support Phone Number:

Not_Applicable Contact Email Address:

support@snti.com Company URL:

http://www.snti.com Primary Contact:

David Kerry

1.106 Consultant ID: Canada, Steve Willer

     1544A Queen St. W.
     Toronto, ON M6R1A6
     Canada

Consulting Specialties:

Custom Programming E-Commerce Internet Server Development Internet (Web) Programming Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

416-516-8470 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.interlog.com/˜ willer/ Primary Contact:

Steve Willer

1.107 Consultant ID: Canada, Technical Consulting Group

     25 - 52061 - RR215
     Sherwood Park, AB T8E-1B2
     Canada

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Linux on Alpha Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

780-922-5755 Company Fax:

780-922-6990 Support Phone Number:

Not_Applicable Contact Email Address:

ken.gehring@gov.ab.ca Company URL:

http://Not_Applicable Primary Contact:

Ken Gehring

1.108 Consultant ID: Canada, TeleDynamics Communications Inc

7 Forest Place RR#1
Sauble Beach, ON N0H 2G0
Canada

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

519-422-1150 Company Fax:

519-422-2723 Support Phone Number:

Not Available Contact Email Address:

teledynamics@canada.com Company URL:

http://http://www.teledyn.com Primary Contact:

Gary Murphy

1.109 Consultant ID: Canada, The Net Result

#10 3450 Coast Meridian Rd
Port Coquitlam, BC V3B7H2
Canada

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

604 942 7257 Company Fax:

604 942 7440 Support Phone Number:

Not_Applicable Contact Email Address:

support@result.com Company URL:

http://Not_Applicable Primary Contact:

Jay Thorne

1.110 Consultant ID: Canada, The People’s Linux

109-288 East 14th Avenue
Vancouver, BC V5T 2M6
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

604-709-9512 Company Fax:

604-709-9512 Support Phone Number:

604-709-9512 Contact Email Address:

fred@nuge.com Company URL:

http://http://www.linuxnut.com Primary Contact:

Bob Strasser

1.111 Consultant ID: Canada, Troika Management Services

Unit 4 - 40 Fairfax Court
London, NON United States N6G3Y3
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-519-670-4567 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

xbill@xbill.com Company URL:

http://www.xbill.com Primary Contact:

William McLean

1.112 Consultant ID: Canada, Troy Ryder

3665 Luxmoore Road
Kelowna, BC V1W 4C7
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

250-764-0917 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

ryder@direct.ca Company URL:

http://www.vanisle.net/ryder Primary Contact:

Troy Ryder Not_Applicable

1.113 Consultant ID: Canada, VL

     221 BRitannia
     Ottawa, ON K2B - 5X1
     Canada

Consulting Specialties:

System Security Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

1 (613) 292-9818 Company Fax:

1 (613) 829-0967 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Hung Vu

1.114 Consultant ID: Canada, Webcon, Inc.

70 Forest Hill Ave.
Ottawa, ON K2C 1P6
Canada

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 613 276 6206 Company Fax:

+1 613 276 8206 Support Phone Number:

+1 613 276 6206 Contact Email Address:

tech@webcon.net Company URL:

http://www.webcon.net Primary Contact:

Robert Hardy

1.115 Consultant ID: Canada, iNsu Innovations

     3465 Thimmens
     St-Laurent, PQ H4R 1V5
     Canada

Consulting Specialties:

E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

514-336-5544 Company Fax:

514-336-8128 Support Phone Number:

514-336-5544 Contact Email Address:

support@insu.com Company URL:

http://www.insu.com Primary Contact:

Joel Pomerleau

1.116 Consultant ID: Colombia, Linux Colombia Ltda

     Avenida 5AN ·23DN68 Oficina 2-118B
     Cali, NON United States Not_Applicable
     Colombia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

57 2 6607908 Company Fax:

57 2 6676596 Support Phone Number:

57 2 6607908 Contact Email Address:

linuxcol@linuxcolombia.com.co Company URL:

http://www.linuxcolombia.com.co Primary Contact:

Juan Diego Bolaños Ramirez

1.117 Consultant ID: Colombia, Msoft

     Cra 24 No 80-52 Apto G-201
     Bucaramanga, NON United States 212
     Colombia

Consulting Specialties:

CIFS (Samba) Internet Server Development Netware Connectivity Network Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

++57-7-6384935 Company Fax:

++57-7-6389374 Support Phone Number:

++57-7-6313428 Contact Email Address:

nelsonmauriciosilva@yahoo.com Company URL:

http://Not_Applicable Primary Contact:

Mauricio Silva

1.118 Consultant ID: Colombia, Skina Ltda.

     Calle 95 #30-61 int 8
     Santafe de Bogota, NON United States Not_Applicable
     Colombia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

5712565008 Company Fax:

5712565008 Support Phone Number:

cel: 2059620 Contact Email Address:

skina@skina.com.co Company URL:

http://www.skina.com.co Primary Contact:

Ricardo Naranjo_Faccini

1.119 Consultant ID: Czech Republic, Martin Knotek, poskytovani software

     Havelkova 23
     Brno, NON United States 625 00
     Czech Republic

Consulting Specialties:

Custom Programming Custom Systems System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

420 5 356605 Company Fax:

Not_Applicable Support Phone Number:

420 5 356605 Contact Email Address:

knotekmartin@email.cz Company URL:

http://www.email.cz Primary Contact:

Martin Knotek

1.120 Consultant ID: Denmark, Arnecke Informatik

     Havnegade 4
     Odense, NON United States 5100
     Denmark

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+45 63130050 Company Fax:

+45 63130051 Support Phone Number:

+45 63130050 Contact Email Address:

info@a-info.dk Company URL:

http://www.a-info.dk Primary Contact:

Allan Hansen

1.121 Consultant ID: Denmark, Dansk Data Elektronik A S

     Herlev Hovedgade 199
     Herlev, NON United States 2730
     Denmark

Consulting Specialties:

Custom Systems Main Distribution:

Other Other Distribution:

Any well known Linux Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+45 44572000 Company Fax:

+45 44572001 Support Phone Number:

+45 44572010 Contact Email Address:

support@dde.dk Company URL:

http://www.dde.dk Primary Contact:

Jens Knudsen

1.122 Consultant ID: Denmark, Plomus

Nøddelunden 110
Smørum, NON United States DK-2765
Denmark

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+45 44 66 24 16 Company Fax:

+45 44 66 25 16 Support Phone Number:

+45 44 66 25 16 Contact Email Address:

support@plomus.dk Company URL:

http://www.plomus.dk Primary Contact:

Claus Sørensen

1.123 Consultant ID: Dominican Republic, microbell

jesus maestr 18, suite A3, mirador sur
santo domingo, ZZ 0000
Dominican Republic

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

809 383 2613 Company Fax:

809 540 4025 Support Phone Number:

809 383 2613 Contact Email Address:

microbell@microbell.com.do Company URL:

http://microbell.com.do Primary Contact:

homero gonzalez

1.124 Consultant ID: Egypt, AB2

     29 Alexander the Great St.
     Alexandria, NON United States 21131
     Egypt

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

Intel Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+20 3 482 4756 Company Fax:

+20 3 483 5030 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://www.ab2.com Primary Contact:

Khalid Baheyeldin

1.125 Consultant ID: Finland, Netsol Network Solutions Oy

     Hiihtotie 3 A
     Vantaa, NON United States 01280
     Finland

Consulting Specialties:

Custom Systems Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+358 9 34090410 Company Fax:

+358 9 34090404 Support Phone Number:

+358 9 34090410 Contact Email Address:

support@netsol.fi Company URL:

http://www.netsol.fi Primary Contact:

Timo Virtaneva

1.126 Consultant ID: Finland, SOT Finnish Sofware Engineering Ltd.

Hermiankatu 8 E
TAMPERE, NON United States 33720
Finland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Best Linux Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+358-3-316 5544 Company Fax:

+358-3-316 5959 Support Phone Number:

0600-95549 Contact Email Address:

linux@bestlinux.net Company URL:

http://www.sot.com Primary Contact:

Santeri Kannisto Kai Valijärvi

1.127 Consultant ID: Finland, Tmi Netics Konsultointi

     Eteläniityntie 6
     MUSTASAARI, NON United States FIN-65610
     Finland

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

planned with custome Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+358 50 501 2132 Company Fax:

Not_Applicable Support Phone Number:

+358 50 501 2132 Contact Email Address:

juhani.puska@netics.fi Company URL:

http://www.netics.fi Primary Contact:

Juhani Puska

1.128 Consultant ID: France, ALCOVE

     12 - 13 place Indira Gandhi
     Gennevilliers, NON United States 92230
     France

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 1 47 33 82 84 Company Fax:

+33 1 47 33 76 98 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Anne Collet

1.129 Consultant ID: France, Alcove

     12-13 place Indira Gandhi
     Gennevilliers, NON United States 92230
     France

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

33147338284 Company Fax:

33147337698 Support Phone Number:

Not_Applicable Contact Email Address:

assistance@alcove.fr Company URL:

http://www.alcove.fr Primary Contact:

Anne Collet

1.130 Consultant ID: France, CRAO

     19, rue Forcrand
     Montpellier, NON United States 34000
     France

Consulting Specialties:

Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 4 67 63 24 27 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

crao@crao.net Company URL:

http://crao.net Primary Contact:

Arnaud Fontaine

1.131 Consultant ID: France, EBC Consulting

16, rue des sorbiers
SORBEY, NON United States 57580
France

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

client specific Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 (0) 3. 87.64.58. Company Fax:

+33 (0) 3. 87.64.58. Support Phone Number:

+33 (0) 3. 87.64.58. Contact Email Address:

ebconsul@altavista.net Company URL:

http://www.ebcip.fr Primary Contact:

Eric Binger

1.132 Consultant ID: France, Hubert Gregoire

118, avenue Gambetta
PARIS, NON United States 75020
France

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 (0) 140312326 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

hubert@dune-concept.com Company URL:

http://dune-concept.com Primary Contact:

Hubert Gregoire

1.133 Consultant ID: France, NTCD

     11 rue A. Carrel
     PARIS, ZZ 75019
     France

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Debian Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

33 1 40 40 77 09 Company Fax:

33 1 40 40 77 09 Support Phone Number:

none Contact Email Address:

http://www.ntcd.com/hotline/index.htm Company URL:

http://ntcd Primary Contact:

Didier Blanchard

1.134 Consultant ID: France, Paralline

     71, avenue des Vosges
     STRASBOURG, NON United States 67000
     France

Consulting Specialties:

Custom Systems Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

33 3 88 14 17 40 Company Fax:

33 3 88 14 17 41 Support Phone Number:

33 3 88 14 17 40 Contact Email Address:

support@paralline.com Company URL:

http://www.paralline.com Primary Contact:

Pierre BRUA

1.135 Consultant ID: France, Paul Boyer Consultants Sarl

     20 rue d'Hauteville
     Paris, NON United States 75010
     France

Consulting Specialties:

Firewalls and Internet Security System Security Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 1 47 700 801 Company Fax:

+33 1 47 700 850 Support Phone Number:

+33 1 47 700 801 Contact Email Address:

paulboyer@usa.net Company URL:

http://www.usa.net Primary Contact:

Paul Boyer

1.136 Consultant ID: France, Thorvax

     39, rue Armand Silvestre
     Courbevoie, NON United States 92400
     France

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33 6 80 08 93 83 Company Fax:

+33 1 43 33 51 33 Support Phone Number:

+33 6 80 08 93 83 Contact Email Address:

Illaire@Club-Internet.fr Company URL:

http://www.penguinpowered.com/˜ tyrann Primary Contact:

Vincent Illaire

1.137 Consultant ID: France, sokram inc

     161 rue de la eoquette
     Paris, NON United States 75011
     France

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+33614704591 Company Fax:

+33140242235 Support Phone Number:

+33614704591 Contact Email Address:

sokram@usa.net Company URL:

http://Not_Applicable Primary Contact:

Bernard Not_Applicable

1.138 Consultant ID: Germany, Andreas Spengler EDV-Service

     Gronauer Strasse 9
     Karben, NON United States 61184
     Germany

Consulting Specialties:

CIFS (Samba) Custom Systems Network Administration System Administration Main Distribution:

Suse Other Distribution:

RedHat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 6039 44401 Company Fax:

+49 6039 44401 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Andreas Spengler

1.139 Consultant ID: Germany, Bigge & Langanke GbR

     Corunnastr. 1
     Iserlohn, NON United States 58636
     Germany

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

support@transdata.de Company URL:

http://www.transdata.de Primary Contact:

Frank Bigge

1.140 Consultant ID: Germany, CHIRON

Neupullach 3
Hohenlinden, NON United States 85664
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Selfmade embedded Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 8124 52 83 44 Company Fax:

+49 8124 52 83 46 Support Phone Number:

Not_Applicable Contact Email Address:

chiron@chiron.de Company URL:

http://www.chiron.de Primary Contact:

M. Brandstetter

1.141 Consultant ID: Germany, Call-a-Server

     Not_Applicable
     Berlin, NON United States D-14129 Berlin
     Germany

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

DLD, RedHat und S.u. Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 +177 7170896 Company Fax:

+49 +30 8017423 Support Phone Number:

+49 +177 7170896 Contact Email Address:

support@call-a-server.de Company URL:

http://www.call-a-server.de Primary Contact:

Joachim von Thadden

1.142 Consultant ID: Germany, Gecko GmbH

     Ringstr. 53
     Bad Oeynhausen, NON United States 32549
     Germany

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49.5731.741010 Company Fax:

+49.5731.741013 Support Phone Number:

Not Available Contact Email Address:

support@lizard.de Company URL:

http://www.lizard.de Primary Contact:

Eike Sieker

1.143 Consultant ID: Germany, Gerd Aschemann

     Osannstr. 49
     Darmstadt, NON United States 64285
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

Redhat, Caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49/6151/16-2259 Company Fax:

+49/6151/16-3052 Support Phone Number:

+49/6151/16-2259 Contact Email Address:

Aschemann@Informatik.TU-Darmstadt.de Company URL:

http://www.Informatik.TU-Darmstadt.de Primary Contact:

Gerd Aschemann

1.144 Consultant ID: Germany, Grosch & Link GmbH

     Friedrich-Ebert-Anlage 18
     Frankfurt, NON United States 60325
     Germany

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

RedHat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49699740130 Company Fax:

+496997401322 Support Phone Number:

Not_Applicable Contact Email Address:

frankfurt@grosch-link.de Company URL:

http://www.grosch-link.de/html Primary Contact:

Mario Oschwald

1.145 Consultant ID: Germany, ID-Pro GmbH

Koenigswinterer Str. 116
Bonn, NON United States 53227
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

All major distributi Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-(0)228-42154-0 Company Fax:

+49-(0)228-42154-29 Support Phone Number:

+49-(0)228-42154-19 Contact Email Address:

support@id-pro.de Company URL:

http://id-pro.de Primary Contact:

Daniel Riek

1.146 Consultant ID: Germany, Jochen Laser IT-Services

     Fliederweg 2a
     Woebbelin, NON United States 19288
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

others as well Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 (0)38753 80181 Company Fax:

on request Support Phone Number:

+49 (0)38753 80181 Contact Email Address:

on request Company URL:

http://Not_Applicable Primary Contact:

Jochen Laser

1.147 Consultant ID: Germany, LINUXHAUS

     RINGBAHNSTR. 13
     BERLIN, NON United States 10711
     Germany

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

SuSE Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 30 890 944 63 Company Fax:

+49 30 890 944 64 Support Phone Number:

Not Available Contact Email Address:

support@linuxhaus.de Company URL:

http://www.linuxhaus.de Primary Contact:

Joerg Fruehbrodt

1.148 Consultant ID: Germany, Linux-Systemhaus Schulz

     Hauptfeld 18
     Dortmund, NON United States 44369
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

none Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

0049 172 2362995 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

schulz@linux-systemhaus.de Company URL:

http://www.Linux-Systemhaus.de Primary Contact:

Karsten Schulz

1.149 Consultant ID: Germany, Manfred Kunde EDV-Beratung

     Hauptstrasse 70
     Karben, NON United States 61184
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Internet (Web) Programming

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-(0)-06039-931064 Company Fax:

+49-(0)-06039-931065 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://http://www.Manfred-Kunde.de Primary Contact:

Manfred Kunde

1.150 Consultant ID: Germany, Matthias Stolte Datentechnik

     Not_Applicable
     33613 Bielefeld, NON United States Not_Applicable
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

SuSE und andere... Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-521-880056 Company Fax:

+49-521-880059 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.151 Consultant ID: Germany, Michael Bussmann EDV-Dienstleistungen

     Im Brook 8
     Haltern/Westf., NON United States 45721
     Germany

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 2364 108537 Company Fax:

+49 2364 4095 Support Phone Number:

n/a Contact Email Address:

michael.bussmann@gmx.net Company URL:

http://Not_Applicable Primary Contact:

Michael Bussmann

1.152 Consultant ID: Germany, Michael Krause Software-Entwicklung

     Auf der Horst 11
     Borgholzhausen, NON United States 33829
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 5425 930310 Company Fax:

+49 5425 930311 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://homepages.teuto.net/mkrause/ Primary Contact:

Michael Krause

1.153 Consultant ID: Germany, NetUSE GmbH

Siemenswall
Kiel, NON United States 24107
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 431 386 435 00 Company Fax:

+49 431 386 435 99 Support Phone Number:

Not_Applicable Contact Email Address:

service@netuse.de Company URL:

http://netuse.de Primary Contact:

Heinz Rohde

1.154 Consultant ID: Germany, PiN - Präsenz im Netz Ges. f. Informationstechnolo

Leimbacher Str. 36
Wuppertal, NON United States 42281
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Alle Distributionen Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-202-2501164 Company Fax:

+49-202-2501165 Support Phone Number:

+49-202-2501164 Contact Email Address:

support@pinserve.de Company URL:

http://www.pinserve.de Primary Contact:

Andre Dressler

1.155 Consultant ID: Germany, SDC GmbH

Schlossstrase 26
Berlin, NON United States 13507
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Redhat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 30 4309430 Company Fax:

+49 30 43094310 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.sdc.de Primary Contact:

Christian Gleich

1.156 Consultant ID: Germany, Simulina GmbH

     Trajanstraße  8
     LADENBURG, NON United States DE-68526
     Germany

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+49 6203 92 20 41 Company Fax:

+49 6203 92 20 42 Support Phone Number:

+49 6203 92 40 41 Contact Email Address:

hk@simulina.se Company URL:

http://www.ncc-mannheim.net/simulina Primary Contact:

Håkan Källberg

1.157 Consultant ID: Germany, Syslab.com oHg

Karl-Theodor Straße 66
München, NON United States 80803
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 89 306358-90 Company Fax:

+49 89 306358-99 Support Phone Number:

+49 89 306358-92 Contact Email Address:

support@syslab.com Company URL:

http://www.syslab.com Primary Contact:

Manfred Lang

1.158 Consultant ID: Germany, TDS Consulting GmbH

     Gustav-Heinemann-Ring 212
     Munich, NON United States 81739
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not Applicable Company Phone Number:

+49 89 62736-0 Company Fax:

+49 89 62736-129 Support Phone Number:

Not_Applicable Contact Email Address:

Peter.Schroff@tds.de Company URL:

http://www.tds.de Primary Contact:

Peter Schroff

1.159 Consultant ID: Germany, WebArtists Internet Services

Steendammwisch 49
Hamburg, NON United States 22459
Germany

Consulting Specialties:

Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-40-555 40 427 Company Fax:

+49-40-555 40 182 Support Phone Number:

0172/4032506 Contact Email Address:

support@linuxberatung.com Company URL:

http://www.webartists.net Primary Contact:

Markus Braasch

1.160 Consultant ID: Germany, Wesemann Software & Consulting

     Heinrich-Plett-Allee 18
     Bremen, NON United States D-28259
     Germany

Consulting Specialties:

System Administration Main Distribution:

Suse Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

++49 421 579 579 3 Company Fax:

++49 421 579 58 28 Support Phone Number:

+49 421 579 579 3 Contact Email Address:

info@wscnet.de Company URL:

http://www.wscnet.de Primary Contact:

Andreas Wesemann

1.161 Consultant ID: Germany, Wolfgang Kirch

     Gustav-Heinemann-Ring 51
     Eisenberg, NON United States 67304
     Germany

Consulting Specialties:

Firewalls and Internet Security Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

49 6351 398761 Company Fax:

49 6351 398762 Support Phone Number:

Not Available Contact Email Address:

Kirch.Wolfgang@t-online.de Company URL:

http://www.t-online.de Primary Contact:

Wolfgang Kirch

1.162 Consultant ID: Germany, WorNet Internetdienste

Bürgermeister-Graf-Ring 28
Geretsried, NON United States D-82538
Germany

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 8171 41809-0 Company Fax:

+49 8171 41809-9 Support Phone Number:

Not Available Contact Email Address:

wornet@wor.net Company URL:

http://www.wor.net Primary Contact:

Christian Eich

1.163 Consultant ID: Germany, ariadne_net

Köpenicker Str. 8
Berlin, NON United States 10997
Germany

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

0493061280438 Company Fax:

0493061280437 Support Phone Number:

0493061280438 Contact Email Address:

aduecker@t-online.de Company URL:

http://www.t-online.de Primary Contact:

Andreas Dücker

1.164 Consultant ID: Germany, b.i.t. beratungsgesellschaft für informationstech

     elisabethenstr. 62
     darmstadt, NON United States 64283
     Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-6151-827575 Company Fax:

+49-6151-827576 Support Phone Number:

+49-6151-827575 Contact Email Address:

support@b-i-t.de Company URL:

http://www.b-i-t.de Primary Contact:

Andy Schmidt

1.165 Consultant ID: Germany, conIT GmbH

Industriering 7
Grosswallstadt, NON United States 63868
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 (0)6022/262-082 Company Fax:

+49 (0)6022/262-101 Support Phone Number:

+49 (0)6022/262-102 Contact Email Address:

support@conIT.com Company URL:

http://www.conIT.com Primary Contact:

Arnold Weis

1.166 Consultant ID: Germany, dev consulting GmbH

Webereistr. 3
Bielefeld, NON United States 33602
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Linux Server Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Not Applicable Company Phone Number:

Not_Applicable Company Fax:

+49 521 1365803 Support Phone Number:

+49 521 1365800 Contact Email Address:

info@devcon.net Company URL:

http://www.devcon.net Primary Contact:

Winfried Motzkus

1.167 Consultant ID: Germany, dynamis EDV-Consulting

Bundschuhstrasse 24
76661 Philippsburg, NON United States 76661
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 7256 9326-0 Company Fax:

+49 7256 9326-90 Support Phone Number:

+49 7256 9326-0 Contact Email Address:

support@dynamis.de Company URL:

http://www.dynamis.de Primary Contact:

Jochen Lillich

1.168 Consultant ID: Germany, envi.con KG

Brandenburgische Str. 69
Berlin, NON United States 10713
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

almost all Specialty Distribution:

Other Other Specialty Distribution:

almost all Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-30-34902670 Company Fax:

+49-30-34902671 Support Phone Number:

+49-30-34902670 Contact Email Address:

yourx@envicon.de Company URL:

http://http://www.envicon.de Primary Contact:

Stefania Rozzi

1.169 Consultant ID: Germany, frontsite AG

     Eschollbrücker Straße 4a
     Darmstadt, NON United States 64283
     Germany

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

high avail. and ISDN Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-6151-300-88-80 Company Fax:

+49-6151-300-88-80 Support Phone Number:

+49-6151-300-88-80 Contact Email Address:

info@frontsite.de Company URL:

http://www.frontsite.de Primary Contact:

Mark Semmler

1.170 Consultant ID: Germany, intraDAT GmbH

     Wilhelm-Leuschner-Str. 9-11
     Frankfurt/Main, NON United States D-60329
     Germany

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet (Web) Programming Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 +69-9740-5703 Company Fax:

+49 +69-9740-5705 Support Phone Number:

+49 +69-9740-5703 Contact Email Address:

email@intradat.com Company URL:

http://www.intradat.com Primary Contact:

Manfred Schmid

1.171 Consultant ID: Germany, ipm intranet project management gmbh

     Huelchrather Strasse 17-23
     Koeln, NON United States 50670
     Germany

Consulting Specialties:

E-Commerce Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

49-221-97353217 Company Fax:

49-221-97353216 Support Phone Number:

49-221-97353217 Contact Email Address:

support@ipm-koeln.de Company URL:

http://http://www.ipm-koeln.de Primary Contact:

Dietmar Horn

1.172 Consultant ID: Germany, loth systemtechnik

Julius-Vosseler-Strasse 136
Hamburg, ZZ D-22527
Germany

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Not_Applicable Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 40 40192181 Company Fax:

+49 40 40192183 Support Phone Number:

+49 171 5406784 Contact Email Address:

support@loth.de Company URL:

http://www.loth.de Primary Contact:

Michael Loth

1.173 Consultant ID: Germany, netServe oHG Münster

Fresnostraße 10a
Münster, NON United States 48159
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

49-251-98725-0 Company Fax:

49-251-866977 Support Phone Number:

49-251-98725-21 Contact Email Address:

info@muenster.com Company URL:

http://www.muenster.com Primary Contact:

Torsten Mueller

1.174 Consultant ID: Germany, ralf geschke internet consulting

     Roemerhofweg 1a
     Erftstadt, NON United States 50374
     Germany

Consulting Specialties:

Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49-177-5262345 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

support@geschke.nt Company URL:

http://www.geschke.nt Primary Contact:

Ralf Geschke

1.175 Consultant ID: Germany, regioconnect GmbH

Bahnhofstr. 11
Steinfurt, NON United States Steinfurt
Germany

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+492551919300 Company Fax:

+492551833169 Support Phone Number:

Not_Applicable Contact Email Address:

info@regioconnect.net Company URL:

http://www.regioconnect.net Primary Contact:

Michael Rüve

1.176 Consultant ID: Germany, unitegs - Unix Systems Technologies

     Wacholderweg 11
     Ilmenau, NON United States 98693
     Germany

Consulting Specialties:

Custom Programming Custom Systems System Administration Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+49 3677 884620 Company Fax:

+49 3677 884621 Support Phone Number:

Not Available Contact Email Address:

support@unitegs.de Company URL:

http://www.unitegs.de Primary Contact:

Gunar Schorcht

1.177 Consultant ID: Hong Kong, Cyber Channel Int’l Ltd

Room 1723 610 Nathan Road Kowloon
Hong Kong SAR, NON United States Not_Applicable
Hong Kong

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not Applicable Company Phone Number:

852 23754341 Company Fax:

852 23754148 Support Phone Number:

852 27823211 Contact Email Address:

support@cyberchn.com.hk Company URL:

http://http://www.cyberchn.com.hk Primary Contact:

Thomas Lee Paul Lee

1.178 Consultant ID: Hong Kong, Genesis Systems Int’l Ltd.

     5/F., 7/F., Honytex Building T.S.T, Kowloon
     Hong Kong, ZZ Not_Applicable
     Hong Kong

Consulting Specialties:

Internet Server Development Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+852 2815 0728 Company Fax:

+852 2815 0729 Support Phone Number:

+852 2199 7255 Contact Email Address:

support@genesis.com.hk Company URL:

http://www.genesis.com.hk Primary Contact:

Gyver Lo

1.179 Consultant ID: Hong Kong, Hong Kong Terminal

72A G/F Tokwawan Road,
Kowloon, NON United States Not_Applicable
Hong Kong

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

852-21420007 Company Fax:

852-27814919 Support Phone Number:

852-21420007 Contact Email Address:

support@zero.com.hk Company URL:

http://http://zero.com.hk Primary Contact:

Peter Wong

1.180 Consultant ID: Hong Kong, Linux Center (Hong Kong) Ltd.

Unit A, 5th Floor, Comet Comeercial Building , 42A
Kowloon, NON United States Not_Applicable
Hong Kong

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+852 2959 3675 Company Fax:

+852 2959 3672 Support Phone Number:

available to custome Contact Email Address:

available to customers only Company URL:

http://Not_Applicable Primary Contact:

Chris Yung

1.181 Consultant ID: Hong Kong, Uniforce System Ltd.

     Rm 903B, Kaiser Center, 18 Centre Street, Sai Ying
     Hong Kong, NON United States Not_Applicable
     Hong Kong

Consulting Specialties:

Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+852 24464316 Company Fax:

+852 21918185 Support Phone Number:

+852 24464316 Contact Email Address:

support@uniforce.net Company URL:

http://www.uniforce.net Primary Contact:

Francis Kam

1.182 Consultant ID: Hong Kong, Vortex Ltd.

Tsim Sha Tsui
Not_Applicable, NON United States Not_Applicable
Hong Kong

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

(852) 23021988 Company Fax:

(852) 23021212 Support Phone Number:

Not_Applicable Contact Email Address:

scwong@ieee.org Company URL:

http://Comming Soon Primary Contact:

Stephen Wong

1.183 Consultant ID: Hong Kong, Yoric Network Ltd.

Rm 801, On Hong Commerical Building, 145 Hennessy
Wanchai, NON United States Not_Applicable
Hong Kong

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

852 2391 4685 Company Fax:

852 2330 4818 Support Phone Number:

852 2391 4685 Contact Email Address:

support@yoric.com Company URL:

http://www.yoric.com Primary Contact:

Vincent Chan

1.184 Consultant ID: Hungary, BalaBit Bt.

     42. Bem Street
     Szekszard, NON United States 7100
     Hungary

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

++36209683602 Company Fax:

Not_Applicable Support Phone Number:

++36209683602 Contact Email Address:

support@balabit.hu Company URL:

http://www.balabit.hu Primary Contact:

Balazs Scheidler

1.185 Consultant ID: India, Atanu M

     H1/55, Mahavir Enclave, Palam-Dabri Road
     New Delhi, NON United States 110045
     India

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

91-11-5033742 91-11-Company Fax:

Not_Applicable Support Phone Number:

91-11-5033742 91-11-Contact Email Address:

atanu@poboxes.com Company URL:

http://Not_Applicable Primary Contact:

Atanu Not_Applicable

1.186 Consultant ID: India, C Consulting

     15/5 18th Cross
     Bangalore, NON United States 560055
     India

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+91 (80) 3440397 Company Fax:

+91 (80) 3341137 Support Phone Number:

+91 (80) 3440397 Contact Email Address:

linux-support@cbconsulting.com Company URL:

http://www.cbconsulting.com Primary Contact:

Atul Chitnis

1.187 Consultant ID: India, Connect Solutions

     Not Applicable
     Not Appicable, NON United States 10
     India

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

Easy, Debian, RedHat Specialty Distribution:

Other Other Specialty Distribution:

Testing all kinds of Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

not applicable Company Fax:

not applicable Support Phone Number:

not applicable Contact Email Address:

sriram@connectsolutions.net Company URL:

http://www.connectsolutions.net Primary Contact:

Sriram V

1.188 Consultant ID: India, DSF Internet Services

C-29 Neeti Bagh
New Delhi, NON United States 110049
India

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

91-11-685-6158 Company Fax:

91-11-685-6157 Support Phone Number:

91-11-685-6158 Contact Email Address:

support@indoworld.com Company URL:

http://www.indoworld.com Primary Contact:

Anmol Taneja

1.189 Consultant ID: India, GratiSource Solutions (P) Ltd.

#308, Shakti Sai Apartments, Chapel Rd, Nampally
Hyderabad, NON United States 500001
India

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Debian, Caldera and Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

91-40-3391937 Company Fax:

91-40-3391937 Support Phone Number:

91-40-3391937 Contact Email Address:

sidcarter@softhome.net Company URL:

http://gratisource.8m.com Primary Contact:

Khader Syed

1.190 Consultant ID: India, K K Software Services

     B 2, Murugan Towers NAD Kotha Rd
     Visakhapatnam, NON United States 530009
     India

Consulting Specialties:

Custom Systems System Administration Main Distribution:

Other Other Distribution:

Not_Applicable Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+91-891-521102 Company Fax:

+91-891-596571 Support Phone Number:

Not_Applicable Contact Email Address:

vijaykittu@yahoo.com Company URL:

http://vizag.8m.com Primary Contact:

Vijaya Kittu Manda

1.191 Consultant ID: India, Lateral Software Technology Pvt Ltd

     86 TTK Road
     Madras, NON United States 600 018
     India

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

Easy Linux Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+91=044-4671560 Company Fax:

+91-044-4910740 Support Phone Number:

Not_Applicable Contact Email Address:

linux@lateralsoftware.com Company URL:

http://lateralsoftware.com Primary Contact:

Roshan Thomas

1.192 Consultant ID: India, Netangle Com Pvt ltd

     # 2 , Vishal Bhavan , Road no 12, Banjara hills
     Hyderabad, ZZ 500034
     India

Consulting Specialties:

CIFS (Samba) Main Distribution:

RedHat Other Distribution:

Debian Specialty Distribution:

Other Other Specialty Distribution:

RPM,TAR Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

091-40-3300394 Company Fax:

091-40-3391412 Support Phone Number:

091-40-3328207 Contact Email Address:

goswami@netangle.us-inc.com Company URL:

http://www.netangle.us-inc.com Primary Contact:

satyakam goswami

1.193 Consultant ID: India, Shah Micro System

     103, Orental House, 229/231, Samuel Street, Musjid
     Bombay, NON United States 400003
     India

Consulting Specialties:

CIFS (Samba) Internet Server Development Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

All Linux Flavours Specialty Distribution:

Other Other Specialty Distribution:

All i386 Distributio Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+91-22-8862711 Company Fax:

+91-22-8868214 Support Phone Number:

+91-22-8862711 Contact Email Address:

dinesh@indiamal.com Company URL:

http://http://members.tripod.com/˜ dineshah/ Primary Contact:

Dinesh Shah

1.194 Consultant ID: Indonesia, Nayaga Reka Media Aplika, PT

Jl. Situsari Wetan no 39
Bandung, NON United States 40265
Indonesia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

RedHat, Slackware, M Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+62-22-302764 Company Fax:

n/a Support Phone Number:

+62-22-302764 Contact Email Address:

nayagapt@bdg.centrin.net.id Company URL:

http://nayagarekamedia.hypermart.net/index1.htm Primary Contact:

Andika Triwidada

1.195 Consultant ID: Indonesia, Netindo Pusaka Prima pt.

Duta Harapan Indah  Blok SS/10, Jl. Kapuk Muara
Jakarta, NON United States 14460
Indonesia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+6221 - 6631868 Company Fax:

+6221 - 6623606 Support Phone Number:

n/a Contact Email Address:

support@netindo.co.id Company URL:

http://www.netindo.co.id Primary Contact:

Andy Luhur

1.196 Consultant ID: Indonesia, PT Cakram DataLingga Duaribu

     Jln Gunung Batu 131
     Bogor, NON United States 16610
     Indonesia

Consulting Specialties:

CIFS (Samba) Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+62-251-332122 Company Fax:

+62-251-332122 Support Phone Number:

+62-251-332122 Contact Email Address:

linux-support@cdl2000.or.id Company URL:

http://http://www.cdl2000.or.id Primary Contact:

Maliana Harsoadi

1.197 Consultant ID: Israel, Breakthrough LTD

     Hatzabar 18/9
     Kiron, NON United States 55106
     Israel

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

972-52-442549 Company Fax:

972-3-6356434 Support Phone Number:

972-52-442549 Contact Email Address:

webmaster@breakt.co.il Company URL:

http://www.breakt.co.il Primary Contact:

Alex Rier Not_Applicable

1.198 Consultant ID: Israel, Init Computers

     35 Hadar St.
     Herzlia, NON United States 46326
     Israel

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+972-9-9582971 Company Fax:

+972-9-9582952 Support Phone Number:

Not_Applicable Contact Email Address:

support@init.co.il Company URL:

http://Not_Applicable Primary Contact:

Paolo Supino

1.199 Consultant ID: Israel, Reuven Lerner – Communications Consulting Ltd.

17 Disraeli Street
Haifa, NON United States 34333
Israel

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+972-4-824-2265 Company Fax:

+972-4-826-1219 Support Phone Number:

Not_Applicable Contact Email Address:

reuven@lerner.co.il Company URL:

http://lerner.co.il Primary Contact:

Reuven Lerner

1.200 Consultant ID: Israel, vPrise

     50, Disengof
     Tel-Aviv, NON United States Not_Applicable
     Israel

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

972-54-654576 Company Fax:

972-151-54-654576 Support Phone Number:

972-54-654576 Contact Email Address:

support@vprise.com Company URL:

http://www.vprise.com Primary Contact:

Shai Almog

1.201 Consultant ID: Italy, Alessandro Rubini

     Viale Golgi 72
     PAvia, NON United States 27100
     Italy

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Network Administration Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

ET-Linux (embedded 3 Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+39 0382 529554 Company Fax:

Not_Applicable Support Phone Number:

+39 0382 529554 Contact Email Address:

alessandro.rubini@linux.it Company URL:

http://www.linux.it Primary Contact:

Alessandro Rubini

1.202 Consultant ID: Italy, ERANET snc

13H, Via S. Giuseppe
Conegliano, NON United States 31015
Italy

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+39(0)438-415887 Company Fax:

+39(0)438-415887 Support Phone Number:

+39(0)438-415887 Contact Email Address:

andrea.girotto@usa.net Company URL:

http://www.era-net.it Primary Contact:

Andrea Girotto

1.203 Consultant ID: Italy, Francesco Patamia

     Via G. Marrazzo, 2
     Cropani Marina, NON United States Not_Applicable
     Italy

Consulting Specialties:

CIFS (Samba) E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+390961-961486 Company Fax:

+390961-961155 Support Phone Number:

+390338-8165310 Contact Email Address:

patamia@biponline.com Company URL:

http://http://www.biponline.com/fpnet Primary Contact:

Francesco Patamia

1.204 Consultant ID: Italy, Luca Perugini

     Via Furio Camillo 27
     Passo Corese (RI), NON United States 02036
     Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+390765488457 Company Fax:

Not_Applicable Support Phone Number:

+3903473392355 Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.205 Consultant ID: Italy, Marco Iannacone

     Not_Applicable
     Milano, NON United States I-20090
     Italy

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Red Hat, Slackware, Specialty Distribution:

Other Other Specialty Distribution:

Securty, Network int Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.pippo.com Primary Contact:

Marco Iannacone

1.206 Consultant ID: Italy, Marco Michelino

     Viale Colli Aminei 36
     Napoli, NON United States 80131
     Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+390817418669 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

mmarco@hempseed.com Company URL:

http://www.hempseed.com Primary Contact:

Marco Michelino

1.207 Consultant ID: Italy, Mauro Barella

Via Roma, 8
UDINE, NON United States 33100
Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+39 0432 512592 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.maurobarella.org Primary Contact:

Mauro Barella

1.208 Consultant ID: Italy, PROFUSO

Not_Applicable
Abano Terme - Padova, NON United States 35031
Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+39 0498059070 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

info@profuso.com Company URL:

http://http://www.profuso.com/ Primary Contact:

Giuseppe Zanetti

1.209 Consultant ID: Italy, PROSA Progettazione Sviluppo Aperto

     via Degli Zabarella, 3
     Padova, NON United States 35121
     Italy

Consulting Specialties:

Custom Programming Main Distribution:

Debian Other Distribution:

PROSA Debian GNU/Lin Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+39 049 66 05 19 Company Fax:

+39 049 87 80 504 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Davide Barbieri

1.210 Consultant ID: Italy, Panservice

Via Nervi - CC LT Fiori - Torre 8
Latina, NON United States 04100
Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

39 0773 410020 Company Fax:

39 0773 410020 Support Phone Number:

39 0773 410084 Contact Email Address:

helpdesk@panservice.it Company URL:

http://http://www.panservice.it Primary Contact:

Giuliano Peritore

1.211 Consultant ID: Italy, Paolo Correnti

     Via dei Monti di S.Paolo 55/O
     Rome, NON United States 00126
     Italy

Consulting Specialties:

Custom Systems Internet Server Development Internet (Web) Programming System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

3965250639 Company Fax:

— Support Phone Number:

—- Contact Email Address:

pcorrenti@iol.it Company URL:

http://users.iol.it/pcorrenti/index.html Primary Contact:

Paolo Correnti

1.212 Consultant ID: Italy, Prometeus di Michele Sciabarrà

     via Marco Polo 31
     Martinsicuro, NON United States 64014
     Italy

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

39 0861 761767 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

michele@sciabarra.com Company URL:

http://www.sciabarra.com Primary Contact:

Michele Sciabarrà

1.213 Consultant ID: Italy, Studio C

via Romeo Romei 23
Rome, NON United States 00136
Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+39639722036 Company Fax:

+39639736598 Support Phone Number:

+39639722036 Contact Email Address:

albe@vdi.net Company URL:

http://www.studioc-advertising.com Primary Contact:

Marco Capecci

1.214 Consultant ID: Italy, Studio LEADER Pro

     Via Pietrastretta 76
     TRENTO, NON United States 38100
     Italy

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Network Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Caldera RedHat Specialty Distribution:

LinuxRouter Other Specialty Distribution:

HAL91 Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera Not Applicable Company Phone Number:

+39 461828229 Company Fax:

+39 461829826 Support Phone Number:

Not Available Contact Email Address:

info@leader.it Company URL:

http://www.leader.it Primary Contact:

Guido Brugnara

1.215 Consultant ID: Japan, M

     1-8-19 Toshimaku Minami Ikebukuro
     Tokyo, NON United States 171
     Japan

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+81 3 5396 5966 Company Fax:

+81 3 3983 4542 Support Phone Number:

+81 3 5396 5966 Contact Email Address:

fisher@m-t.com Company URL:

http://www.m-t.com Primary Contact:

Masahiro Otsuka

1.216 Consultant ID: Kuwait, Internet Universe - Network Solutions

     Dasma
     P.O.Box 2, NON United States non
     Kuwait

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

General Linux networ Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+965-262-5678 Company Fax:

+965-262-5679 Support Phone Number:

+965-262-5678 Contact Email Address:

support@kuwait-net.com Company URL:

http://www.kuwait-net.com Primary Contact:

Faisal Al-Abhoul

1.217 Consultant ID: Malaysia, JirehBuilderZ Web Techs Snd. Bhd.

Unit 5.06, Fifth Menara PSCI,
Penang, NON United States 10050
Malaysia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

SlackWare Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

604-3700378 Company Fax:

604-3700377 Support Phone Number:

604-3700378 Contact Email Address:

info@jirehtechs.com Company URL:

http://www.jirehtechs.com Primary Contact:

Wai Boon Ng

1.218 Consultant ID: Malaysia, Linux Resources Sdn. Bhd.

No. 33-1, Jalan 46A/26, Pusat Bandar Taman Sri Ram
Setapak, NON United States 53300
Malaysia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+603-4139010 Company Fax:

+603-4139011 Support Phone Number:

+603-4139010 Contact Email Address:

drliwa@pc.jaring.my Company URL:

http://www.ihsan.com/linux Primary Contact:

Dr. Liwaudin Muhamad

1.219 Consultant ID: Malaysia, MAGNIFIX

     2102 Wisma Rampai, Jln 34/26
     Kuala Lumpur, NON United States 53300
     Malaysia

Consulting Specialties:

Internet Server Development Main Distribution:

RedHat Other Distribution:

SUSE Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Linux Certifications:

RedHat Company Phone Number:

603 412 1775 Company Fax:

603 412 1550 Support Phone Number:

603 412 1776 Contact Email Address:

info@magnifix.com.my Company URL:

http://www.magnifix.com.my Primary Contact:

Izauddin Isa

1.220 Consultant ID: Mexico, BUFETE CONSULTOR DE MEXICO

APDO POSTAL 105-336
Mexico CityMMexico City, NON United States 11590
Mexico

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

DEBIAN SUSE SLACKWAR Specialty Distribution:

Other Other Specialty Distribution:

EPHYRA CLUSTERS, RDB Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

+(52)5247-0272 Company Fax:

+(52)5247-0272 Support Phone Number:

Not_Applicable Contact Email Address:

support@ephyra.com Company URL:

http://ephyra.com Primary Contact:

FELIPE BAROUSSE

1.221 Consultant ID: Mexico, BUFETE CONSULTOR DE MEXICO

APDO POSTAL 105-336
Mexico CityMMexico City, NON United States 11590
Mexico

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

DEBIAN SUSE SLACKWAR Specialty Distribution:

Other Other Specialty Distribution:

EPHYRA CLUSTERS, RDB Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

+(52)5247-0272 Company Fax:

+(52)5247-0272 Support Phone Number:

Not_Applicable Contact Email Address:

support@ephyra.com Company URL:

http://ephyra.com Primary Contact:

FELIPE BAROUSSE

1.222 Consultant ID: Mexico, Ephyra Systems (BCM)

APDO POSTAL 105336
Mexico City, NON United States 11590
Mexico

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

DEBIAN SUSE SLACKWAR Specialty Distribution:

Other Other Specialty Distribution:

EPHYRA CLUSTERS, RDB Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

+(52)5247-0272 Company Fax:

+(52)5247-0272 Support Phone Number:

Not_Applicable Contact Email Address:

info@ephyra.com Company URL:

http://ephyra.com Primary Contact:

Phillip BAROUSSE

1.223 Consultant ID: Mexico, Gustavo Cavazos

Not_Applicable
Mexico City, NON United States Not_Applicable
Mexico

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

gcavazos@yahoo.com Company URL:

http://Not_Applicable Primary Contact:

Gustavo Cavazos

1.224 Consultant ID: Mexico, Linux Michoacan OpenSource

186 Gral Antonio Cabales
Morelia, NON United States 58000
Mexico

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

CHARAL Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+5243 246824 Company Fax:

+5243 246824 Support Phone Number:

+5243 246824 Contact Email Address:

linux@ciencia.dcc.umich.mx Company URL:

http://www.ciencia.dcc.umich.mx Primary Contact:

Cederik De Leon

1.225 Consultant ID: Mexico, Servicios en Alta Tecnologia

     Jose Ma. Chavez #420, Centro
     Aguascalientes, NON United States 20000
     Mexico

Consulting Specialties:

System Administration Main Distribution:

RedHat Other Distribution:

SUSE Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

52(49)15-70-26 Company Fax:

52(49)15-70-26 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Hector Vega

1.226 Consultant ID: Mexico, Virtual.Net

     Martin Torres #200
     Xalapa, Veracruz, NON United States 91020
     Mexico

Consulting Specialties:

E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware 4.0, Suse Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+52 (28) 135187 Company Fax:

+52 (28) 135187 Support Phone Number:

+52 (28) 249063 Contact Email Address:

soporte@virtual.net.mx Company URL:

http://www.virtual.net.mx Primary Contact:

Roman Torres

1.227 Consultant ID: Morocco, ANNOUR Technologies

     89 Bd Anfa, Appt. 9 A,
     CASABLANCA, NON United States 20000
     Morocco

Consulting Specialties:

E-Commerce Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

RedHat Company Phone Number:

212 2 48 11 34 Company Fax:

212 2 48 70 56 Support Phone Number:

212 2 48 68 09 Contact Email Address:

info@annour.com Company URL:

http://www.annour.com Primary Contact:

Karim BENABID Hicham BENABID

1.228 Consultant ID: Netherlands, Abe I.T. B.V.

     Oude haagweg 295
     Den Haag, NON United States 2552 EP
     Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31704400368 Company Fax:

+31704400369 Support Phone Number:

+31704400368 Contact Email Address:

info@aboveit.nl Company URL:

http://www.aboveit.nl Primary Contact:

Alex Schreuder

1.229 Consultant ID: Netherlands, BitWizard B.V.

     Van Bronckhorststraat 12
     Delft, NON United States 2612 XV
     Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31-15-2137555 Company Fax:

+31-15-2138217 Support Phone Number:

+31-15-2137555 Contact Email Address:

info@BitWizard.nl Company URL:

http://www.BitWizard.nl Primary Contact:

Roger Wolff

1.230 Consultant ID: Netherlands, CMG

Kralingseweg 241-249
Rotterdam, NON United States 3009 AN
Netherlands

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

+31.10.2537000 Company Fax:

+31.2537032 Support Phone Number:

Not_Applicable Contact Email Address:

Henk.Drenth@cmg.nl Company URL:

http://www.cmg.nl Primary Contact:

Henk Drenth

1.231 Consultant ID: Netherlands, CistroN

38, Ondernemingsweg
Alphen aan den Rijn, NON United States 2404 HN
Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31-172-421071 Company Fax:

+31-172-430979 Support Phone Number:

n/a Contact Email Address:

linux@cistron.nl Company URL:

http://www.cistron.nl Primary Contact:

Danny ter Haar

1.232 Consultant ID: Netherlands, ComCare IT BV

     Papaverweg 34
     Amsterdam, NON United States 1032 KJ
     Netherlands

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

RedHat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31 20 4942165 Company Fax:

+31 20 4942164 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.comcare.nl Primary Contact:

Sjoerd Venema

1.233 Consultant ID: Netherlands, Le Reseau netwerksystemen BV

Bieslookstraat 31, Unit A3
Groningen, NON United States NL-9731 HH
Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Secure Linux (own di Specialty Distribution:

UltraPenguin Other Specialty Distribution:

and LinuxRouter Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31 50 549 2701 Company Fax:

+31 50 549 2310 Support Phone Number:

+31 50 549 4548 Contact Email Address:

support@reseau.nl Company URL:

http://www.reseau.nl/ Primary Contact:

Arthur Donkers Rudo de Jong

1.234 Consultant ID: Netherlands, Linvision

     Delftechpark 26
     Delft, NON United States 2628 XH
     Netherlands

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet (Web) Programming System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31 15 260 04 33 Company Fax:

+31 15 260 04 05 Support Phone Number:

Not_Applicable Contact Email Address:

info@linvision.com Company URL:

http://www.linvision.com Primary Contact:

Geerten Schram

1.235 Consultant ID: Netherlands, Manna Automatisering

     Meteorenstraat 116
     Hilversum, NON United States 1223 EV
     Netherlands

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Advise Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

31356423980 Company Fax:

31356423981 Support Phone Number:

31356423982 Contact Email Address:

info@manna.nl Company URL:

http://www.manna.nl Primary Contact:

Not_Applicable Not_Applicable

1.236 Consultant ID: Netherlands, Maxfort Consultancy

Albardastraat 11
Schiedam, NON United States 3119 PA
Netherlands

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Debian, TurboLinux, Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Linux Certifications:

Not Applicable Company Phone Number:

0031102469224 Company Fax:

0031102469225 Support Phone Number:

Not_Applicable Contact Email Address:

helpdesk@maxfort.com Company URL:

http://www.maxfort.com Primary Contact:

Arjan van Eersel

1.237 Consultant ID: Netherlands, Murphy Software BV

     P.O. Box 285
     Enschede, NON United States NL-7500 AG
     Netherlands

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

31 53 4320055 Company Fax:

31 53 5360448 Support Phone Number:

Not_Applicable Contact Email Address:

office@murphy.nl Company URL:

http://www.murphy.nl Primary Contact:

Marc Roelofs

1.238 Consultant ID: Netherlands, Netland Internet Services

     Kruislaan 419
     Amsterdam, NON United States 1098 VA
     Netherlands

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

our own Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31 20 562 8282 Company Fax:

+31 20 562 8281 Support Phone Number:

+31 20 562 8280 Contact Email Address:

support@netland.nl Company URL:

http://www.netland.nl Primary Contact:

Ron Arts

1.239 Consultant ID: Netherlands, Romme SE

     Cornelis Trompstraat 90
     Delft, NON United States 2628 RS
     Netherlands

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31-15-2855904 Company Fax:

+31-15-2855905 Support Phone Number:

Not_Applicable Contact Email Address:

akram@xs4all.nl Company URL:

http://Not_Applicable Primary Contact:

Johan Romme

1.240 Consultant ID: Netherlands, Root Systems

van Kinsbergenstraat 73
's Gravenhage, ZZ 2518 GW
Netherlands

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(+31) 70 3460010 Company Fax:

(+31) 70 3460118 Support Phone Number:

(+31) 703460126 Contact Email Address:

support@rootsystems.nl Company URL:

http://www.rootsystems.nl Primary Contact:

Conrad Langford

1.241 Consultant ID: Netherlands, Stone IT Services BV

     Jan Steenlaan 3
     Bilthoven, NON United States 3723 BS
     Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

S.u.S.E. Reseller Authorizations:

Compaq Linux Certifications:

Caldera RedHat Company Phone Number:

+31 30 2287995 Company Fax:

+31 30 2287997 Support Phone Number:

+31 30 2294653 Contact Email Address:

support@stonebv.com Company URL:

http://www.stonebv.com Primary Contact:

M.J. Smit

1.242 Consultant ID: Netherlands, X OS Experts in Open Systems BV

     Kruislaan 419
     Amsterdam, NON United States 1098 VA
     Netherlands

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+31 20 6938364 Company Fax:

+31 20 6948204 Support Phone Number:

+31 20 6938364 Contact Email Address:

support@xos.nl Company URL:

http://www.xos.nl Primary Contact:

Jos Vos

1.243 Consultant ID: New Zealand, Catalyst IT Ltd

PO Box 10-225, Level 21, Morrison Kent House, 105
Wellington, NON United States 6000
New Zealand

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+64 (4) 499 2267 Company Fax:

+64 (4) 499 5596 Support Phone Number:

+64 (4) 499 2267 Contact Email Address:

linux@cat-it.co.nz Company URL:

http://www.cat-it.co.nz Primary Contact:

Andrew McMillan

1.244 Consultant ID: New Zealand, Egressive Limited

P.O. Box 24 162
Christchurch, NON United States 8006
New Zealand

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

General Linux instal Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

64 25 229 8147 Company Fax:

Not_Applicable Support Phone Number:

64 25 229 8147 Contact Email Address:

support@egressive.co.nz Company URL:

http://www.egressive.co.nz Primary Contact:

David Lane

1.245 Consultant ID: Nigeria, CyberAge Systems

6, Rasaq Balogun Str,
Surulere, Lagos, NON United States Not_Applicable
Nigeria

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Others Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

234 01 2602080 Company Fax:

234 01 2602080 Support Phone Number:

Not_Applicable Contact Email Address:

cyberservice@eudoramail.com Company URL:

http://Not_Applicable Primary Contact:

Edwin Okugbo

1.246 Consultant ID: Norway, AQS AS

P.B. 121
Eidsvoll, NON United States 2080
Norway

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

47 63962296 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.aqs.as Primary Contact:

Martin Langsjøen

1.247 Consultant ID: Norway, Initio IT-losninger

Holtermannsv. 31B
Trondheim, NON United States 7031
Norway

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+47 - 73 83 71 71 Company Fax:

+47 - 73 83 71 70 Support Phone Number:

+47 - 73 83 71 71 Contact Email Address:

initio@initio.no Company URL:

http://www.initio.no Primary Contact:

Pal Loberg

1.248 Consultant ID: Not_Applicable, Cybernetrics and Systems Co. Ltd.

34-36 Cipero Rd, Cross Crossing,
San Fernando, NON United States TTWI
Not_Applicable

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 868 653 5960 Company Fax:

1 868 659 6604 Support Phone Number:

Not_Applicable Contact Email Address:

info@cysco.com Company URL:

http://www.cysco.com Primary Contact:

Kerry Panchoo

1.249 Consultant ID: Not_Applicable, L and L Develo Oy

     Sturenkatu 37-41
     HELSINKI, NON United States 00550
     Not_Applicable

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

v5.2 Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+358-40-500 5635 Company Fax:

+358-9-7269 2988 Support Phone Number:

+358-9-7269 2980 Contact Email Address:

develo@saunalahti.fi Company URL:

http://www.develo.fi Primary Contact:

Tomi Lind

1.250 Consultant ID: Pakistan, Atlantic 2000 Micro Care

Al-Habib Arcade, Near Shell Petrol Pump,  Clifton
Karachi, NON United States xxx
Pakistan

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

92 21 5872379 Company Fax:

92 21 5872367 Support Phone Number:

92 21 5872379 Contact Email Address:

info@atlantic2000.com Company URL:

http://www.atlantic2000.com Primary Contact:

Farhan Ghafoor

1.251 Consultant ID: Pakistan, GeneralSoft (Pvt) Ltd

     179-C-1 Model Town
     Lahore, NON United States 54700
     Pakistan

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+92 42 5836152 Company Fax:

+92 42 5834920 Support Phone Number:

+92 42 5880851 Contact Email Address:

linux@generalsoft.com Company URL:

http://www.generalsoft.com Primary Contact:

Abid Nasim

1.252 Consultant ID: Pakistan, Instec Digital Systems

904 Fortune Centre, 45A Block 6 PECHS
Karachi, NON United States 75400
Pakistan

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

9221-4547281 Company Fax:

9221-4547282 Support Phone Number:

9221-4547283 Contact Email Address:

support@instecdigital.com Company URL:

http://www.instecdigital.com Primary Contact:

Mudassir Khan

1.253 Consultant ID: Philippines, MISTRONIX

153 Julio Pacana St
Cagayan de Oro, NON United States 9000
Philippines

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

SLACKWARE Specialty Distribution:

Other Other Specialty Distribution:

SLACKWARE Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

63 (8822) 724249 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

mistronix@linuxstart.com Company URL:

http://www.linuxstart.com/˜ mistronix Primary Contact:

Noel Mistula

1.254 Consultant ID: Poland, AM ELEKTRONIK s. c.

     ul. Biala 7
     Gdansk, NON United States 80-435
     Poland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+48 58 3440061 Company Fax:

+48 58 3440061 Support Phone Number:

Not Available Contact Email Address:

amelektr@itnet.pl Company URL:

http://www.itnet.pl/amelektr Primary Contact:

Marek Michalkiewicz

1.255 Consultant ID: Poland, Internet Designers s.c.

     Rynek 58
     Wroclaw, NON United States 50-116
     Poland

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+48713410099 Company Fax:

+48713410099 Support Phone Number:

+48713460827 Contact Email Address:

support@id.pl Company URL:

http://www.id.pl Primary Contact:

Ryszard Lach

1.256 Consultant ID: Poland, RG Studio

Limanowskiego 23
Warszawa, NON United States 02-934
Poland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Cobalt Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+48 22 651 66 38 Company Fax:

+48 22 651 66 38 Support Phone Number:

+48 22 651 66 38 Contact Email Address:

info@rgstudio.com.pl Company URL:

http://www.rgstudio.com.pl Primary Contact:

Jan Rychter

1.257 Consultant ID: Poland, Softomat Ltd.

     Krynicka 1
     Gdansk, NON United States 80-393
     Poland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+48 58 5543 730 Company Fax:

+48 58 556 11 76 Support Phone Number:

+48 58 5543 731 Contact Email Address:

softomat@softomat.com.pl Company URL:

http://www.softomat.com.pl Primary Contact:

Jacek Kijewski

1.258 Consultant ID: Portugal, COM2000 Sistemas Informaticos

     APARTADO 13098
     LISBOA, NON United States 1019
     Portugal

Consulting Specialties:

Internet Server Development Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

+351 1 8478838 Company Fax:

+351 1 8475438 Support Phone Number:

+351 1 8478838 Contact Email Address:

linux@com2000si.com Company URL:

http://www.com2000si.com Primary Contact:

Not_Applicable Not_Applicable

1.259 Consultant ID: Portugal, S CONSULTORES LDA

     R. Prof. Fernando Fonseca, 16
     Lisboa, NON United States 1600-618
     Portugal

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not_Applicable Company Phone Number:

351 1 757 17 18 Company Fax:

351 1 757 17 33 Support Phone Number:

351 1 7520017 Contact Email Address:

sf@sfconsultores.pt Company URL:

http://www.sfconsultores.pt Primary Contact:

José Ferradeira

1.260 Consultant ID: Puerto Rico, EZ Networks

PO Box 3534
Guaynabo, PR 00970
Puerto Rico

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

787-289-7801 Company Fax:

787-289-8779 Support Phone Number:

NA Contact Email Address:

info@ez-networks.com Company URL:

http://www.ez-networks.com Primary Contact:

Jose Torres

1.261 Consultant ID: Romania, EVERCOMM Serv SRL

Str. Ciucea nr.8, bl.L18, ap.189
Bucharest, NON United States 7400
Romania

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+40-9-254-4355 Company Fax:

+40-1-345-4355 Support Phone Number:

+40-9-254-4355 Contact Email Address:

support@evercomm.ro Company URL:

http://www.evercomm.ro Primary Contact:

Mihnea Stoenescu

1.262 Consultant ID: Russia, Sofia Ltd.

PO Box 188
Kazan, NON United States 420101
Russia

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Mandrake Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+7-8432-343469 Company Fax:

NA Support Phone Number:

Not_Applicable Contact Email Address:

sherba@sofia.cgb18.kcn.ru Company URL:

http://NA Primary Contact:

Andrew Sherba

1.263 Consultant ID: Singapore, Aeon Xe Technology Pte Ltd

     38 Edgware Road
     Singapore, NON United States 799781
     Singapore

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

Redhat, Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

654846834 Company Fax:

654842687 Support Phone Number:

Not Available Contact Email Address:

support@aeonxe.com Company URL:

http://www.aeonxe.com Primary Contact:

Eddie Tan

1.264 Consultant ID: Singapore, LinuxAsia Pte Ltd

     No. 2 Alexandra Road, #05-03 Delta House
     Singapore, NON United States 159919
     Singapore

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

65-272 5482 Company Fax:

65-234 2400 Support Phone Number:

65-272 5482 Contact Email Address:

support@linuxasia.com Company URL:

http://www.linuxasia.com Primary Contact:

Stephane Benoist

1.265 Consultant ID: Singapore, Webworks Pte Ltd

37, Lorong 23 Geylang, #03-04, Yu Li Industrial Bu
Singapore, NON United States 388371
Singapore

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Linux Certifications:

Not Applicable Company Phone Number:

+65 741 9526 Company Fax:

+65 749 3806 Support Phone Number:

+65 741 9526 Contact Email Address:

support@webworks.com.sg Company URL:

http://www.webworks.com.sg Primary Contact:

Lon Voon Ong

1.266 Consultant ID: South Africa, Computer Development Systems

P.O.Box 40012
Cleveland, NON United States 2022
South Africa

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

Industrial Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+27-11-615-7983 Company Fax:

+27-11-615-8650 Support Phone Number:

+27-11-615-7983 Contact Email Address:

support@lrlabs.com Company URL:

http://www.lrlabs.com Primary Contact:

Linda Lessing

1.267 Consultant ID: South Africa, Neogenix

Pineslopes Office Park; East Block; Witkoppen Road
Johannesburg, NON United States Not_Applicable
South Africa

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

RedHat Company Phone Number:

+27 (0)11 465-0798 Company Fax:

+27 (0)11 465-9688 Support Phone Number:

+27 (0)11 465-0796 Contact Email Address:

marius@neogenix.co.za Company URL:

http://www.neogenix.co.za Primary Contact:

Theunis Botha

1.268 Consultant ID: South Africa, Obsidian Systems

Unit 3, Boskruin Business Park, Boskruin
Johannesburg, NON United States 2156
South Africa

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+27-11-792-6500 Company Fax:

+27-11-792-6522 Support Phone Number:

+27-11-792-6500 Contact Email Address:

support@obsidian.co.za Company URL:

http://www.obsidian.co.za Primary Contact:

Anton de Wet

1.269 Consultant ID: South Africa, Qualica Technologies (Pty) Ltd

28 8th Avenue, Parktown North
Gauteng, NON United States 2193
South Africa

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Jurix Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+27 11 442 9090 Company Fax:

+27 11 442 9090 Support Phone Number:

+27 11 442 9090 Contact Email Address:

support@qualica.com Company URL:

http://www.qualica.com Primary Contact:

Kezia Cousins

1.270 Consultant ID: Spain, ACCESSUS - Proyectos y Consultores en Telecomunica

Moctezuma, 4 - 1
Santander - Cantabria, NON United States 39003
Spain

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+34 942 31 80 55 Company Fax:

+34 942 31 80 56 Support Phone Number:

+34 942 31 80 55 Contact Email Address:

seguridad@grupoaccessus.com Company URL:

http://www.grupoaccessus.com Primary Contact:

Jose Antonio Revilla

1.271 Consultant ID: Spain, Atica Sistemas, SL

Ronda Pío XII, 1-A, 2, mod. 5
Sevilla, NON United States 41008
Spain

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+34-954368107 Company Fax:

+34-954367159 Support Phone Number:

+34-954368107 Contact Email Address:

soporte-linux@aticasistemas.com Company URL:

http://www.aticasistemas.com Primary Contact:

Marco Antonio Blanco

1.272 Consultant ID: Spain, Innovacion y Futuro s.l.

     Plaza de S'Antigor, 3
     Manacor, NON United States 07500
     Spain

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Netware Connectivity Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Caldera Company Phone Number:

+34 971846014 Company Fax:

+34 971846014 Support Phone Number:

+34 971846014 Contact Email Address:

info@i-futuro.com Company URL:

http://www.i-futuro.com Primary Contact:

Pascal Peregrina

1.273 Consultant ID: Spain, LANGRE INGENIEROS S.L.

     Castilla 95
     Santander, NON United States 39009
     Spain

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

34-942-33-56-33 Company Fax:

34-942-35-42-75 Support Phone Number:

34-942-33-56-33 Contact Email Address:

linux@langre.com Company URL:

http://www.langre.com Primary Contact:

Arturo Limon

1.274 Consultant ID: Spain, Linux Sistemas

     C/ Emilio Carrere, 8 5izq
     Madrid, ZZ Not_Applicable
     Spain

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+34914459494 Company Fax:

Not_Applicable Support Phone Number:

+34914459494 Contact Email Address:

vaceituno@hotmail.com Company URL:

http://www.hotmail.com Primary Contact:

Vicente Aceituno

1.275 Consultant ID: Sweden, Ewing Data AB

Götgatan 39
Stockholm, NON United States 11621
Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

RedHat, Debian, Cald Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+46 (0)8 643-1106 Company Fax:

+46 (0)8 643-1106 Support Phone Number:

+46 (0)8 643-1106 Contact Email Address:

admin@ewingdata.com Company URL:

http://www.ewingdata.com Primary Contact:

James Ewing

1.276 Consultant ID: Sweden, Mandator Sverige AB

     Box 6074
     Göteborg, NON United States 400 60
     Sweden

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

031-7076900 Company Fax:

031-7076969 Support Phone Number:

031-7076990 Contact Email Address:

helpdesk@mandator.se Company URL:

http://www.mandator.se Primary Contact:

Ulrik Holmén

1.277 Consultant ID: Sweden, Omicron Software Systems AB

Borgarfjordsgatan 13
Kista, NON United States S-164 40
Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

ALL Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

46 8 7510180 Company Fax:

46 8 7039413 Support Phone Number:

46 8 7512940 Contact Email Address:

support@omicron.se Company URL:

http://www.omicron.se Primary Contact:

Johan Nordin

1.278 Consultant ID: Sweden, South Pole AB

Banvaktsvägen 12
Solna, NON United States 171 48
Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+46 (0)8 566 106 50 Company Fax:

+46 (0)8 566 106 01 Support Phone Number:

Not_Applicable Contact Email Address:

support@southpole.se Company URL:

http://southpole.se Primary Contact:

Jakob Sandgren

1.279 Consultant ID: Sweden, Tarkus

     Flitiga Lisas v. 8
     Västerås, NON United States 72246
     Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

+33 4 50602749 Support Phone Number:

+46 70 75 22246 Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Carlos Vidal

1.280 Consultant ID: Sweden, Textalk AB

     Krokslätts fabriker 64
     Göteborg, NON United States 43137
     Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

ALL Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+46 31 872920 Company Fax:

+46 31 674894 Support Phone Number:

NA Contact Email Address:

info@textalk.se Company URL:

http://www.textalk.se Primary Contact:

Henryk Rubinstein

1.281 Consultant ID: Sweden, TheLogic

Lingonstigen 24
SKÖVDE, NON United States 541 33
Sweden

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+46-500-459772 Company Fax:

+46-500-459773 Support Phone Number:

+46-500-459772 Contact Email Address:

thomas@thelogic.com Company URL:

http://www.thelogic.com Primary Contact:

Thomas Ekström

1.282 Consultant ID: Sweden, Xantus AB

Bredgatan 10
Lund, NON United States S-222 21
Sweden

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+46 46 399200 Company Fax:

+46 46 399201 Support Phone Number:

+46 46 399200 Contact Email Address:

info@xantus.se Company URL:

http://www.xantus.se Primary Contact:

Henrik Carling

1.283 Consultant ID: Switzerland, 3T

Gassackerweg 47
Frenkendorf, NON United States CH-4402
Switzerland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

RedHat, Stampede Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 61 303 91 91 Company Fax:

+41 61 303 91 92 Support Phone Number:

+41 61 303 91 91 Contact Email Address:

info@t-t-t.com Company URL:

http://www.t-t-t.com Primary Contact:

Stefan Schaefer

1.284 Consultant ID: Switzerland, ALPHANET NF

     Battieux 6c
     Colombier, NON United States 2013
     Switzerland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41328414014 Company Fax:

+41328414081 Support Phone Number:

Not_Applicable Contact Email Address:

schaefer@alphanet.ch Company URL:

http://www-internal.alphanet.ch Primary Contact:

Marc SCHAEFER

1.285 Consultant ID: Switzerland, Born Consulting

     Brunnadernstrasse 3
     Berne, NON United States 3006
     Switzerland

Consulting Specialties:

CIFS (Samba) Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

MkLinux and SPARC Li Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 (0)79 338 95 79 Company Fax:

- Support Phone Number:

+41 (0)79 338 95 79 Contact Email Address:

flborn@spectraweb.ch Company URL:

http://www.spectraweb.ch Primary Contact:

Fritz Lorenz Born

1.286 Consultant ID: Switzerland, CONNEXT Communications

Lebrigstrasse 1
Goetighofen, NON United States 8583
Switzerland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41-71 640 0410 Company Fax:

+41-71 640 0411 Support Phone Number:

+41-71 640 0410 Contact Email Address:

support@connext.ch Company URL:

http://www.connext.ch Primary Contact:

Mauro Canal

1.287 Consultant ID: Switzerland, Coxbit GmbH

Postgasse 6
Zollikofen, NON United States 3052
Switzerland

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 31 9811 60 35 Company Fax:

+41 31 9811 60 35 Support Phone Number:

+41 31 9811 60 35 Contact Email Address:

info@coxbit.ch Company URL:

http://http://www.coxbit.ch Primary Contact:

Henriette Fux

1.288 Consultant ID: Switzerland, Didier Frick, Freelance Software Engineer

     Chasselas 32a
     Peseux, NON United States 2034
     Switzerland

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 (32) 731 49 29 Company Fax:

Not_Applicable Support Phone Number:

+41 (32) 731 49 29 Contact Email Address:

dfrick@dial.eunet.ch Company URL:

http://www.dfr.ch Primary Contact:

Didier Frick

1.289 Consultant ID: Switzerland, ETV Software Engineering S.A.

Ch. du Pavillon 2, PO Box 242
Grand Saconnex, NON United States 1218
Switzerland

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 22 717 64 00 Company Fax:

+41 22 717 64 01 Support Phone Number:

+41 22 717 64 18 Contact Email Address:

linux-support@software-engineering.ch Company URL:

http://www.software-engineering.ch Primary Contact:

Joel Margot

1.290 Consultant ID: Switzerland, Erik Rossen

     Rue Centrale 26
     Chavannes-Renens, NON United States CH-1022
     Switzerland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet (Web) Programming System Administration

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+41 21 636 42 80 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Erik Rossen

1.291 Consultant ID: Switzerland, InfoDevice Sagl

     V. Castelrotto 20
     Locarno, NON United States CH-6600
     Switzerland

Consulting Specialties:

Not_Applicable Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 91 7600168 Company Fax:

+41 91 7600169 Support Phone Number:

+41 91 7600168 Contact Email Address:

infodevice@ticino.com Company URL:

http://Not_Applicable Primary Contact:

Giuseppe Abbatiello

1.292 Consultant ID: Switzerland, Penta Group AG

     Badenerstrasse 701
     ZURICH, NON United States CH-8048
     Switzerland

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Network Administration System Administration

Main Distribution:

Other Other Distribution:

ALL (whatever Custom Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

+41 1 438 50 50 Company Fax:

+41 1 438 50 60 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Benedikt Hochstrasser

1.293 Consultant ID: Switzerland, SciOS Scientific Operating Systems GmbH

     Not_Applicable
     Islikon, NON United States 8546
     Switzerland

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

++41-52-366-52-15 Company Fax:

++41-52-366-52-14 Support Phone Number:

++41-52-366-52-15 Contact Email Address:

support@scios.ch Company URL:

http://www.scios.ch Primary Contact:

Not_Applicable Not_Applicable

1.294 Consultant ID: Switzerland, Technology Solutions

Not_Applicable
Pfaffhausen, NON United States 8118
Switzerland

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Debian, Suse Specialty Distribution:

Other Other Specialty Distribution:

Linux/AXP Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 1 825 28 36 Company Fax:

+41 1 825 28 36 Support Phone Number:

Not Available Contact Email Address:

linux-support@feigin.com Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.295 Consultant ID: Switzerland, Wescom GmbH

     Wesemlinstr. 37
     Luzern, NON United States CH-6002
     Switzerland

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Debian Other Distribution:

RedHat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+41 41 4221050 Company Fax:

+41 41 4221051 Support Phone Number:

+41 41 4221050 Contact Email Address:

mail@wescom.net Company URL:

http://www.wescom.ch Primary Contact:

Markus Fischer

1.296 Consultant ID: Taiwan, Daniel C. Chao

     PO Box 84-511
     Taipei, NON United States 100
     Taiwan

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+886935630111 Company Fax:

+88628731535 Support Phone Number:

+886935630111 Contact Email Address:

danchao@o3.net Company URL:

http://www.o3.net Primary Contact:

Daniel Chao

1.297 Consultant ID: Taiwan, WHINSTONE INFO. CO., LTD.

Not_Applicable
TAIPEI, NON United States 105
Taiwan

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

886-2-27492626 Company Fax:

886-2-27468790 Support Phone Number:

886-2-27492626 Contact Email Address:

info@whinstone.com Company URL:

http://www.whinstone.com Primary Contact:

Anthony Hsia

1.298 Consultant ID: Thailand, Software Advanced Solutions Co., Ltd.

     48 Villa 1, Yaowarat Soi 1
     Phuket, NON United States 83000
     Thailand

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+66 76 218 826 Company Fax:

+66 76 214 041 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.sas.co.th Primary Contact:

Rudiger Koch

1.299 Consultant ID: Togo, HTP Communication Pvt. Ltd

     1/236, Kalikasthan, Dillibazar
     Kathmandu, NON United States 11188
     Togo

Consulting Specialties:

Internet Server Development Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

IBM Not Applicable Linux Certifications:

RedHat Not Applicable Company Phone Number:

9771472072 Company Fax:

9771496620 Support Phone Number:

9771415772 Contact Email Address:

ssupport@htp.com.np Company URL:

http://www.namaste.com.np Primary Contact:

Ranjan Koirala

1.300 Consultant ID: Turkey, Gelecek A.S

     Perpa Ticaret Merkezi Kat:11 B Blok No:1932
     Istanbul, NON United States Not_Applicable
     Turkey

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

EasyLinux Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+90 212 2107194 Company Fax:

+90 212 2100442 Support Phone Number:

Not_Applicable Contact Email Address:

destek@gelecek.com.tr Company URL:

http://www.gelecek.com.tr Primary Contact:

Gorkem Cetin

1.301 Consultant ID: Turkey, YORE Elektronik Yayimcilik A.S.

     Halaskargazi C 301K 5  Sisli
     Istanbul, NON United States 80260
     Turkey

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+90 212 234 0090 Company Fax:

+90 212 230 7142 Support Phone Number:

+90 212 234 0090 Contact Email Address:

destek@yore.com.tr Company URL:

http://http://www.yore.net.tr/ Primary Contact:

Tunc Ozbilgin

1.302 Consultant ID: US, AK, HB Technical Services

     PO Box 60731
     Fairbanks, AK 99706
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

907-479-2403 Company Fax:

none Support Phone Number:

907-479-2403 Contact Email Address:

hbtech@mosquitonet.com Company URL:

http://www.mosquitonet.com/˜ hbtech Primary Contact:

Howard Benson

1.303 Consultant ID: US, AL, Linux Support Services

923 Stage Road Suite F
Auburn, AL 36830
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(877) LINUX-ME (toll Company Fax:

(334) 887-2299 Support Phone Number:

(334) 887-1042 Contact Email Address:

staff@linux-support.net Company URL:

http://www.linux-support.net Primary Contact:

Mark Spencer

1.304 Consultant ID: US, AL, Megabyte Information Services, Inc.

     PO Box 122
     Pell City, AL 35125
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

2052227706 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@meginfo.com Company URL:

http://Not_Applicable Primary Contact:

David Miller

1.305 Consultant ID: US, AL, Technology Unlimited, Inc.

     3336 Summit Drive
     Birmingham, AL 35243
     US

Consulting Specialties:

CIFS (Samba) Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

205-969-1110 Company Fax:

205-969-7113 Support Phone Number:

205-969-1110 Contact Email Address:

support@tuial.com Company URL:

http://www.tuial.com Primary Contact:

Edward McLain

1.306 Consultant ID: US, AL, The Net Effect, LLC

P O Box 885
Mobile, AL 36601
US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

USA (334) 433 0196 Company Fax:

USA (334) 433 5371 Support Phone Number:

USA (334) 433 0196 Contact Email Address:

grs@theneteffect.com Company URL:

http://www.theneteffect.com Primary Contact:

Glenda Snodgrass

1.307 Consultant ID: US, AZ, Brindlesoft

     PO Box 237
     St. David, AZ 85630-0237
     US

Consulting Specialties:

Custom Programming System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(520) 720-9431 Company Fax:

Not_Applicable Support Phone Number:

(520) 720-9431 Contact Email Address:

thegrendel@theriver.com Company URL:

http://www.theriver.com Primary Contact:

Mendel Cooper

1.308 Consultant ID: US, AZ, Jay Ts

P.O. Box 410
Sedona, AZ 86339
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(888) 282-0549 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://http://jayts.cx Primary Contact:

Jay Ts

1.309 Consultant ID: US, AZ, LucidImage Technologies

P.O. Box 875
Chandler, AZ 85224-0875
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet (Web) Programming Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

480 967 6899 Company Fax:

480 641 0248 Support Phone Number:

480 967 6899 Contact Email Address:

support@lucid-image.com Company URL:

http://www.lucid-image.com Primary Contact:

Bill Williams

1.310 Consultant ID: US, AZ, MagusNet, Inc.

     1201 East Washington Ave.
     Gilbert, AZ 85234
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

602-892-8186 Company Fax:

NA Support Phone Number:

NA Contact Email Address:

support@magusnet.com Company URL:

http://www.magusnet.com Primary Contact:

Jean Francois

1.311 Consultant ID: US, AZ, Tapestry Software

     7990 E. Maguey Drive
     Tucson, AZ 85750
     US

Consulting Specialties:

Custom Programming Internet (Web) Programming System Administration System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

01-520-749-2728 Company Fax:

01-520-749-2885 Support Phone Number:

Not_Applicable Contact Email Address:

sbw@tapestry.tucson.az.us Company URL:

http://Not_Applicable Primary Contact:

Steve Wampler

1.312 Consultant ID: US, CA, Aedra Associates, Inc.

Not_Applicable
San Jose, CA 95117
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

408-246-0875 Company Fax:

Not_Applicable Support Phone Number:

408-246-0875 Contact Email Address:

support@aedra.com Company URL:

http://www.aedra.com Primary Contact:

Steve Otto

1.313 Consultant ID: US, CA, Aztronic Enterpises

     Not_Applicable
     San Diego, CA Not_Applicable
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

619 585-1063 Company Fax:

619 585-1063 Support Phone Number:

Not_Applicable Contact Email Address:

support@aztronic.com Company URL:

http://www.aztronic.com Primary Contact:

Randy Church

1.314 Consultant ID: US, CA, Beyond 2000 Solutions

     2750 E. Spring St. Suite 230
     Long Beach, CA 90806
     US

Consulting Specialties:

Firewalls and Internet Security Network Administration System Administration System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Not Applicable Company Phone Number:

562-997-9500 Company Fax:

562-997-9504 Support Phone Number:

n/a Contact Email Address:

abusir@b2000s.com Company URL:

http://www.b2000s.com Primary Contact:

Asem Abusir

1.315 Consultant ID: US, CA, Carumba

55 West 20th Ave Suite #113
San Mateo, CA 94402
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

650 572 2644 Company Fax:

650 572 2644 Support Phone Number:

650 572 2644 Contact Email Address:

support@carumba.com Company URL:

http://Not_Applicable Primary Contact:

Jauder Ho

1.316 Consultant ID: US, CA, Cyber Sleuths

     715 Tender Ln.
     Foster City, CA 94404
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

650-571-7771 Company Fax:

650-571-7771 Support Phone Number:

Not_Applicable Contact Email Address:

evan715@sirius.com Company URL:

http://www.cyber-sleuths.com Primary Contact:

Evan Klinger

1.317 Consultant ID: US, CA, David Fetter Consulting

     888 O'Farrell Street, Suite E1205
     San Francisco, CA 94109
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 415 567 2690 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

dfetter@best.com Company URL:

http://www.best.com Primary Contact:

David Fetter

1.318 Consultant ID: US, CA, Gregg Weber

     Not_Applicable
     El Sobrante, CA 94803
     US

Consulting Specialties:

Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

510 223-3310 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.319 Consultant ID: US, CA, Herlein Engineering

     2034 Filbert Street
     San Francisco, CA 94123
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

415-519-3650 Company Fax:

415-440-9015 Support Phone Number:

415-519-3650 Contact Email Address:

support@herlein.com Company URL:

http://www.herlein.com Primary Contact:

Greg Herlein

1.320 Consultant ID: US, CA, Hiverworld, LLC

     1918 University Av e
     Berkeley, CA 94704
     US

Consulting Specialties:

System Security Virtual Private Networking Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

510.848.0740 Company Fax:

n/a Support Phone Number:

510.848.0740 Contact Email Address:

support@hiverworld.com Company URL:

http://www.hiverworld.com Primary Contact:

Gretchen Flowers

1.321 Consultant ID: US, CA, Intrusoft

     1433 Superior Ave #373
     Newport Beach, CA 92663
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

949-515-3554 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

info@intrusoft.com Company URL:

http://www.intrusoft.com Primary Contact:

Jean-Christophe Smith

1.322 Consultant ID: US, CA, Inverselogic, LLC

     2947A Honolulu Ave,
     Glendale, CA 91214q=
     US

Consulting Specialties:

Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

all Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Caldera RedHat Company Phone Number:

818-542-3941 Company Fax:

818-542-3942 Support Phone Number:

Not_Applicable Contact Email Address:

support@inverselogic.com Company URL:

http://www.inverselogic.com Primary Contact:

Ara Aslanian

1.323 Consultant ID: US, CA, Jalan Network Services

24664 Shepardson Drive
Loma Linda, CA 92354
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Customized Redhat fo Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

909-799-3805 Company Fax:

909-799-3285 Support Phone Number:

909-799-3805 Contact Email Address:

ops@jalan.com Company URL:

http://www.jalan.com Primary Contact:

Matthew Marlowe

1.324 Consultant ID: US, CA, KSI

1677 Azusa Ave, #240
Hacienda Heights, CA 91745
US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

714-937-2469 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@klabs.net Company URL:

http://www.klabs.net Primary Contact:

Dan Kuykendall

1.325 Consultant ID: US, CA, KrystalBox Technologies, Inc.

555 Bryant Street #359
Palo Alto, CA 94301
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

650-327-8929 Company Fax:

650-327-8949 Support Phone Number:

Not_Applicable Contact Email Address:

support@krystalbox.com Company URL:

http://www.krystalbox.com Primary Contact:

Travis Godbout

1.326 Consultant ID: US, CA, Linuxcare

650 Townsend Street
San Francisco, CA 94103
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

We support ALL distr Specialty Distribution:

Other Other Specialty Distribution:

We support ALL distr Reseller Authorizations:

Not_Applicable Linux Certifications:

Caldera RedHat Company Phone Number:

888-546-4878 Company Fax:

415-701-7457 Support Phone Number:

888-546-4878 Contact Email Address:

support@linuxcare.com Company URL:

http://www.linuxcare.com Primary Contact:

David LaDuke

1.327 Consultant ID: US, CA, Lone Peak Software

     164 N. Katherine Drive
     Ventura, CA 93003-1509
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

805-652-1301 Company Fax:

805-652-1301 Support Phone Number:

Not_Applicable Contact Email Address:

lonepeak@pacbell.net Company URL:

http://www.lonepeak.net Primary Contact:

Joe Duran

1.328 Consultant ID: US, CA, MK Information System Services

     Not_Applicable
     Roseville, CA Not_Applicable
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Suse Other Distribution:

Caldera, RedHat Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(916)765-2253 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

blaw@ns.net Company URL:

http://Not_Applicable Primary Contact:

Bob Law

1.329 Consultant ID: US, CA, Mountain Ranch Design

     P.O. Box 679
     Mountain Ranch, CA 95246
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

209-754-3108 Company Fax:

209-754-0273 Support Phone Number:

1-888-551-3003 Contact Email Address:

support@mtnranch.com Company URL:

http://www.mtnranch.com Primary Contact:

Mark Hansen

1.330 Consultant ID: US, CA, Network Presence, LLC

548 S Spring St, Ste 1015
Los Angeles, CA 90013
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

213-627-4356 Company Fax:

213-627-4357 Support Phone Number:

213-627-4356 Contact Email Address:

support@netpr.com Company URL:

http://www.netpr.com Primary Contact:

Robert Brown

1.331 Consultant ID: US, CA, Newyen Technology

3535 El Portal Dr C206
El Sobrante, CA 94803
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

1-510-758-5950 Company Fax:

Not_Applicable Support Phone Number:

1-510-758-5950 Contact Email Address:

sales@newyen.com Company URL:

http://www.newyen.com Primary Contact:

Kent Nguyen

1.332 Consultant ID: US, CA, Oration

     7 West 41st Avenue, Suite 78
     San Mateo, CA 94403
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

877-ORATION Company Fax:

650-572-8315 Support Phone Number:

877-ORATION Contact Email Address:

support@oration.com Company URL:

http://www.oration.com Primary Contact:

Pramod John

1.333 Consultant ID: US, CA, Ozlin Network Designs

     600-A Anton Blvd. #302
     Costa Mesa, CA 92626
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

714-546-4250 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

info@ozlin.com Company URL:

http://www.ozlin.com Primary Contact:

Mark Linn

1.334 Consultant ID: US, CA, PSO Systems Inc

120 Linden Street
Oakland, CA 94607
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

510 663 1156 Company Fax:

510 663 1157 Support Phone Number:

510 663 1156 Contact Email Address:

support@pso.com Company URL:

http://www.pso.com Primary Contact:

Hwajin Bae

1.335 Consultant ID: US, CA, Predictive Science

     2008 Conquista Ave
     Long Beach, CA 90815
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

RedHat & Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1.562.592.6884 Company Fax:

1.562.592.6884 Support Phone Number:

1.562.592.6884 Contact Email Address:

lganders@gte.net Company URL:

http://www.gte.net Primary Contact:

Leonard Anderson

1.336 Consultant ID: US, CA, Softcraft Impresa

     2923 Cohansey Drive
     San Jose, CA 95132
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

Other Other Distribution:

Generations Linux Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@softcraft.com Company URL:

http://www.softcraft.com Primary Contact:

Mauro DePalma

1.337 Consultant ID: US, CA, South Coast Software

     Not_Applicable
     Anaheim, CA 92802
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet Server Development Network Administration System Administration Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

All distributions Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-714-774-6952 Company Fax:

1-714-774-7512 Support Phone Number:

Not_Applicable Contact Email Address:

support@sc-software.com Company URL:

http://www.sc-software.com Primary Contact:

John Heil

1.338 Consultant ID: US, CA, Starshine Technical Services

     903 Harriet Avenue
     Campbell, CA 95008
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Internet Server Development Internet (Web) Programming System Administration System Security

Main Distribution:

Other Other Distribution:

PC-based but includi Specialty Distribution:

Other Other Specialty Distribution:

will aid requirement Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

800-938-4078 Company Fax:

408-374-LUGS Support Phone Number:

(leave voice mail at Contact Email Address:

consulting@starshine.org Company URL:

http://www.starshine.org Primary Contact:

Heather Stern Not_Applicable

1.339 Consultant ID: US, CA, Tesno Systems

3435 Ocean Park Bl Ste 112
Santa Monica, CA 90405
US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

310 2047875 Company Fax:

310 5589221 Support Phone Number:

310 2047875 Contact Email Address:

info@tesnosystems.com Company URL:

http://www.tesnosystems.com Primary Contact:

Jerry Leahy

1.340 Consultant ID: US, CA, U.S. Technical Services

     9121 Atlanta Ave. Suite 230
     Huntington Beach, CA 92646
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(714) 374-6300 Company Fax:

(714) 242-1346 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

brian harvey

1.341 Consultant ID: US, CA, US CustomLogic, LLC

11292 Coloma Road, Suite C
Gold River, CA 95670
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Intel Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-916-853-4920 Company Fax:

1-916-853-4924 Support Phone Number:

1-916-853-4920 Contact Email Address:

info@customlogic.com Company URL:

http://www.customlogic.com Primary Contact:

Tom Henderson

1.342 Consultant ID: US, CA, Vennerable Consulting

1563 Solano Avenue, Suite 516
berkeley, CA 94707
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

510 528 8072 Company Fax:

510 528 8072 Support Phone Number:

510 528 8072 Contact Email Address:

support@vennerable.com Company URL:

http://www.vennerable.com Primary Contact:

jason venner

1.343 Consultant ID: US, CA, interactivate consulting group

2244b carmel valley road
del mar, CA 92014
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

red hat, slackware, Specialty Distribution:

Other Other Specialty Distribution:

custom Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

please use email Company Fax:

619 793 4069 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

sean dreilinger

1.344 Consultant ID: US, CO, BDA Consulting LLC

     758 English Sparrow Trail Suite 100
     Highlands Ranch, CO 80126
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

firewalls Reseller Authorizations:

Not Applicable Linux Certifications:

Not_Applicable Company Phone Number:

303-683-9252 Company Fax:

303-683-7439 Support Phone Number:

303-683-9252 Contact Email Address:

support@bdaconsulting.com Company URL:

http://www.bdaconsulting.com Primary Contact:

James Davis

1.345 Consultant ID: US, CO, Eklektix, Inc.

     5624 Rim Rock Ct
     Boulder, CO 80301-3553
     US

Consulting Specialties:

Custom Programming Network Administration System Administration System Security Main Distribution:

RedHat Other Distribution:

All others Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

303 581 0750 Company Fax:

303 581 0385 Support Phone Number:

303 581 0750 Contact Email Address:

info@eklektix.com Company URL:

http://www.eklektix.com Primary Contact:

Elizabeth Coolbaugh

1.346 Consultant ID: US, CO, FourThought LLC

     3355 Sam Houston
     Ft. Collins, CO 80526
     US

Consulting Specialties:

Custom Programming E-Commerce Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Debian Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

01-970-481-0805 Company Fax:

01-970-513-2884 Support Phone Number:

Not_Applicable Contact Email Address:

consulting@fourthought.com Company URL:

http://FourThought.com Primary Contact:

Uche Ogbuji

1.347 Consultant ID: US, CO, ISP, Ltd.

302 Grand Ave., Box 132
Paonia, CO 81428
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1-970-527-6756 Company Fax:

+1-970-527-4120 Support Phone Number:

Not published Contact Email Address:

techsupport@paonia.com Company URL:

http://www.paonia.com Primary Contact:

Angelo Babudro

1.348 Consultant ID: US, CO, MW Software

     P.O. Box 481762
     Denver, CO 80248-1762
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

303-832-5595 Company Fax:

303-832-3579 Support Phone Number:

303-832-5595 Contact Email Address:

mwinter@mwsoftware.com Company URL:

http://mwsoftware.com Primary Contact:

Mark Winter

1.349 Consultant ID: US, CO, PC WorkStations

     Not_Applicable
     Colorado Springs, CO Not_Applicable
     US

Consulting Specialties:

Custom Systems Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

719-330-3563 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

sales@pcworkstations.com Company URL:

http://www.pcworkstations.com Primary Contact:

lance williams

1.350 Consultant ID: US, CO, Paramin Consulting Group

     6424 South Quebec Street
     Englewood, CO 80111
     US

Consulting Specialties:

E-Commerce Main Distribution:

Other Other Distribution:

Slakware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not_Applicable Company Phone Number:

888 337 8227 Company Fax:

303 721 7522 Support Phone Number:

303 773 1361 Contact Email Address:

support@paramin.com Company URL:

http://www.paramin.com Primary Contact:

Amy Strauser

1.351 Consultant ID: US, CO, Paramin Consulting Group, Inc.

     6424 South Quebec Street
     Englewood, CO 80111
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Not Applicable Company Phone Number:

303 773 1361 Company Fax:

303 721 7522 Support Phone Number:

303 773 1361 Contact Email Address:

support@paramin.com Company URL:

http://www.paramin.com Primary Contact:

Amy Strauser

1.352 Consultant ID: US, CO, Travis Schack

     6105 Chase St
     Arvada, CO 80003
     US

Consulting Specialties:

Firewalls and Internet Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

303-456-2951 Company Fax:

303-414-6299 Support Phone Number:

303-456-2951 Contact Email Address:

tschack@uswest.net Company URL:

http://Not_Applicable Primary Contact:

Travis Schack

1.353 Consultant ID: US, CO, tummy.com, ltd.

5400 Fossil Court North
Fort Collins , CO 80525
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Own Dist. Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

970 223-8215 Company Fax:

408 490-2728 Support Phone Number:

Not Available Contact Email Address:

support@tummy.com Company URL:

http://www.tummy.com Primary Contact:

Evelyn Mitchell

1.354 Consultant ID: US, CT, DCA Online (Dutch Computer Association)

26 Tobacco Road
Weston, CT 06883-1612
US

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 203 222 8267 Company Fax:

+1 203 226 7037 Support Phone Number:

+1 203 222 8267 Contact Email Address:

support@dcaonline.com Company URL:

http://www.dcaonline.com Primary Contact:

Peter Leonard Krebs

1.355 Consultant ID: US, CT, DownCity, LLC

131 Broadway
Norwich, CT 06360
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-860-823-3000 Company Fax:

1-860-859-9401 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Robert Szarka

1.356 Consultant ID: US, CT, GNUTEC

1655 North Colony Road, Suite 3002
Meriden, CT 06450
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

All Distributions Specialty Distribution:

Other Other Specialty Distribution:

All Distributions Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-203-379-0125 Company Fax:

1-203-379-0126 Support Phone Number:

1-203-668-UNIX Contact Email Address:

support@gnutec.com Company URL:

http://http://www.gnutec.com/ Primary Contact:

Kyle Amon

1.357 Consultant ID: US, CT, Information Design, Inc.

     145 Durham Road, Suite 11
     Madison, CT 06443
     US

Consulting Specialties:

Custom Programming Internet Server Development Internet (Web) Programming Main Distribution:

Other Other Distribution:

All Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(203)245-0772 Company Fax:

(203)245-1885 Support Phone Number:

Not_Applicable Contact Email Address:

sales@idi-middleware.com Company URL:

http://www.idi-middleware.com Primary Contact:

Dennis Faust

1.358 Consultant ID: US, CT, MANHATTAN GROUP

     925 WHITE PLAINS ROAD
     TRUMBULL, CT 06611
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

203-445-0011 Company Fax:

203-445-0009 Support Phone Number:

203-445-0011 Contact Email Address:

SUPPORT@MANHATTANGROUP,COM Company URL:

http://WWW.MANHATTANGROUP.COM Primary Contact:

JOHN LOMBARDI

1.359 Consultant ID: US, CT, Peter Santoro Computing

     66 Farmbrook Lane
     South Windsor, CT 06074
     US

Consulting Specialties:

Custom Programming E-Commerce Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

Suse, Caldera, Slack Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

860 648-5737 Company Fax:

NA Support Phone Number:

860 648-5737 Contact Email Address:

peter@pscomp.com Company URL:

http://www.connix.com/˜ psantoro/ Primary Contact:

Peter Santoro

1.360 Consultant ID: US, CT, Tempest Harding, Inc.

     322 Main Street, Suite 1204
     Willimantic, CT 06226
     US

Consulting Specialties:

Custom Systems Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

860-456-3639 Company Fax:

860-456-2814 Support Phone Number:

860-456-3639 Contact Email Address:

info@thinc.net Company URL:

http://www.thinc.net Primary Contact:

Paul Herbstritt

1.361 Consultant ID: US, DC, BKP, Inc.

     P.O. Box 21026
     Washington, DC 20009
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Internet Server Development Netware Connectivity Network Administration System Administration

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(202) 332-5609 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

John Thomas Not_Applicable

1.362 Consultant ID: US, DC, Haisam K. Ido

     Not_Applicable
     Washington, DC 20015
     US

Consulting Specialties:

Internet (Web) Programming Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 202 537 5064 Company Fax:

Not_Applicable Support Phone Number:

1 202 537 5064 Contact Email Address:

idoh@cais.com Company URL:

http://Not_Applicable Primary Contact:

Haisam Ido

1.363 Consultant ID: US, DE, W. E. Hughes, INC

     1527 Vandyke Greenspring Road
     Smyrna, DE 19977
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security System Administration System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

wally@dca.net Company URL:

http://members.dca.net/wally Primary Contact:

Wally Hughes

1.364 Consultant ID: US, FL, Anthony Awtrey Consulting

781 Marsaille Dr Suite D
Indialantic, FL 32903
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

407-777-7164 Company Fax:

407-779-7269 Support Phone Number:

407-777-7164 Contact Email Address:

support@awtrey.com Company URL:

http://www.awtrey.com Primary Contact:

Anthony Awtrey

1.365 Consultant ID: US, FL, Bellstedt Custom Computers And Consulting

     3865 Corey Road
     Malabar, FL 32950-4118
     US

Consulting Specialties:

Custom Programming Custom Systems Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

001 407 729 0932 Company Fax:

001 407 956 8629 Support Phone Number:

Not_Applicable Contact Email Address:

olaf@bellstedt.com Company URL:

http://www.bellstedt.com Primary Contact:

Olaf Bellstedt

1.366 Consultant ID: US, FL, Byte Technologies, Inc.

     224 Datura St., Suite 808
     West Palm Beach, FL 33401
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

561-833-2330 Company Fax:

561-833-3135 Support Phone Number:

Not_Applicable Contact Email Address:

info@bytetechinc.com Company URL:

http://www.bytetechinc.com Primary Contact:

Adam Steinhoff

1.367 Consultant ID: US, FL, Inicom, Inc.

     6255 SW 69 Street
     South Miami, FL 33143
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Mandrake Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

305-439-9437 Company Fax:

305-665-1139 Support Phone Number:

305-439-9437 Contact Email Address:

aem@inicom.com Company URL:

http://www.inicom.com Primary Contact:

Andrew Mossberg

1.368 Consultant ID: US, FL, Linux Florida

6344 Raleigh St #1113
Orlando, FL 32835
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

407-296-0680 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

cshepher@linux-florida.com Company URL:

http://www.linux-florida.com Primary Contact:

Christopher Shepherd

1.369 Consultant ID: US, FL, Mackintosh Web, Inc.

3318 Clark Rd
Sarasota, FL 34231
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-941-316-5691 Company Fax:

1-941-921-5148 Support Phone Number:

1-941-921-5148 Contact Email Address:

terry@terrym.com Company URL:

http://www.terrym.com Primary Contact:

Terry Mackintosh

1.370 Consultant ID: US, FL, Richard J Kinch, PhD

6994 Pebble Beach Ct
Lake Worth, FL 33467
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 (561) 966-8400 Company Fax:

+1 (561) 966-0962 Support Phone Number:

+1 (561) 966-8400 Contact Email Address:

kinch@truetex.com Company URL:

http://www.truetex.com Primary Contact:

Richard Kinch

1.371 Consultant ID: US, FL, Spectrum Enterprises

     1609 Puritan Ave
     Winter Park, FL 32792
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

1-407-257-1437 Company Fax:

1-407-257-1437 Support Phone Number:

Not_Applicable Contact Email Address:

support@spectrumenterprises.com Company URL:

http://www.spectrumenterprises.com Primary Contact:

Paul Marble

1.372 Consultant ID: US, FL, TSM Services

108 Sunset Blvd
Boynton Beach, FL 33426
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

561-374-8874 Company Fax:

561-375-7877 Support Phone Number:

Not Available Contact Email Address:

support@tsmservices.com Company URL:

http://www.tsmservices.com Primary Contact:

Steve Grevemeyer

1.373 Consultant ID: US, FL, Timothy Jones

     PO Box 292371
     Temple Terrace, FL 33687
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

813-914-0653 Company Fax:

n/a Support Phone Number:

813-914-0653 Contact Email Address:

tim@worldfax.com Company URL:

http://www.timjones.com Primary Contact:

Timothy Jones

1.374 Consultant ID: US, FL, Yup Dot Com

     531 NE 2nd Street
     Gainesville, FL 32601
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

352/379-8921 Company Fax:

Not_Applicable Support Phone Number:

352/379-8921 Contact Email Address:

info@yup.com Company URL:

http://yup.com Primary Contact:

Daniel Butler

1.375 Consultant ID: US, FL, eBuilders Network

     8204 Crystal Clear Lane Suite 1200
     Orlando , FL 32809
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

Mandrake Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

407-857-6800 Company Fax:

407-857-2333 Support Phone Number:

407-857-6800 Contact Email Address:

sales@ebuildernet.com Company URL:

http://www.ebuildernet.com Primary Contact:

Pablo Bryan

1.376 Consultant ID: US, FL, psiCat Software, Inc.

     571-U Haverty Court
     Rockledge, FL 32955
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

407-632-1924 Company Fax:

407-632-1925 Support Phone Number:

Not_Applicable Contact Email Address:

support@psicat.com Company URL:

http://www.psicat.com Primary Contact:

Nathan Harris

1.377 Consultant ID: US, GA, Corvusware, Inc.

     PMB 138, 3245 Peachtree Pkwy
     Suwanee, GA 30024
     US

Consulting Specialties:

Custom Programming Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

All - we support LIN Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

770-886-7385 Company Fax:

770-886-7295 Support Phone Number:

770-886-7385 Contact Email Address:

sales@corvusware.com Company URL:

http://www.corvusware.com Primary Contact:

David Corbin

1.378 Consultant ID: US, GA, Fly-By-Day Consulting, Inc.

     4642 Bentley Place
     Duluth, GA 30096
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 770 662 8321 Company Fax:

Not_Applicable Support Phone Number:

1 770 662 8321 Contact Email Address:

bob@cavu.com Company URL:

http://www.cavu.com Primary Contact:

Bob Toxen

1.379 Consultant ID: US, GA, Hacom Corp

2477 Wrightsboro Road
Augusta, GA 30904
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

706-736-8717 Company Fax:

706-736-8717 Support Phone Number:

Not_Applicable Contact Email Address:

support@hacom.net Company URL:

http://www.hacom.net Primary Contact:

Bao Ha

1.380 Consultant ID: US, GA, John Meister

2350 Cobb Parkway, #34J
Smyrna, GA 30080
US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(770) 980-0665 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

meisterj@acm.org Company URL:

http://www.acm.org Primary Contact:

John Meister

1.381 Consultant ID: US, GA, PWR Group

     4330 North Smoke Ridge Court
     Roswell, GA 30075
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

770-650-8144 Company Fax:

Not_Applicable Support Phone Number:

1-770-650-8144 Contact Email Address:

prooker@pwrgroup.com Company URL:

http://www.pwrgroup.com Primary Contact:

Penn Rooker

1.382 Consultant ID: US, GA, Worldwind Technologies, Inc.

     8520 River Walk Landing
     Suwanee, GA 30024
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

Mandrake Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(770) 307-8590 Company Fax:

(770) 814-9967 Support Phone Number:

(770) 307-8590 Contact Email Address:

info@worldwindtech.com Company URL:

http://www.worldwindtech.com Primary Contact:

Chris Woodruff

1.383 Consultant ID: US, HI, Cytronix Computer Services, Inc.

     350 Ward Ave. 106-288
     Honolulu, HI 96814
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Netware Connectivity Virtual Private Networking

Main Distribution:

Other Other Distribution:

Custom In-House Dist Specialty Distribution:

Other Other Specialty Distribution:

Customer In-House Di Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(808) 383-0558 Company Fax:

Not_Applicable Support Phone Number:

(800) 939-9247 Contact Email Address:

deven@cytronix.net Company URL:

http://www.cytronix.net Primary Contact:

Deven Phillips

1.384 Consultant ID: US, HI, United Writers PC Support

     205-C1 Rose St
     Wahiawa, HI 96786
     US

Consulting Specialties:

Internet (Web) Programming Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Will look at all Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

808-622-0043 Company Fax:

808-622-1345 Support Phone Number:

808-622-0043 Contact Email Address:

unitedwriters.org@juno.com Company URL:

http://hawaiidakine.com (Linux Forum) Primary Contact:

AL PLANT

1.385 Consultant ID: US, IA, JEONET

PO Box 1282
Iowa City, IA 52241
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

DNS and Firewall con Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

319-338-6353 Company Fax:

319-338-6353 Support Phone Number:

800 number for clien Contact Email Address:

linuxsupport@jeo.net Company URL:

http://www.jeo.net Primary Contact:

Afan Ottenheimer

1.386 Consultant ID: US, IL, Cobler Corporation

     3013 S. Wolf Rd.
     Westchester, IL 60154
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

708-205-3313 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Kenneth Cobler

1.387 Consultant ID: US, IL, Dataline

105 South Wright St. Suite #309
Champaign, IL 61820
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

2173512647 Company Fax:

2173512647 Support Phone Number:

2173512647 Contact Email Address:

support@jay.cx Company URL:

http://Not_Applicable Primary Contact:

Jay Thomas

1.388 Consultant ID: US, IL, Ossmann Consulting

651 W. Belmont Ave.
Chicago, IL 60657
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

773-294-0250 Company Fax:

Not_Applicable Support Phone Number:

773-294-0250 Contact Email Address:

mike@ossmann.com Company URL:

http://www.ossmann.com Primary Contact:

Michael Ossmann

1.389 Consultant ID: US, IL, Promethan Consulting

18 Skyview Drive
Springfield, IL 62702
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Red Hat, Debian Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

217/698-6234 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Jeff Licquia

1.390 Consultant ID: US, IL, QLITech Linux Computers

     3717 43rd St. Suite #112
     Moline, IL 61265
     US

Consulting Specialties:

Custom Systems E-Commerce Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-877-24-LINUX Company Fax:

1-309-764-1142 Support Phone Number:

1-877-24-linux Contact Email Address:

support@qlitech.net Company URL:

http://www.qlitech.net Primary Contact:

Ray Sanders

1.391 Consultant ID: US, IL, Remington*Microshare

1834 Walden Office Square, Suite 100
Schaumburg, IL 60173
US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

847-221-0200 Company Fax:

847-221-0399 Support Phone Number:

847-221-0200 Contact Email Address:

chrisd@remingtonltd.com Company URL:

http://www.remingtonmicroshare.com Primary Contact:

Jeremy Simon

1.392 Consultant ID: US, IL, Urbana Computer Works

     701 Devonshire Dr. C-8
     Champaign, IL 61820
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

217-355-5936 Company Fax:

217-355-5953 Support Phone Number:

217-355-5936 Contact Email Address:

dpointer@ucworks.com Company URL:

http://www.ucworks.com Primary Contact:

David Pointer

1.393 Consultant ID: US, IL, WellThot Inc.

305 Van Damin Ave.
Glen Ellyn, IL 60137-5215
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

630-545-1470 Company Fax:

630-545-1471 Support Phone Number:

630-545-1470 Contact Email Address:

support@wellthot.com Company URL:

http://wellthot.com Primary Contact:

Mark Schwenk

1.394 Consultant ID: US, IL, onShore, Inc

     1407 W Chicago Av
     Chicago, IL 60622
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Virtual Private Networking

Main Distribution:

Debian Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Pacific HiTech HA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

RedHat Company Phone Number:

312 850-5200 Company Fax:

312 850-5208 Support Phone Number:

312 850-5200 Contact Email Address:

support-managers@onshore.com Company URL:

http://www.onshore.com Primary Contact:

Adam Di Carlo

1.395 Consultant ID: US, IN, CLFINC

P.O.Box 1271
Columbus, IN 47202-1271
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

812-372-1911 Company Fax:

Not_Applicable Support Phone Number:

812-372-1911 Contact Email Address:

cfiene@iquest.net Company URL:

http://www.iquest.net Primary Contact:

Curtis Fiene

1.396 Consultant ID: US, IN, IndyBox Systems, Inc.

PO Box 3564
Carmel, IN 46082-3564
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera RedHat Company Phone Number:

01 317 8469762 Company Fax:

Not_Applicable Support Phone Number:

+01 317 8469762 Contact Email Address:

support@indybox.com Company URL:

http://www.indybox.com Primary Contact:

Charles Martin

1.397 Consultant ID: US, IN, ValCom

     8098 Woodland Drive
     Indianapolis, IN 46278
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

(317)870-8020 Company Fax:

(317)870-8035 Support Phone Number:

(317)870-8020 Contact Email Address:

mreba@pccval.com Company URL:

http://www.pccval.com Primary Contact:

Matthew Reba

1.398 Consultant ID: US, KY, Computer Associates Network, Inc.

     218 South Porter Dr. Suite 2
     Richmond, KY 40475
     US

Consulting Specialties:

Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Mandrake Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-800-803-6706 Company Fax:

1-606-626-9559 Support Phone Number:

1-800-803-6706 Contact Email Address:

service@richmondsolutions.com Company URL:

http://www.richmondsolutions.com Primary Contact:

Jerry Hensley

1.399 Consultant ID: US, KY, WSI.com Consulting

     Woodchase Drive
     Erlanger, KY 41018
     US

Consulting Specialties:

Custom Programming Internet (Web) Programming System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(877) 506-6762 Company Fax:

(209) 882-9554 Support Phone Number:

(877) 506-6762 Contact Email Address:

support@wsicnslt.com Company URL:

http://www.wsicnslt.com Primary Contact:

Justin Rains

1.400 Consultant ID: US, LA, Advanced Network Technologies

     10795 Mead Rd. #104
     Baton Rouge, LA 70816
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

225-296-5815 Company Fax:

Not_Applicable Support Phone Number:

225-296-5815 Contact Email Address:

support@advanced-net-tech.com Company URL:

http://www.advanced-net-tech.com Primary Contact:

Benjamin Curtis

1.401 Consultant ID: US, LA, Connectivity Solutions, Inc.

365 Canal Street, Suite 2300
New Orleans, LA 70130
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

504-561-6540 Company Fax:

504-561-6694 Support Phone Number:

504-561-6540 Contact Email Address:

support@conn-sol.com Company URL:

http://www.conn-sol.com Primary Contact:

Kevin Wilson

1.402 Consultant ID: US, MA, Antarctica IT, Inc.

48 Simpson Drive
Framingham, MA 01701
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Caldera Company Phone Number:

1-877-365-4689 (1-87 Company Fax:

1-603-215-4709 Support Phone Number:

Not_Applicable Contact Email Address:

support@antarcticait.com Company URL:

http://www.antarcticait.com Primary Contact:

Scott Shaw

1.403 Consultant ID: US, MA, CYWAYS, Incorporated

19 Westchester Road
Newton, MA 02468
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+1 800 5-CYWAYS Company Fax:

+1 617 796-8997 Support Phone Number:

+1 617 796-8995 Contact Email Address:

support@cyways.com Company URL:

http://www.cyways.com Primary Contact:

Peter Lemieux

1.404 Consultant ID: US, MA, ComputeAmerica.com

     120C Broadmeadow Rd.
     Marlborough, MA 01752
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Intel Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

877-672-2171 Company Fax:

508-624-0120 Support Phone Number:

508-460-8897 Contact Email Address:

support@computeamerica.com Company URL:

http://www.computeamerica.com/WebDesign Primary Contact:

Sonny parlin Steve Doherty

1.405 Consultant ID: US, MA, DataProfit

     330 Whitney Ave
     Holyoke, MA 01040
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

800-437-7634 Company Fax:

413-536-3354 Support Phone Number:

NA Contact Email Address:

rcolumbia@dataprofit.com Company URL:

http://www.dataprofit.com Primary Contact:

Richard Columbia

1.406 Consultant ID: US, MA, LSC

205 West Granville RD
Tolland, MA 01034
US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

413-258-4074 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

c5141@iname.com Company URL:

http://www.iname.com Primary Contact:

Richard Columbia

1.407 Consultant ID: US, MA, NetInterface Consulting

56 Houghton
Hudson, MA 01749
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1.978.567.0972 Company Fax:

1.978.567.0972 Support Phone Number:

1.978.567.0972 Contact Email Address:

support@netinterface.com Company URL:

http://www.netinterface.com Primary Contact:

Lenny Forziati

1.408 Consultant ID: US, MA, New England Linux Resource

     Not_Applicable
     Newton, MA 02468
     US

Consulting Specialties:

E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

617 332 0387 Company Fax:

Not_Applicable Support Phone Number:

617 332 0837 Contact Email Address:

nelinuxresources@hotmail.com Company URL:

http://www.nelinuxresourses.com Primary Contact:

curt Not_Applicable

1.409 Consultant ID: US, MA, New England Linux resources

     Not_Applicable
     Newton, MA 02468
     US

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

caldera Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

617 332 0837 Company Fax:

Not_Applicable Support Phone Number:

617 332 0837 Contact Email Address:

nelinuxresources@hotmail.com Company URL:

http://www.nelinuxresources.com Primary Contact:

curt Not_Applicable

1.410 Consultant ID: US, MA, Worldmachine Technologies Corporation

44 Winter Street, 3rd Floor
Boston, MA 02108-4745
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

all Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(617) 357-4040 Company Fax:

(617) 357-4949 Support Phone Number:

(617) 357-4040 Contact Email Address:

support@worldmachine.com Company URL:

http://www.worldmachine.com Primary Contact:

Eric Hansen

1.411 Consultant ID: US, MD, JAMUX

4520 Yates Road
Beltsville, MD 20705-2629
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

1 301 595 2013 Company Fax:

Not_Applicable Support Phone Number:

1 301 595 2013 Contact Email Address:

jam@jamux.com Company URL:

http://www.jamux.com Primary Contact:

John Martin

1.412 Consultant ID: US, MD, Quest

10555 Metropolitan Avenue
Kensington, MD 20895-2627
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera, SUSE Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Caldera RedHat Company Phone Number:

800 876 4000 Company Fax:

301 942 0175 Support Phone Number:

800 876 4000 Contact Email Address:

superman@xrayvision.com Company URL:

http://qhq.com Primary Contact:

David Schwering

1.413 Consultant ID: US, MD, RABA Technologies, Inc.

10500 Little Patuxent Pkwy, Ste. 190
Columbia, MD 21044
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera RedHat Company Phone Number:

(410) 715-9399 Company Fax:

(410) 715-9397 Support Phone Number:

(410) 715-9399 Contact Email Address:

stuart.chandler@raba.com Company URL:

http://www.raba.com Primary Contact:

Stuart Chandler

1.414 Consultant ID: US, ME, TransWeb ToolBox

     PO Box 788
     Windham, ME 04062
     US

Consulting Specialties:

Not_Applicable Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Caldera Reseller Authorizations:

Not_Applicable Linux Certifications:

RedHat Company Phone Number:

207-450-7914 Company Fax:

207-892-8745 Support Phone Number:

207-450-7914 Contact Email Address:

support@transwebtools.com Company URL:

http://www.transwebtools.com Primary Contact:

Henri Brady

1.415 Consultant ID: US, MI, Jim Willette

     18190 - 15 mile Rd.
     Big Rapids, MI 49307
     US

Consulting Specialties:

CIFS (Samba) Main Distribution:

Other Other Distribution:

Any Specialty Distribution:

Other Other Specialty Distribution:

Any Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(616) 592 - 0405 Company Fax:

Not_Applicable Support Phone Number:

(616) 592-0405 Contact Email Address:

jimw@ctec.net Company URL:

http://www.ctec.net Primary Contact:

Jim Willette

1.416 Consultant ID: US, MI, Nugent Telecommunications

3081 Braeburn Circle
Ann Arbor, MI 48108
US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

734-971-1076 Company Fax:

734-971-4529 Support Phone Number:

734-971-1076 Contact Email Address:

jjn@nuge.com Company URL:

http://www.nuge.com Primary Contact:

Jay Nugent

1.417 Consultant ID: US, MI, Radiomind Studios

213 Schuring Rd
Portage, MI 49024-5012
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

616-323-9417 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@radiomind.com Company URL:

http://www.radiomind.com Primary Contact:

John Beimler

1.418 Consultant ID: US, MI, Struction Consulting

     3181 Baylis Dr.
     Ann Arbor, MI 48108-1709
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+1 (734) 975-0782 Company Fax:

+1 (917) 463-1256 Support Phone Number:

Not_Applicable Contact Email Address:

support@struction.com Company URL:

http://www.struction.com Primary Contact:

Brian Shensky

1.419 Consultant ID: US, MN, Anvi Corp

     2610 Grand St. NE
     Minneapolis, MN 55418
     US

Consulting Specialties:

Custom Programming Custom Systems Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Slackware Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(612) 782-1955 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://menon.com/ Primary Contact:

Kps Menon

1.420 Consultant ID: US, MN, Donnal Group Inc.

     19820 Cottagewood Road
     Excelsior, MN 55331
     US

Consulting Specialties:

Custom Systems Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera Company Phone Number:

612-474-7785 Company Fax:

612-474-7793 Support Phone Number:

888-617-0002 Contact Email Address:

support@donnalgroup.com Company URL:

http://www.donnalgroup.com Primary Contact:

Tim Burke

1.421 Consultant ID: US, MN, Geek Work

     634 Providence Ct
     Chaska, MN 55318
     US

Consulting Specialties:

Custom Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

612-368-0331 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

randym@acm.org Company URL:

http://www.hamline.edu/˜ rcmaas Primary Contact:

Randall Maas

1.422 Consultant ID: US, MO, Caviness Technologies, Inc.

4940 S. Peck Ave. Independence
Kansas City, MO 64055
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

816 478-7078 Company Fax:

816 350-7001 Support Phone Number:

816 478-7078 Contact Email Address:

linuxsupport@maxbaud.net Company URL:

http://www.maxbaud.net Primary Contact:

Chris Caviness

1.423 Consultant ID: US, MO, Caviness Technologies, Inc.

4940 S. Peck Ave. Independence
Kansas City, MO 64055
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

816 478-7078 Company Fax:

816 350-7001 Support Phone Number:

816 478-7078 Contact Email Address:

linuxsupport@maxbaud.net Company URL:

http://www.maxbaud.net Primary Contact:

Chris Caviness

1.424 Consultant ID: US, MO, Firefly Enterprises, Inc.

     922 Rochdale Dr . Suite 100
      Kirkwood , MO 63122
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

314-835-1590 Company Fax:

314-835-1591 Support Phone Number:

314-835-1590 Contact Email Address:

rsponholtz@earthlink.net Company URL:

http://Not_Applicable Primary Contact:

Ross Sponholtz

1.425 Consultant ID: US, MO, Omnitec Corporation

PO Box 411010
St. Louis, MO 63141
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

Debian/LinuxPro Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

3149974600 Company Fax:

3149970153 Support Phone Number:

3149974600 Contact Email Address:

admin@omnitec.net Company URL:

http://www.omnitec.net Primary Contact:

Lee Lammert

1.426 Consultant ID: US, MO, PhazeOne, Inc.

     3101 Broadway, Suite 605
     Kansas City, MO 64111
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(816)756-1300 Company Fax:

(816)756-1355 Support Phone Number:

(816)756-1300 Contact Email Address:

support@phazeone.com Company URL:

http://www.phazeone.com Primary Contact:

Steven Heutinck

1.427 Consultant ID: US, MS, BCCS

     307 Clear View dr
     Florence, MS 39073
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

601-845-3191 Company Fax:

NA Support Phone Number:

NA Contact Email Address:

rbelk@netdoor.com Company URL:

http://www2.netdoor.com/˜ rbelk Primary Contact:

Randy Belk

1.428 Consultant ID: US, NC, ApeX Systems Integration Corp.

     PO Box 338
     Apex, NC 27502
     US

Consulting Specialties:

Custom Systems Firewalls and Internet Security Network Administration System Administration Main Distribution:

Suse Other Distribution:

Red Hat, and others Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

919-468-8150 Company Fax:

919-468-5288 Support Phone Number:

NA Contact Email Address:

apex@2boot.com Company URL:

http://www.2boot.com Primary Contact:

Joe Ferguson

1.429 Consultant ID: US, NC, Certified Services

     4911 Benthaven Lane
     Charlotte, NC 28269
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

704-509-4585 Company Fax:

Not_Applicable Support Phone Number:

704-509-4585 Contact Email Address:

support@certifiedservices.com Company URL:

http://wwww.certifiedservices.com Primary Contact:

Kurt Brust

1.430 Consultant ID: US, NC, Computer Network Consulting Inc.

102 Shepton Drive
Morrisville, NC 27560
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 919 380 9687 Company Fax:

+1 919 380 0409 Support Phone Number:

+1 919 380 9687 Contact Email Address:

Support@CNConsulting.com Company URL:

http://www.CNConsulting.com Primary Contact:

Tuananh Nguyen

1.431 Consultant ID: US, NC, Jim Bray

     (not posted on web)
     Brevard, NC 28712
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Debian Other Distribution:

Any Specialty Distribution:

Other Other Specialty Distribution:

Any Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

828-862-5407 Company Fax:

NA Support Phone Number:

NA Contact Email Address:

jb@as220.org Company URL:

http://as220.org/jb/freeware Primary Contact:

Jim Bray

1.432 Consultant ID: US, NC, Linux Workshop

105 B. Milton Dr.
Carrboro, NC 27514
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Not_Applicable Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Not Applicable Company Phone Number:

919-960-8030 Company Fax:

919-942-0793 Support Phone Number:

919-960-8030 Contact Email Address:

support@linux-workshop.com Company URL:

http://www.linux-workshop.com Primary Contact:

Not_Applicable Not_Applicable

1.433 Consultant ID: US, NC, Moongroup Consulting, Inc.

     134 Grande Meadow Way
     Cary, NC 27513
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

01-919-462-9426 Company Fax:

01-919-462-9427 Support Phone Number:

01-919-462-9426 Contact Email Address:

support@moongroup.com Company URL:

http://www.moongroup.com Primary Contact:

Chuck Mead

1.434 Consultant ID: US, NC, Network Computing Solutions

3733 National Dr.
Raleigh, NC 27612
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-919-510-6970 Company Fax:

1-919-510-6971 Support Phone Number:

1-919-510-6970 Contact Email Address:

linux-support@networks.com Company URL:

http://www.networks.com Primary Contact:

Pat Hughes

1.435 Consultant ID: US, NC, UTSL dot com

     1007 North Buchanan Blvd.
     Durham, NC 27701-1139
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Debian Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

919-286-1441 Company Fax:

Not_Applicable Support Phone Number:

919-286-1441 Contact Email Address:

biggers@utsl.com Company URL:

http://www.utsl.com Primary Contact:

Mark Biggers

1.436 Consultant ID: US, NC, Whole Systems Solutions, Inc.

     2 Terrace Way, Suite A
     Greensboro, NC 27403
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Any Distribution Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 336-297-4977 Company Fax:

+1 336-297-9787 Support Phone Number:

+1 336-297-4977 Contact Email Address:

support@wss.net Company URL:

http://www.wss.net Primary Contact:

Jay Eisenberg

1.437 Consultant ID: US, NC, World Wide Wizdom

104 Quartermaine Ct.
Cary, NC 27513-5152
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

919-818-1855 Company Fax:

919-468-4684 Support Phone Number:

919-468-4680 Contact Email Address:

support@wiZdom.net Company URL:

http://www.wiZdom.net Primary Contact:

David Boone

1.438 Consultant ID: US, NE, Paktronix Systems LLC

     1506 North 59th Street
     Omaha, NE 68104-4830
     US

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Trinux, Pretty Secur Reseller Authorizations:

IBM Linux Certifications:

Not_Applicable Company Phone Number:

1-402-932-7250 Company Fax:

1-402-932-7258 Support Phone Number:

1-402-932-7250 Contact Email Address:

tech@paktronix.com Company URL:

http://www.paktronix.com Primary Contact:

Paul Greunke

1.439 Consultant ID: US, NH, AKCP, Inc

379 Amherst Street, Suite 182
Nashua, NH 03063
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

603-881-7343 Company Fax:

Not_Applicable Support Phone Number:

603-881-7343 Contact Email Address:

info@kcpinc.com Company URL:

http://www.kcpinc.com Primary Contact:

Brad Klein

1.440 Consultant ID: US, NH, KORKSOFT

8 Olde Bedford Way
Bedford, NH 03110
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

603 472 8262 Company Fax:

603 472 8262 Support Phone Number:

603 472 8262 Contact Email Address:

support@korksoft.com Company URL:

http://www.korksoft.com Primary Contact:

Jason Korkin

1.441 Consultant ID: US, NJ, Etc Services

     31 N. Vivyen St.
     Bergenfield, NJ 07621-1529
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

201-385-7113 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Scott Webster

1.442 Consultant ID: US, NJ, Mitchell C. Nelson, Ph.D.

     Not_Applicable
     Morristown, NJ 07960
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Any and All Specialty Distribution:

Other Other Specialty Distribution:

Any and All Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(973) 644-3577 Company Fax:

please phone first Support Phone Number:

(973) 644-3577 Contact Email Address:

mcnelson@mindspring.com Company URL:

http://Not_Applicable Primary Contact:

Mitch Nelson

1.443 Consultant ID: US, NJ, Voortek Software, Inc

     1 Las Brisas Blvd
     Voorhees, NJ 08043
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Caldera Company Phone Number:

David J Adamson Company Fax:

609-346-3172 Support Phone Number:

888-346-4616 Contact Email Address:

voorteksoftware@erols.com Company URL:

http://www.erols.com Primary Contact:

David J Adamson

1.444 Consultant ID: US, NON United States, ASA Computers, Inc.

     23
     Not_Applicable, NON United States Not_Applicable
     US

Consulting Specialties:

Not_Applicable Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.445 Consultant ID: US, NON United States, Antonio Marcelo F. Fonseca

     Rua Haddock Lobo 131 - c7 -  Rio Comprido
     Rio de Janeiro, NON United States 20260-130
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

55-21-2932523 Company Fax:

55-21-2932523 Support Phone Number:

Not_Applicable Contact Email Address:

amarcelo@iis.com.br Company URL:

http://Not_Applicable Primary Contact:

Antonio Marcelo

1.446 Consultant ID: US, NON United States, Aztronic Enterprises

     Not_Applicable
     San , NON United States Not_Applicable
     US

Consulting Specialties:

Not_Applicable Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.447 Consultant ID: US, NON United States, Coldweb Consulting

     2200 A Montclair Rd
     Greensboro, NON United States 27410
     US

Consulting Specialties:

E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Caldera Other Distribution:

Redhat and Other Specialty Distribution:

Other Other Specialty Distribution:

Turbo Linux Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

336-323-0253 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.448 Consultant ID: US, NON United States, CyberMarketplace

     1106 Coolidge Blvd
     Lafayette, NON United States LA
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Cobalt Intranet Serv Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(318) 232-5759 Company Fax:

(318) 234-7376 Support Phone Number:

support@cybermp.net Contact Email Address:

support@cybermp.net Company URL:

http://www.cybermp.net Primary Contact:

Theron Lutz

1.449 Consultant ID: US, NON United States, Developpement Strategique Sophos inc.

1305 Monette
Montreal, NON United States J7K 3C1
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(450) 966-9676 Company Fax:

(450) 474-7161 Support Phone Number:

Not_Applicable Contact Email Address:

info@Sophos.ca Company URL:

http://Sophos.ca Primary Contact:

Ludovic Marcotte

1.450 Consultant ID: US, NON United States, Emit Sp z o.o.

ul. Koszarowa 48/9
Wroclaw, NON United States 51-149
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Custom Specialty Distribution:

Other Other Specialty Distribution:

Single floppy Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+48713267026 Company Fax:

+48713289733 Support Phone Number:

+48713288967 Contact Email Address:

kontakt@emit.pl Company URL:

http://www.emit.pl Primary Contact:

Ian Carr-de Avelon

1.451 Consultant ID: US, NON United States, Gatekeeper Technology, Ltd.

     26 All Saints' Avenue
     COLCHESTER, NON United States CO3 4NZ
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 7050 368852 Company Fax:

Not_Applicable Support Phone Number:

+44 7050 368852 Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Tom Yates

1.452 Consultant ID: US, NON United States, Jobsoft Design and Development, Inc.

118 South Maple Street
Murfreesboro, NON United States 37130
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

6159049559 Company Fax:

6158908941 Support Phone Number:

Not_Applicable Contact Email Address:

info@jobsoft.com Company URL:

http://www.jobsoft.com Primary Contact:

Andy Fort

1.453 Consultant ID: US, NON United States, Magiclink

Av. ACM, Sala 304 Edf. Maxcenter - Pituba
Salvador - BA, NON United States 41858-900
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+55-71-351-1615 Company Fax:

+55-71-353-2220 Support Phone Number:

+55-71-351-1615 Contact Email Address:

admin@magiclink.com.br Company URL:

http://www.magiclink.com.br Primary Contact:

Isamar Maia

1.454 Consultant ID: US, NON United States, Not_Applicable

     Not_Applicable
     Not_Applicable, NON United States Not_Applicable
     US

Consulting Specialties:

Not_Applicable Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.455 Consultant ID: US, NON United States, Onlinet Com. Serv. Informática Ltda.

     R. Pará, 35 - Cid. Industrial
     Lorena, NON United States 12.600-000
     US

Consulting Specialties:

Internet Server Development Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+55 (012) 553-1234 Company Fax:

+55 (012) 553-2568 Support Phone Number:

+55 (012) 553-1234 Contact Email Address:

suporte@onlinet.com.br Company URL:

http://www.onlinet.com.br Primary Contact:

Denys Sene

1.456 Consultant ID: US, NON United States, Psand Limited

     Not_Applicable
     Not_Applicable, NON United States Not_Applicable
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

RedHat Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://www.psand.net Primary Contact:

Mike Harris

1.457 Consultant ID: US, NON United States, Tksoft, OY, Inc.

     Sepantie 10
     Porvoo, NON United States 06400
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet (Web) Programming System Security

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+358-19-575-4081 Company Fax:

+358-19-575-4133 Support Phone Number:

Not_Applicable Contact Email Address:

admin@tksoft.com Company URL:

http://www.tksoft.com Primary Contact:

Troy Korjuslommi

1.458 Consultant ID: US, NON United States, Venkat

     1,1810,river court
     jersey, NON United States 07310
     US

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

avsrao_india@yahoo.com Company URL:

http://avsrao.8m.com Primary Contact:

venkat rao

1.459 Consultant ID: US, NON United States, iOpen Technologies Ltd.

PO Box 13-917
Christchurch, NON United States 8004
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+64 21 636 661 Company Fax:

Not_Applicable Support Phone Number:

+64 21 636 661 Contact Email Address:

support@iopen.co.nz Company URL:

http://www.iopen.co.nz Primary Contact:

Richard Waid

1.460 Consultant ID: US, NON United States, spinalware technologies

     33,st.pauls ave
     jersey, NON United States 07310
     US

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

201-626-3184 Company Fax:

n/a Support Phone Number:

201-626-3184 Contact Email Address:

avsrao_india@yahoo.com Company URL:

http://avsrao.8m.com Primary Contact:

venkat rao

1.461 Consultant ID: US, NY, Aegis Information Systems, Inc.

     P.O. Box 730
     Hicksville, NY 11802
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

various Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

800-234-4700 Company Fax:

516-937-3999 Support Phone Number:

800-344-8165 Contact Email Address:

help@AegisInfoSys.com. Company URL:

http://www.AegisInfoSys.com. Primary Contact:

David Harari

1.462 Consultant ID: US, NY, Andrew Madloch

     152 Madison Avenue, Suite 1104
     New York, NY 10016
     US

Consulting Specialties:

Custom Programming E-Commerce Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

+1 212 214 0558 Company Fax:

+1 212 214 0558 Support Phone Number:

+1 212 214 0558 Contact Email Address:

support@intermarket3d.com Company URL:

http://www.intermarket3d.com Primary Contact:

Andrew Madloch

1.463 Consultant ID: US, NY, Donahue Consulting

100 John St., #1410
New York, NY 10038
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

212.998.3158 Company Fax:

212.995.4120 Support Phone Number:

212.998.3158 Contact Email Address:

adam@donahueinc.net Company URL:

http://www.donahueinc.net Primary Contact:

Adam Donahue

1.464 Consultant ID: US, NY, Instrumental Software Technologies, Inc.

     70 CEREUS WAY
     NEW PALTZ , NY 12561
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

914 256-9290 Company Fax:

914 256-9299 Support Phone Number:

Not_Applicable Contact Email Address:

info@isti.com Company URL:

http://www.isti.com Primary Contact:

Paul Friberg

1.465 Consultant ID: US, NY, Internet Server COnnections, Inc.

     3616 Edgehill Road
     Yorktown Heights, NY 10598-1104
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Intel Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1-914-243-5675 Company Fax:

+1-914-243-4965 Support Phone Number:

+1-914-243-5675 Contact Email Address:

support@internetserver.com Company URL:

http://www.internetserver.com Primary Contact:

Homayoon Beigi

1.466 Consultant ID: US, NY, JAP Consulting

     82 Beach 2nd Street
     Lawrence, NY 11559
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration

Main Distribution:

Suse Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-516-371-6886 Company Fax:

1-516-371-6886 Support Phone Number:

1-516-371-6886 Contact Email Address:

support@japconsulting.com Company URL:

http://www.japconsulting.com Primary Contact:

Jason Pearsall

1.467 Consultant ID: US, NY, Logicept Corporation

41 Chenango St, Suite 201
Binghamton, NY 13901-2901
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

info@logicept.com Company URL:

http://www.logicept.com Primary Contact:

Jaron Rubenstein

1.468 Consultant ID: US, NY, Lrw.Net

     188 2nd Ave, #9
     New York, NY 10003
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Suse Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-212-254-3551 Company Fax:

1-212-254-6175 Support Phone Number:

1-212-254-3551 Contact Email Address:

rw26@lrw.net Company URL:

http://lrw.net Primary Contact:

Randy Wright

1.469 Consultant ID: US, NY, Paul Dlug

     6 Alilah Ct.
     Miller Place, NY 11764
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

pdlug@portjeff.net Company URL:

http://Not_Applicable Primary Contact:

Paul Dlug

1.470 Consultant ID: US, NY, Realtime Software Solutions, Inc.

     44 Northfield Road
     Rcchester, NY 14617
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Internet Server Development Internet (Web) Programming Network Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-716-467-8889 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

tom@realtime1.com Company URL:

http://realtime1.com Primary Contact:

Thomas Bond

1.471 Consultant ID: US, NY, RetiComp, Inc.

     75 West End Ave. Suite R32C
     New York, NY 10023
     US

Consulting Specialties:

Firewalls and Internet Security Internet Server Development Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

212-957-6710 Company Fax:

212-957-6711 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://www.reticomp.com Primary Contact:

Donn Morrill

1.472 Consultant ID: US, NY, SuperGeek Consolidated

80 Miller Rd.
Latham, NY 12110
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Red Hat & Debian Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1+ 518-786-1594 Company Fax:

1+ 518-786-1594 Support Phone Number:

Not_Applicable Contact Email Address:

supergeek@supergeek.com Company URL:

http://www.supergeek.com Primary Contact:

Ed Huott

1.473 Consultant ID: US, NY, Templar Studios, Inc.

157 Ludlow St.
New York, NY 10002
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Or any other... Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

212-982-9360 Company Fax:

212-982-9370 Support Phone Number:

212-982-9360 Contact Email Address:

dwood@templar.com Company URL:

http://www.templar.com Primary Contact:

David Wood

1.474 Consultant ID: US, NY, Zenez, Inc.

     40 Paine Street
     Lindenhurst, NY 11757
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-516-957-1907 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

ben@amvalue.com Company URL:

http://Not_Applicable Primary Contact:

Benjamin Suto

1.475 Consultant ID: US, OH, 800Linux

     1357 West Lane Ave suite 211
     Columbus, OH 43221
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

614-324-1610 Company Fax:

614-324-1617 Support Phone Number:

614-324-1610 Contact Email Address:

suppor Company URL:

http://www.linux800.com Primary Contact:

Not_Applicable Not_Applicable

1.476 Consultant ID: US, OH, Advanced Unix and Network Consulting

     7652 Crimson Ct.
     Mentor, OH 44060
     US

Consulting Specialties:

Firewalls and Internet Security Network Administration System Administration System Security Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(440) 255-0421 Company Fax:

(440) 255-0421 Support Phone Number:

Not_Applicable Contact Email Address:

Dennis.Dumont@AUNC.com Company URL:

http://www.aunc.com Primary Contact:

Dennis Dumont

1.477 Consultant ID: US, OH, Bradley M. Kuhn

     P.O. Box 20042
     Cincinnati, OH 45220
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

Red Hat Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

513-475-0311 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

bkuhn@acm.org Company URL:

http://www.ebb.org/bkuhn Primary Contact:

Bradley Kuhn

1.478 Consultant ID: US, OH, CinLAN Systems

     1054 Kreis Lane
     Cincinnati, OH 45205
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Network Administration System Administration Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

513-471-7272 Company Fax:

Not_Applicable Support Phone Number:

513-471-7272 Contact Email Address:

crdorr@briefcase.com Company URL:

http://members.xoom.com/CinLAN Primary Contact:

Christopher Dorr

1.479 Consultant ID: US, OH, Cleveland Internet Association

     4876 East 85th
     Cleveland, OH 44125
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Will consider any Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

216-552-4368 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Jeff Garvas

1.480 Consultant ID: US, OH, Computer Wizards Consulting

     148 Apricot Ln
     Dayton, OH 45433-1109
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

937-754-0458 Company Fax:

937-754-0194 Support Phone Number:

937-754-0458 Contact Email Address:

support@computerwizards.com Company URL:

http://www.computerwizards.com Primary Contact:

Jeff Lamb

1.481 Consultant ID: US, OH, Essex Software Development, Inc.

     3452 Meadowbrook Blvd.
     Cleveland Heights, OH 44118
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

216-932-1230 Company Fax:

216-397-3380 Support Phone Number:

216-932-1230 Contact Email Address:

Support@esdi.com Company URL:

http://http://www.esdi.com Primary Contact:

Brendan O’Haire

1.482 Consultant ID: US, OH, LANshark Systems, Inc

784 Morrison Road
Columbus, OH 43230
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

All others Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

614-751-1111 Company Fax:

614-751-1112 Support Phone Number:

614-751-1111 Contact Email Address:

support@lanshark.com Company URL:

http://www.lanshark.com Primary Contact:

Scott Sharkey

1.483 Consultant ID: US, OH, Lightstream Internet, Inc.

     7700 Saint Clair
     Mentor, OH 44060
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(440)/269-2352 Company Fax:

(440)/942-6387 Support Phone Number:

(440)/269-2352 Contact Email Address:

tech@lightstream.net Company URL:

http://www.lightstream.net Primary Contact:

Jon Changnon

1.484 Consultant ID: US, OH, Saraber Solutions

     110 Diehl st.
     Bettsville, OH 44815
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

Any, Redhat, Debian, Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(419)986-5028 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@saraber.dhs.org Company URL:

http://saraber.dhs.org Primary Contact:

Gerard Saraber

1.485 Consultant ID: US, OH, TES, Inc.

     2903 Ravogli Ave.
     Cincinnati, OH 45211
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 513 661 3200 Company Fax:

+1 513 661 3732 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Larry Townsend

1.486 Consultant ID: US, OH, The Computer Underground

1357 West Lane Ave suite 210
Columbus, OH 43221
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not_Applicable Company Phone Number:

614-485-0506 Company Fax:

Not_Applicable Support Phone Number:

614-485-0506 Contact Email Address:

support@tcu-inc.com Company URL:

http://http://www.tcu-inc.com Primary Contact:

Mark Nielsen

1.487 Consultant ID: US, OH, The Computer Underground, Inc.

1357 West Lane Ave suite 210
Columbus, OH 43221
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

614-485-0506 Company Fax:

614-485-9617 Support Phone Number:

614-485-0506 Contact Email Address:

suport@tcu-inc.com Company URL:

http://www.tcu-inc.com Primary Contact:

Mark Nielsen

1.488 Consultant ID: US, OH, Unique Systems, Inc.

     6920 Spring Valley Dr, #106
     Holland, OH 43528
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Internet (Web) Programming System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Mandrake Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Not Applicable Company Phone Number:

419 861-3331 Company Fax:

419 861-3340 Support Phone Number:

419 861-3331 Contact Email Address:

info@uniqsys.com Company URL:

http://www.uniqsys.com Primary Contact:

Glenn Jacobson

1.489 Consultant ID: US, OH, dkFlyer

3962 East 55th St
Newburgh Hts, OH 44125-4864
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

All Specialty Distribution:

Other Other Specialty Distribution:

WiniX Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(216)883-0222 Company Fax:

(216)883-0221 Support Phone Number:

Not_Applicable Contact Email Address:

dkflyer@dkflyer.com Company URL:

http://www.dkflyer.com Primary Contact:

Dennis Kennedy

1.490 Consultant ID: US, OK, After Hours Computer Consulting,LLC.

     2 West 6th Street, Suite 450
     Tulsa, OK 74119
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

918.583.1004 Company Fax:

918.587.2422 Support Phone Number:

918.583.1004 Contact Email Address:

support@ahcctulsa.com Company URL:

http://www.ahcctulsa.com Primary Contact:

Kenneth Hess

1.491 Consultant ID: US, OR, 3D Technology, Inc.

     9150 SW Pioneer Ct. Suite E
     Portland, OR 97070
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet (Web) Programming Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

RedHat Company Phone Number:

877-333-2677 Company Fax:

503-582-9102 Support Phone Number:

503-582-1190 Contact Email Address:

support@3dtech.net Company URL:

http://www.3dtech.net Primary Contact:

Kerry Lafferty

1.492 Consultant ID: US, OR, Avalon Technology Group, Inc.

0607 SW Idaho St.
Portland, OR 97201
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

(503) 246-3630 Company Fax:

(503) 246-3124 Support Phone Number:

(503) 246-3630 Contact Email Address:

support@avalongroup.net Company URL:

http://www.avalongroup.net Primary Contact:

Timothy Grant

1.493 Consultant ID: US, OR, Fargus.Net

     326 sw 7th street
     Corvallis, OR 97333
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

1 541 754 6231 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

nick@fargus.net Company URL:

http://www.fargus.net Primary Contact:

Nick Rusnov

1.494 Consultant ID: US, OR, Fence-Post Technology

     5117 NE 31st Ave
     Portland, OR 97211
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

503-287-5244 Company Fax:

503-335-9306 Support Phone Number:

503-287-5244 Contact Email Address:

info@fence-post.com Company URL:

http://www.fence-post.com Primary Contact:

Art Opiela-Young

1.495 Consultant ID: US, OR, Hurrah Internet Services

     28 SW 1st Ave.
     Portland, OR 97204
     US

Consulting Specialties:

Custom Programming E-Commerce Internet Server Development Internet (Web) Programming System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-503-380-3218 Company Fax:

1-503-223-3140 Support Phone Number:

1-503-380-3218 Contact Email Address:

support@hurrah.com Company URL:

http://www.hurrah.com Primary Contact:

Tracy Camp David Elkins

1.496 Consultant ID: US, OR, Naked Ape Consulting

4404 SE Nehalem St.
Portland, OR 97206
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 503 788 4688 Company Fax:

+1 503 788 4688 Support Phone Number:

+1 503 788 4688 Contact Email Address:

wcooley@nakedape.navi.net Company URL:

http://nakedape.navi.net Primary Contact:

W. Reilly Cooley

1.497 Consultant ID: US, PA, 23rd Century Computing

     271 Austin Drive
     Grantville, PA 17028
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Slackware Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(717) 469-0614 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@23cc.com Company URL:

http://Not_Applicable Primary Contact:

Brian Reifsnyder

1.498 Consultant ID: US, PA, Associated Micro Systems, Inc.

     140 South Village Avenue Suite 20
     Exton, PA 19341
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

610-524-9816 Company Fax:

610-524-9821 Support Phone Number:

610-524-9816 Contact Email Address:

mmarcini@amicrosys.com Company URL:

http://www.amicrosys.com Primary Contact:

Mike Marciniszyn

1.499 Consultant ID: US, PA, Azaroff Consulting

2 Brandywine
Hershey, PA 17033
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Caldera Other Distribution:

Red Hat Specialty Distribution:

Other Other Specialty Distribution:

Extreme Linux Reseller Authorizations:

IBM Linux Certifications:

Caldera Company Phone Number:

7175200423 Company Fax:

none Support Phone Number:

7175200423 Contact Email Address:

aazaroff@redrose.net Company URL:

http://none Primary Contact:

Andre Azaroff

1.500 Consultant ID: US, PA, IQ Group

     325 Chestnut Street
     Philadelphia, PA 19106
     US

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Linux Certifications:

Not Applicable Company Phone Number:

215-925-4500 Company Fax:

215-925-1970 Support Phone Number:

Not_Applicable Contact Email Address:

mcdonc@iqgroup.com Company URL:

http://www.iqgroup.com Primary Contact:

Chris McDonough

1.501 Consultant ID: US, PA, JAV-Associates Consulting

     239 South West Street
     York, PA 17404
     US

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

7178540425 Company Fax:

3058322662 Support Phone Number:

7178540425 Contact Email Address:

tech@src-router.tytek.net Company URL:

http://src-router.tytek.net Primary Contact:

John Vorchak

1.502 Consultant ID: US, PA, Megaglobal Corp.

922 South 46th Street
Philadelphia, PA 19143
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

215-681-8316 Company Fax:

Not_Applicable Support Phone Number:

215-681-8316 Contact Email Address:

support@megaglobal.com Company URL:

http://megaglobal.com Primary Contact:

andrew jastremski

1.503 Consultant ID: US, PA, Michael Jastremski

922 South 46ths
Philadelphia, PA 19143
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Any Specialty Distribution:

Other Other Specialty Distribution:

Our Own Distro :) Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

215-681-8316 Company Fax:

215-555-1212 Support Phone Number:

215-681-8316 Contact Email Address:

support@westphila.net Company URL:

http://http://westphila.net Primary Contact:

Michael Jastremski

1.504 Consultant ID: US, PA, Outcome Technology Associates, Inc.

612 California Avenue
Pittsburgh, PA 15202
US

Consulting Specialties:

Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not Applicable Company Phone Number:

412.766.0300 Company Fax:

412.766.8245 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.otai.net Primary Contact:

Sean McLinden

1.505 Consultant ID: US, PA, Reed Consulting

     26 courtney st
     pittsburgh, PA 15202
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

412-766-2373 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Joel Reed

1.506 Consultant ID: US, PA, Rich Westcott

     4053 Boone St
     Phila, PA 19127
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

215-487-7529 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

rwestcot@ctlinux.ecogchair.org Company URL:

http://www.ctlinux.ecogchair.org Primary Contact:

Rich Westcott

1.507 Consultant ID: US, PR, OG Consulting

     Urb. Santa Clara; W-1 Anamu
     Guaynabo, PR 00969
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

787-731-4332 Company Fax:

787-731-4331 Support Phone Number:

787-731-4332 Contact Email Address:

support@og-consulting.com Company URL:

http://www.og-consulting.com Primary Contact:

James O’Malley

1.508 Consultant ID: US, RI, New England Computer

     847 Mendon Rd.
     Woonsocket, RI 02895
     US

Consulting Specialties:

CIFS (Samba) Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

4017629980 Company Fax:

4017650031 Support Phone Number:

4017629980 Contact Email Address:

jaxon6@home.com Company URL:

http://www.home.com Primary Contact:

Patrick Paul

1.509 Consultant ID: US, SC, Crux

729 Martin Raod
Starr, SC 29684
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

864-375-3077 Company Fax:

Not_Applicable Support Phone Number:

864-375-0377 Contact Email Address:

carlk@carol.net Company URL:

http://www.carol.net Primary Contact:

Carl A. Kolasky Not_Applicable

1.510 Consultant ID: US, SC, GAME.NET

     1209 Midvale Avenue
     Charleston, SC 29412
     US

Consulting Specialties:

Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

Other Other Distribution:

slackware Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-843-762-7131 Company Fax:

1-843-762-7131 Support Phone Number:

1-843-762-7131 Contact Email Address:

info@game.net Company URL:

http://game.net Primary Contact:

Jim Sims

1.511 Consultant ID: US, SC, GAME.NET, Inc.

1209 Midvale Avenue
Charleston, SC 29412
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

SlackWare Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1-843-762-7131 Company Fax:

fax is for losers - Support Phone Number:

+1-843-762-7131 Contact Email Address:

support@game.net Company URL:

http://www.game.net Primary Contact:

Rick Bauman

1.512 Consultant ID: US, SC, Linux Empower

     Not_Applicable
     Sumter, SC 29040
     US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

consulting@linuxempower.com Company URL:

http://www.linuxempower.com Primary Contact:

Rob Smith

1.513 Consultant ID: US, SC, Michael W. Shaffer

P.O. Box 12215
Columbia, SC 29211
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

803-252-8079 Company Fax:

Not_Applicable Support Phone Number:

803-252-8079 Contact Email Address:

mwshaffer@earthlink.net Company URL:

http://Not_Applicable Primary Contact:

Michael Shaffer

1.514 Consultant ID: US, SC, PenguinXpress

1752 Potter Road
Lancaster, SC 29720
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-803-285-1350 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

rclong@hubcap.clemson.edu Company URL:

http://www.penguinxpress.net Primary Contact:

Robert Long

1.515 Consultant ID: US, SC, eDivision.net

     PMB# 274; 33 Office Park Rd 4A
     Hilton Head Island, SC 29928
     US

Consulting Specialties:

E-Commerce Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

770.216.1516 Company Fax:

770.216.1516 Support Phone Number:

Not_Applicable Contact Email Address:

support@eDivision.net Company URL:

http://www.eDivision.net Primary Contact:

ski kinsey

1.516 Consultant ID: US, TN, Michael Chaney Consulting Corporation

     2624 Oak Forest Drive
     Antioch, TN 37013
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 615 361 3254 Company Fax:

+1 615 361 1244 Support Phone Number:

+1 615 361 3254 Contact Email Address:

mdchaney@michaelchaney.com Company URL:

http://www.michaelchaney.com/serious Primary Contact:

Michael Chaney

1.517 Consultant ID: US, TN, Sheer Genius

2016 Rosecliff Dr.
Nashville, TN 37206
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 615.226.4040 Company Fax:

+1 615.228.9900 Support Phone Number:

+1 615.364.4040 Contact Email Address:

consulting@sheergenius.com Company URL:

http://www.sheergenius.com Primary Contact:

Matt Kenigson

1.518 Consultant ID: US, TN, TeleVideo Global, Inc.

     448 N. Cedar Bluff Rd. #350
     Knoxville, TN 37923
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Mandrake-RedHat Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

423-777-0080 Company Fax:

423-777-0098 Support Phone Number:

423-777-0080 Contact Email Address:

televid@televid.com Company URL:

http://linux.televid.com Primary Contact:

Stephen Nodvin

1.519 Consultant ID: US, TX, Aegis Data Systems

PO Box 28554
Dallas, TX 75228
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

214.394.4213 Company Fax:

972.857.9909 Support Phone Number:

214.394.4213 Contact Email Address:

operations@aegisdata.com Company URL:

http://www.aegisdata.com Primary Contact:

Mark Stingley

1.520 Consultant ID: US, TX, Computer Innovations International

2145 NW Military Highway, Suite 103
San Antonio, TX 78213
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Caldera Specialty Distribution:

Other Other Specialty Distribution:

Intel/Sparc/Alpha Reseller Authorizations:

Compaq HP IBM Linux Certifications:

Caldera Company Phone Number:

210-348-0850 or 210-Company Fax:

210-348-0851 Support Phone Number:

210-348-0850 or 210-Contact Email Address:

techsales@txcii.com Company URL:

http://www.txcii.com Primary Contact:

Waldemar Vega

1.521 Consultant ID: US, TX, Corlogic

     1707A Brackenridge
     Austin, TX 78704
     US

Consulting Specialties:

Custom Programming Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

512-707-2347 Company Fax:

Not_Applicable Support Phone Number:

512-707-2347 Contact Email Address:

corey@corlogic.com Company URL:

http://www.corlogic.com Primary Contact:

Corey Huinker

1.522 Consultant ID: US, TX, Daily Data, Inc.

Post Office Box 140465
Dallas, TX 75214-0465
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

214.827.2170 Company Fax:

214.821.5051 Support Phone Number:

214.827.2170 Contact Email Address:

custsvc@dailydatainc.com Company URL:

http://www.dailydatainc.com Primary Contact:

Rod Rodolico

1.523 Consultant ID: US, TX, DartW Network Consulting

     P.O.Box 270362
     Flower Mound, TX 75027-0362
     US

Consulting Specialties:

Not_Applicable Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

214.793.3053 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

dart@dartw.com Company URL:

http://www.dartw.com Primary Contact:

Dart West

1.524 Consultant ID: US, TX, Internet Power And Light

19009 Preston Rd #215-240
Dallas, TX 75252-2496
US

Consulting Specialties:

Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 972 407 0727 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Phil Howard

1.525 Consultant ID: US, TX, LEAP _Linux Enterprise Application Products

P.O. Box 850905
Richardson, TX 75085-0905
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Linux Host with ASCI Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1.972.475.8440 Company Fax:

+1.972.463.2998 Support Phone Number:

+1.972.475.8440 Contact Email Address:

adcoker@leap4linux.com Company URL:

http://www.leap4linux.com Primary Contact:

Dr. Stephen F. Wheeler

1.526 Consultant ID: US, TX, PSW Technologies

6300 Bridgepoint Parkway, Bldg 3, Ste 200
Austin, TX 78730
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Not Applicable Company Phone Number:

512-343-6666 Company Fax:

512-343-9650 Support Phone Number:

512-343-6666 Contact Email Address:

linux@pswtech.com Company URL:

http://www.psw.com Primary Contact:

Satish Rao

1.527 Consultant ID: US, TX, ProSoft Plus, Inc.

     11113 Cherry Park
     San Antonio, TX 78249
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Network Administration System Administration Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

IBM Linux Certifications:

Caldera Company Phone Number:

(210)694-0801 Company Fax:

(210) 699-9416 Support Phone Number:

Not Available Contact Email Address:

prosoft@satx.net Company URL:

http://Not_Applicable Primary Contact:

Ed Valentine

1.528 Consultant ID: US, TX, Quinn Team Incorporated

12941 North Freeway
Houston, TX 77060
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Mandrake Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

281.875.3883 #211 Company Fax:

281.875.3890 Support Phone Number:

Not Available Contact Email Address:

linux@quinnteam.com Company URL:

http://www.quinnteam.com Primary Contact:

Micah Quinn

1.529 Consultant ID: US, TX, R & B Consulting

14019 SW FWY #301  PMB 411
Sugar land, TX 77478
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1-281-494-UNIX Company Fax:

1-281-494-4329 Support Phone Number:

Not_Applicable Contact Email Address:

Support@TexasComputers.com Company URL:

http://www.TexasComputers.com Primary Contact:

Brian Walters

1.530 Consultant ID: US, TX, Z Consulting

Route 1 Box 68 A-1
Hockley, TX 77447
US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

409-931-2838 Company Fax:

Not_Applicable Support Phone Number:

409-931-2838 Contact Email Address:

wczimmer@earthlink.net Company URL:

http://www.earthlink.net Primary Contact:

Chris Zimmerman

1.531 Consultant ID: US, TX, heights Open Systems

517 harvard St.
Houston, TX 77007
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Debian/GNU Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

713-864-5676 Company Fax:

713-864-5676 Support Phone Number:

713-864-5676 Contact Email Address:

perry@open-systems.com Company URL:

http://www.open-systems.com Primary Contact:

Perry Piplani

1.532 Consultant ID: US, UT, Open Source Consulting Group

5136 W. 6435 S.
West Jordan, UT 84084
US

Consulting Specialties:

CIFS (Samba) Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Any Specialty Distribution:

LinuxRouter Other Specialty Distribution:

Any Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(801) 966-1297 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

info@oscg.net Company URL:

http://www.oscg.net Primary Contact:

Mark Whitley

1.533 Consultant ID: US, VA, 2RAD

14402 Haystack Court
Centreville, VA 20120
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

703 502-0640 Company Fax:

703 502 8870 Support Phone Number:

703 502-0640 Contact Email Address:

rad@2rad.net Company URL:

http://www.2rad.net Primary Contact:

Mark Rorabaugh

1.534 Consultant ID: US, VA, Atlantis Consulting

     206B 88th Street
     Virginia Beach, VA 23451
     US

Consulting Specialties:

Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

757 425-5950 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

mskar@zasoftware.com Company URL:

http://www.zasoftware.com Primary Contact:

Marc Skarshinski

1.535 Consultant ID: US, VA, Atlantis

     206B 88th Street
     Virginia Beach, VA 23451
     US

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

757 425-5950 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

mskar@zasoftware.com Company URL:

http://www.zasoftware.com Primary Contact:

Marc Skarshinski

1.536 Consultant ID: US, VA, Branch & Company

     P.O. Box 547
     Powhatan, VA 23139
     US

Consulting Specialties:

Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

804-598-2153 USA Company Fax:

804-225-8023 USA Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Carlisle Branch

1.537 Consultant ID: US, VA, Brown Technolgies Network, Inc.

     43953 Bruceton Mills Circle
     Ashburn, VA 20147-4806
     US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

703-858-0642 Company Fax:

703-858-0642 Support Phone Number:

703-858-0642 Contact Email Address:

pbrown@btechnet.com Company URL:

http://www.btechnet.com Primary Contact:

Paul Brown

1.538 Consultant ID: US, VA, Flamingo Internet Navigators

310-D Patriot lane
Williamsburg, VA 23185-2055
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+1 757 269-0350 Company Fax:

+1 757 269-0348 Support Phone Number:

+1 757 269-0350 Contact Email Address:

support@fini.net Company URL:

http://http://www.fini.net/ Primary Contact:

Christopher Hicks

1.539 Consultant ID: US, VA, HBruch & Associates

     12809 Reams Station Place
     Herndon, VA 20170
     US

Consulting Specialties:

Custom Programming Custom Systems Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

1.703.742.9546 Company Fax:

Not_Applicable Support Phone Number:

1.703.742.9546 Contact Email Address:

harris@bruch.net Company URL:

http://www.mindspring.com/˜ hbruch Primary Contact:

Harris Bruch

1.540 Consultant ID: US, VA, Jade Enterprises, Inc.

12085 Cheroy Woods Ct.
Ashland, VA 23005
US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

804-752-6688 Company Fax:

804-752-6788 Support Phone Number:

n/a Contact Email Address:

jay@jade-ent.com Company URL:

http://www.jade-ent.com Primary Contact:

Jay Campbell

1.541 Consultant ID: US, VA, Linux & Web Professionals

     PO Box 2763
     Lynchburg, VA 24501
     US

Consulting Specialties:

Custom Programming Firewalls and Internet Security Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 (973) 283-4712 Company Fax:

1 (973) 283-4712 Support Phone Number:

1 (973) 283-4712 Contact Email Address:

support@linuxwebpro.com Company URL:

http://www.linuxwebpro.com Primary Contact:

Floyd Jennings

1.542 Consultant ID: US, VA, NetSource of Virginia, Inc.

     7639 Hull Street Rd.   STE 202
     Richmond, VA 23235-6438
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

Slackware and Red Ha Specialty Distribution:

Other Other Specialty Distribution:

none currently Reseller Authorizations:

HP Linux Certifications:

Not_Applicable Company Phone Number:

1-804-674-6040 Company Fax:

1-804-674-6041 Support Phone Number:

Not_Applicable Contact Email Address:

netsource@erols.com Company URL:

http://www.erols.com/netsource Primary Contact:

Jeff Davis

1.543 Consultant ID: US, VA, Noblestar Systems Corporation

     12021 Sunset Hills Road
     Reston, VA 20190
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

703.464.4000 Company Fax:

703.464.4001 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://www.noblestar.com Primary Contact:

Jerry Matus

1.544 Consultant ID: US, VA, Phaedo Consulting, Inc.

8000 Towers Crescent Drive, Suite 1350
Vienna, VA 22182
US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

SparcLinux, AlphaLin Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

703-847-3639 Company Fax:

540-370-4185 Support Phone Number:

703-847-3639 Contact Email Address:

info@phaedo.com Company URL:

http://www.phaedo.com Primary Contact:

Ryan Quick

1.545 Consultant ID: US, VA, Signet Computers, Inc.

1901 N. Moore St. Suite 805
Arlington, VA 22209
US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Slackware Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

703 812 8585 Company Fax:

703 812 8547 Support Phone Number:

703 812 8585 Contact Email Address:

sci-info@sigcomp.com Company URL:

http://www.sigcomp.com Primary Contact:

Chuck Olczak

1.546 Consultant ID: US, VA, Spectrum Office Systems, Inc.

     11320 Random Hills Road, Suite 630
     Fairfax, VA 22030-6001
     US

Consulting Specialties:

Firewalls and Internet Security Netware Connectivity Network Administration System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

703-591-7400 Company Fax:

703-591-9780 Support Phone Number:

703-591-7400 Contact Email Address:

tech@spectrum-systems.com Company URL:

http://www.spectrum-systems.com Primary Contact:

Nancy Fallon

1.547 Consultant ID: US, VA, Two Radical Technologies, Inc.

14119-M Sullyfield Circle
Chantilly, VA 20151-1626
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

1 888 321-2RAD Company Fax:

1 703 502 8870 Support Phone Number:

1 888 321-2RAD Contact Email Address:

rad@2rad.net Company URL:

http://www.2rad.net Primary Contact:

Mark Rorabaugh

1.548 Consultant ID: US, VA, Zentropic Computing, LLC

     441-B Carlisle Dr.
     Herndon, VA 20170
     US

Consulting Specialties:

Custom Programming Custom Systems Main Distribution:

Other Other Distribution:

Zentropix RealTime L Specialty Distribution:

Other Other Specialty Distribution:

Zentropix RealTime L Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

703-471-6690 Company Fax:

703-471-2108 Support Phone Number:

Not_Applicable Contact Email Address:

info@zentropix.com Company URL:

http://www.zentropix.com Primary Contact:

David Beal

1.549 Consultant ID: US, VT, Everest Systems

     Route 100 South
     Morrisville, VT 05661
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Compaq IBM Linux Certifications:

Not Applicable Company Phone Number:

802-888-7857 Company Fax:

802-888-7857 Support Phone Number:

802-888-7857 Contact Email Address:

braman@world.std.com Company URL:

http://world.std.com/˜ braman/ Primary Contact:

Balu Raman, MSEE.,CC Veda Aranyam

1.550 Consultant ID: US, WA, Cheek Consulting

620 5th Avve W, Suite 309
Seattle, WA 98119
US

Consulting Specialties:

CIFS (Samba) E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

206 282-2892 Company Fax:

Not_Applicable Support Phone Number:

206 282-2892 Contact Email Address:

support@cheek.com Company URL:

http://www.cheek.com Primary Contact:

joseph cheek

1.551 Consultant ID: US, WA, Dreamcorp Computer Solutions

4137 University Way NE Suite 201
Seattle, WA 98105
US

Consulting Specialties:

Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

HP Linux Certifications:

Not Applicable Company Phone Number:

(206) 634-1355 Company Fax:

(206) 634-1332 Support Phone Number:

Not_Applicable Contact Email Address:

support@dreamcorp.com Company URL:

http://www.dreamcorp.com Primary Contact:

Steve Glor

1.552 Consultant ID: US, WA, Foresight Technologies

     705 204th DR SE
     Snohomish, WA 98290
     US

Consulting Specialties:

Not_Applicable Main Distribution:

Other Other Distribution:

slakware Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

425-334-0933 Company Fax:

Not_Applicable Support Phone Number:

425-334-0933 Contact Email Address:

dagnabit2@yahoo.com Company URL:

http://Not_Applicable Primary Contact:

Dan Thacker

1.553 Consultant ID: US, WA, K Development

2103 N 106th St. #C-402
Seattle, WA 98133
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

206 365 5585 Company Fax:

Not_Applicable Support Phone Number:

206 365 5585 Contact Email Address:

info@kd-dev.com Company URL:

http://www.kd-dev.com Primary Contact:

Derek Simkowiak

1.554 Consultant ID: US, WA, PeakServ

104 Prospect Street
Bellingham, WA 98225
US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

800-453-5322 Company Fax:

360-733-7818 Support Phone Number:

800-453-5322 Contact Email Address:

info@peakserv.com Company URL:

http://www.peakserv.com Primary Contact:

John Blanford

1.555 Consultant ID: US, WA, Tacoma & Puget Sound Company

4419 Harbor Ridge Road NE
Tacoma, WA 98422
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Slackware Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(253)927-4051 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

spj@tapsco.com Company URL:

http://Not_Applicable Primary Contact:

Steven Johnson

1.556 Consultant ID: US, WA, Veriteam, Inc.

     3039 173rd Ct NE
     Redmond, WA 98052-6228
     US

Consulting Specialties:

Custom Programming Internet Server Development Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

425.882.0938 Company Fax:

425.882.0938 Support Phone Number:

425.882.0938 Contact Email Address:

support@veriteam.com Company URL:

http://www.veriteam.com Primary Contact:

Bernie Thompson

1.557 Consultant ID: US, WA, buzzwerks, Inc.

1000 Second Ave, Suite 4060
Seattle, WA 98104-1094
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

UltraPenguin Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+001 206 625 0135 Company Fax:

+001 206 625 0304 Support Phone Number:

+001 206 625 0135 Contact Email Address:

support@buzzwerks.com Company URL:

http://www.buzzwerks.com Primary Contact:

Mitch Pirtle

1.558 Consultant ID: US, WI, BadgerData Internet Solutions

P.O. Box 1526
Appleton, WI 54912
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

920-830-1343 Company Fax:

920-830-1341 Support Phone Number:

920-830-1343 Contact Email Address:

support@bdis.net Company URL:

http://www.bdis.net Primary Contact:

Garrett Meiers

1.559 Consultant ID: US, WI, Midnight Software

     N11546 Old 41 Road
     Daggett,  (Michigan), WI 49821
     US

Consulting Specialties:

Custom Programming Custom Systems Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

906-639-2882 Company Fax:

906-639-2558 Support Phone Number:

Not_Applicable Contact Email Address:

midnight@alphacomm.net Company URL:

http://Not_Applicable Primary Contact:

Aaron Tuinstra

1.560 Consultant ID: US, WI, Network Integration Services

S29W30139 S. Bethesda Cir.
Waukesha, WI 53188
US

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

414-256-9617 Company Fax:

414-968-3044 Support Phone Number:

414-256-9617 Contact Email Address:

nis@pobox.com Company URL:

http://Not_Applicable Primary Contact:

Michael Bloxham

1.561 Consultant ID: US, WI, Perry Systems

     PO Box 207
     Milwaukee, WI 53201-0207
     US

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

414/803-4797 Company Fax:

NA Support Phone Number:

NA Contact Email Address:

help@perrysystems.com Company URL:

http://www.perrysystems.com Primary Contact:

James Offutt

1.562 Consultant ID: US, WI, Rindy Consulting

125 North Executive Drive; Suite 380
Brookfield, WI 53005
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq HP Linux Certifications:

Not Applicable Company Phone Number:

414-252-4273 Company Fax:

414-252-4274 Support Phone Number:

414-252-4273 Contact Email Address:

support@rindy.com Company URL:

http://www.rindy.com Primary Contact:

Troy Rindy

1.563 Consultant ID: US, WI, WhitePine Consulting

     606 N. Main St
     Fall River, WI 53932
     US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Internet (Web) Programming Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

920.484.6005 Company Fax:

Not_Applicable Support Phone Number:

920.484.6005 Contact Email Address:

mjmc@whitepine.com Company URL:

http://www.whitepine.com Primary Contact:

Not_Applicable Not_Applicable

1.564 Consultant ID: US, WY, INTERNATIONAL SCIENCE & TEKOLOGY

740 STATE HIGHWAY 230
LARAMIE, WY 82072
US

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

REDHAT && SUSE Specialty Distribution:

Other Other Specialty Distribution:

LINUX BEOWULF CLUSTE Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

307 742 5822 Company Fax:

na Support Phone Number:

307 742 5822 Contact Email Address:

maximum_linux@hotmail.com Company URL:

http://na in process Primary Contact:

Keith Schmauss

1.565 Consultant ID: US, ZZ, ETA Associates, Inc.

     2018 31st Ave S
     Seattle, ZZ Not_Applicable
     US

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

425-765-3471 Company Fax:

Not_Applicable Support Phone Number:

Not_Applicable Contact Email Address:

support@ultracode.com Company URL:

http://www.ultracode.com/ Primary Contact:

Eric Anderson

1.566 Consultant ID: United Arab Emirates, GoldenSun Internet Consulting & Research

P.O.Box 26038
Dubai, NON United States Dubai
United Arab Emirates

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

SUSE Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

971-4-2221780 Company Fax:

971-4-2210124 Support Phone Number:

971-4-2221780 Contact Email Address:

consult@goldensun.com Company URL:

http://www.goldensun.com Primary Contact:

Prabhakar James GSC

1.567 Consultant ID: United Arab Emirates, JM Infotech Solutions STPI

     P.O.Box 2687
     Dubai, NON United States 2687
     United Arab Emirates

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

971 50 6326822 Company Fax:

971 4 359295 Support Phone Number:

971 50 6326822 Contact Email Address:

mammen@emirates.net.ae Company URL:

http://www.emirates.net.ae Primary Contact:

John Mammen

1.568 Consultant ID: United Kingdom, Acetechno Solutions Ltd

     77 Delapre Drive
     Banbury, Oxon, NON United States OX16 7WS
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems System Administration Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

0044 1295 259 523 Company Fax:

0044 1295 259 523 Support Phone Number:

0044 1295 259 523 Contact Email Address:

support@acetechno.freeserve.co.uk Company URL:

http://acetechno.hypermart.net Primary Contact:

Junior Williamson

1.569 Consultant ID: United Kingdom, CiteCall Ltd.

     Claremont, Coppice Lane
     Chorley, Lancashire, NON United States PR6 9DA
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

Suse Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 1257 230 638 Company Fax:

+44 870 164 0846 Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Kevin Mulholland

1.570 Consultant ID: United Kingdom, Dome Computer Consultants Ltd

     10 Manor Road, Whitcharch-on-Thames,
     Reading, NON United States RG8 7EU
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Firewalls and Internet Security Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

Redhat Specialty Distribution:

Other Other Specialty Distribution:

Mandrake Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 (0) 7887 955 981 Company Fax:

Not_Applicable Support Phone Number:

+44 (0) 7887 955 981 Contact Email Address:

domecc@email.com Company URL:

http://Not_Applicable Primary Contact:

Phil Lewis

1.571 Consultant ID: United Kingdom, Enterprise Management Consulting

Barnaby House, Charles Dickens Terrace, Maple Rd.
London, NON United States SE20 8RE
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

RedHat Company Phone Number:

44 181 659 2000 Company Fax:

44 181 778 0101 Support Phone Number:

44 181 659 2000 Contact Email Address:

emcweb@btinternet.com Company URL:

http://www.emcuk.com Primary Contact:

Malcolm Macsween

1.572 Consultant ID: United Kingdom, FGM Technology LTD EuroDev Cybernetics LTD

     1 South Terrace
     Thirsk, NON United States YO7 1RH
     United Kingdom

Consulting Specialties:

Network Administration System Administration System Security Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

EuroDev Linux & ARM Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

0441 197 869 / 01845 Company Fax:

NA Support Phone Number:

ONLY GIVEN TO CLIENT Contact Email Address:

ONLY GIVEN TO CLIENTS Company URL:

http://Not_Applicable Primary Contact:

Not_Applicable Not_Applicable

1.573 Consultant ID: United Kingdom, GBdirect

27 Park Drive
Bradford, NON United States BD9 4DS
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Caldera Other Distribution:

ALL Specialty Distribution:

LinuxPPC Other Specialty Distribution:

ALL Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 01274 772277 Company Fax:

+44 01274 772281 Support Phone Number:

Not_Applicable Contact Email Address:

linux@gbdirect.co.uk Company URL:

http://www.gbdirect.co.uk Primary Contact:

Mike Banahan

1.574 Consultant ID: United Kingdom, J.S Unix Consultants Ltd

     1073 Manchester Road, Castleton
     Rochdale, NON United States OL11 2XJ
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

Suse Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

(+44) 958 402172 Company Fax:

(+44) 970 0548678 Support Phone Number:

(+44) 958 402172 Contact Email Address:

support@santoroj.demon.co.uk Company URL:

http://www.netspear.co.uk/js_unix Primary Contact:

Joe Santoro

1.575 Consultant ID: United Kingdom, Jason Clifford

264 Ripon Road
Stevenage, NON United States SG1 4NF
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

Other Other Distribution:

All distributions Specialty Distribution:

Other Other Specialty Distribution:

All Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 7957 356660 Company Fax:

+44 1438 223433 Support Phone Number:

Not Available Contact Email Address:

support@dlsl.demon.co.uk Company URL:

http://www.dlsl.demon.co.uk Primary Contact:

Jason Clifford

1.576 Consultant ID: United Kingdom, Labrat Computing Ltd

     86 Crayford Way
     Dartford, NON United States DA1 4LQ
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 370 837124 Company Fax:

+44 870 050 9749 Support Phone Number:

n/a Contact Email Address:

support@labrat.cx Company URL:

http://www.labrat.cx Primary Contact:

Russ Spooner

1.577 Consultant ID: United Kingdom, Lightwood Consultancy Ltd

     PO Box 1387
     Stoke-on-Trent, NON United States ST3 5YH
     United Kingdom

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming System Administration

Main Distribution:

Other Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 870 7402748 Company Fax:

+44 8700 523955 Support Phone Number:

Not_Applicable Contact Email Address:

support@lightwood.net Company URL:

http://www.lightwood.net Primary Contact:

Chris Newman

1.578 Consultant ID: United Kingdom, MPC Data Limited

     The Old Gas Warehouse, Frome Road
     Bradford on Avon, Wiltshi, NON United States BA15 1HA
     United Kingdom

Consulting Specialties:

Custom Programming Custom Systems Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

Embedded Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 1225 868228 Company Fax:

+44 1225 868191 Support Phone Number:

Not_Applicable Contact Email Address:

jback@mpc-data.co.uk Company URL:

http://www.mpc-data.co.uk Primary Contact:

Julian Back

1.579 Consultant ID: United Kingdom, Michiel Dethmers

Quipu, 91 Brick Lane
London, NON United States E1 6QL
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

44 171 3776601 Company Fax:

44 171 3776601 Support Phone Number:

Not_Applicable Contact Email Address:

support@quipu.co.uk Company URL:

http://www.quipu.co.uk Primary Contact:

Michiel Dethmers

1.580 Consultant ID: United Kingdom, Oakhaven Consultants Ltd

     Bunces Lane, Burghfield Common
     Reading., NON United States RG7 3DL
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet (Web) Programming Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+441189831707 Company Fax:

+448700523200 Support Phone Number:

Not_Applicable Contact Email Address:

Not_Applicable Company URL:

http://www.thennion.demon.co.uk Primary Contact:

Trevor Hennion BSc.

1.581 Consultant ID: United Kingdom, Philip Hands Computing Ltd

     Unit 1, Cherry Close, Caversham
     Reading, NON United States RG4 8UP
     United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development System Security

Main Distribution:

Debian Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not_Applicable Linux Certifications:

Not_Applicable Company Phone Number:

+44 118 9545656 Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

Not_Applicable Company URL:

http://Not_Applicable Primary Contact:

Philip Hands

1.582 Consultant ID: United Kingdom, RedAmazon

Not_Applicable
London, NON United States Not_Applicable
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxRouter Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

Not_Applicable Company Fax:

Not_Applicable Support Phone Number:

Not Available Contact Email Address:

support@redamazon.com Company URL:

http://www.redamazon.com Primary Contact:

Jim Thoreaux

1.583 Consultant ID: United Kingdom, Share Net International Ltd

Belle Vue, Crowborough
East Sussex, NON United States TN6 1ST
United Kingdom

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+044 01892 661 044 Company Fax:

+044 01892 665 731 Support Phone Number:

+044 01892 661 044 Contact Email Address:

support@share-net.co.uk Company URL:

http://www.share-net.co.uk Primary Contact:

Grace Nunn

1.584 Consultant ID: United Kingdom, iCode Systems Ltd

     Consort Court, High Street
     Fareham, NON United States PO16 7AL
     United Kingdom

Consulting Specialties:

Custom Programming E-Commerce Firewalls and Internet Security Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Compaq Linux Certifications:

Not_Applicable Company Phone Number:

+44 1329 829898 Company Fax:

+44 1329 829899 Support Phone Number:

+44 1329 829898 Contact Email Address:

linux@icode.co.uk Company URL:

http://www.icode.co.uk Primary Contact:

Darren Durbin

1.585 Consultant ID: United Kingdom, thorNET

     34 St Mary Street, Thornbury
     Bristol, NON United States BS35 2AT
     United Kingdom

Consulting Specialties:

Custom Programming Custom Systems E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+44 1454854413 Company Fax:

+44 1454854412 Support Phone Number:

+44 1454 854413 Contact Email Address:

info@thornet.co.uk Company URL:

http://www.thornet.co.uk Primary Contact:

Steve Heaven

1.586 Consultant ID: Venezuela, Ernesto Hernández-Novich

Av. Ppal. La Bonita Res. La Montaña Apto. 7-F
Caracas, NON United States 1080
Venezuela

Consulting Specialties:

CIFS (Samba) Custom Programming Custom Systems Firewalls and Internet Security Internet Server Development Internet (Web) Programming Netware Connectivity Network Administration System Administration System Security

Main Distribution:

Other Other Distribution:

All of them Specialty Distribution:

Other Other Specialty Distribution:

None Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

58-2-9414198 Company Fax:

Not_Applicable Support Phone Number:

58-14-9254779 Contact Email Address:

emhn@telcel.net.ve Company URL:

http://www.telcel.net.ve Primary Contact:

Ernesto Hernández-Novich

1.587 Consultant ID: Venezuela, Grupo INABOX,C.A.

     Torre Capriles, Planta Baja, Local 13
     Caracas, NON United States Not_Applicable
     Venezuela

Consulting Specialties:

System Administration Main Distribution:

Caldera Other Distribution:

NA Specialty Distribution:

LinuxPPC Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Caldera Company Phone Number:

005827934692 Company Fax:

005827934692 Support Phone Number:

005827934692 Contact Email Address:

inaboxca@yahoo.com Company URL:

http://www.yahoo.com Primary Contact:

Gilberto Hernandez Meza

1.588 Consultant ID: Venezuela, Linux Sistemas, C.A.

     Paseo Los Ilustres, Ed. Don Paco, Local 1, Los Cha
     Caracas, NON United States 1041
     Venezuela

Consulting Specialties:

CIFS (Samba) Firewalls and Internet Security Internet Server Development Netware Connectivity Network Administration System Administration System Security

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Compaq IBM Linux Certifications:

Not Applicable Company Phone Number:

+58 2 6934233 Company Fax:

+58 2 6623580 Support Phone Number:

+58 2 6934233 Contact Email Address:

linuxsistemas@mercared.com Company URL:

http://http://www.mercared.com/linuxsistemas Primary Contact:

Jose Rey

1.589 Consultant ID: Vietnam, Nam Ltd

     14 Chuong Duong
     Hanoi, NON United States HAN
     Vietnam

Consulting Specialties:

CIFS (Samba) E-Commerce Internet Server Development Internet (Web) Programming Network Administration System Administration

Main Distribution:

RedHat Other Distribution:

NA Specialty Distribution:

Other Other Specialty Distribution:

NA Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

84-4-8242294 Company Fax:

84-4-8242294 Support Phone Number:

NA Contact Email Address:

xuanbn@vol.vnn.vn Company URL:

http://NA Primary Contact:

Bui nam Xuan

1.590 Consultant ID: Yugoslavia, Berislav Todorovic

     Elektrotehnicki fakultet, Bulevar Revolucije 73
     Belgrade SERBIA, NON United States 11000
     Yugoslavia

Consulting Specialties:

CIFS (Samba) Custom Systems Firewalls and Internet Security Internet Server Development Network Administration System Administration System Security Virtual Private Networking

Main Distribution:

RedHat Other Distribution:

Not_Applicable Specialty Distribution:

Other Other Specialty Distribution:

Slackware Reseller Authorizations:

Not Applicable Linux Certifications:

Not Applicable Company Phone Number:

+381-11-3221-419 Company Fax:

+381-11-3248-681 Support Phone Number:

Not_Applicable Contact Email Address:

beri@etf.bg.ac.yu Company URL:

http://ubbg.etf.bg.ac.yu/˜ beri/ Primary Contact:

Berislav Todorovic